### v1

# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

Abstract
Agentic AI systems, built on large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligent autonomy, collaboration and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based agentic multi-agent systems (AMAS). We begin by examining the conceptual foundations of agentic AI, its architectural differences from traditional AI agents, and the emerging system designs that enable scalable, tool-using autonomy. The TRiSM in the agentic AI framework is then detailed through four pillars governance, explainability, ModelOps, and privacy/security each contextualized for agentic LLMs. We identify unique threat vectors and introduce a comprehensive risk taxonomy for the agentic AI applications, supported by case studies illustrating real-world vulnerabilities. Furthermore, the paper also surveys trust-building mechanisms, transparency and oversight techniques, and state-of-the-art explainability strategies in distributed LLM agent systems. Additionally, metrics for evaluating trust, interpretability, and human-centered performance are reviewed alongside open benchmarking challenges. Security and privacy are addressed through encryption, adversarial defense, and compliance with evolving AI regulations. The paper concludes with a roadmap for responsible agentic AI, proposing research directions to align emerging multi-agent systems with robust TRiSM principles for safe, accountable, and transparent deployment.
00footnotetext: *Shaina Raza and Ranjan Sapkota contributed equally and are corresponding authors.
Emails: shaina.raza@torontomu.ca, rs2672@cornell.edu, mk2684@cornell.edu, c.emmanouilidis@rug.nl.
Keywords Agentic AI, LLM-based Multi-Agent Systems, TRISM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, Autonomous Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
1 Introduction
1.1 Background and Motivation
The global market for AI agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  miller2025aiagents . By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models lyzr2025state . An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  russell1997rationality , has undergone a fundamental transformation. Whereas early agents were task-specific and deterministic, modern systems, referred to as Agentic AI, now integrate large language models (LLMs), tools, and persistent memory to support complex planning and coordination sapkota2025ai ; de2023emergent .
From isolated agents to Agentic AI. The autonomous software agents date back to the 1990s castelfranchi1998modelling , the advent of LLM-driven multi-agent ecosystems yang2024multi since 2023 has made “Agentic AI” a distinctly new and rapidly expanding research focus sapkota2025ai . Traditional agents automated narrow tasks, such as information retrieval chiu2007biomimetic ; lew2006content , data summarization tanasa2005advanced ; devault2014simsensei ; hao2012mri , and dialogue response griol2013automatic . Driven by single step logic genesereth2012logical or scripted rules, they lacked deep reasoning, adaptability, and persistence spronck2006adaptive ; szafir2012pay .
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use acharya2025agentic . These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines qian2023chatdev ; wu2023autogen ; hong2023metagpt . This shift represents not just a technological upgrade but a qualitative leap in complexity and autonomy, marking the emergence of machine collectives capable of emergent, decentralized behavior.
Trust, risk, and security challenges. However, this evolution introduces serious challenges. Unlike deterministic agents zeigler2014object , Agentic AI systems can produce non-linear, opaque decisions, increasing the risks of failure, bias, and unintended consequences. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents, yet inadvertently leak sensitive information or violate compliance protocols if safeguards are absent lim2012memory .
Traditional evaluation and safety frameworks, built for static or single-function AI , are no longer sufficient. This underscores the urgent need for a new paradigm that integrates trust, risk, and security as core design principles. To address this gap, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) IBM2025AITRiSM ; habbal2024artificial propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management. These are essential for deploying agentic systems in high-stakes domains such as finance, healthcare, and defense.
Problem Statement Agentic AI systems now coordinate multiple LLM-powered agents, yet no unified framework exists for managing their trust, risk, and security (TRiSM). High-profile lapses, e.g., an autonomous research agent citing unverifiable medical sources or a forecasting swarm drifting off-target due to memory inconsistencies show that traditional safeguards do not scale to dynamically evolving, multi-agent workflows. A structured TRiSM perspective is therefore crucial to mitigate orchestration failures, collusion, data leakage, and opaque decision paths. This review distills current findings and outlines practical guidelines to build secure, reliable, and trustworthy Agentic AI ecosystems.
Scope and Objectives This survey maps the TRiSM (Trust, Risk, and Security Management) landscape for LLM-powered, multi-agent “agentic” AI. We frame four pillars explainability, security, lifecycle governance, and privacy and show why safeguards designed for bounded, rule-based agents cannot contain open-world, self-orchestrating collectives sapkota2025ai ; acharya2025agentic . Our goal is to give researchers and practitioners a concise, unified reference for building and governing trustworthy Agentic AI systems.
Necessity of this Survey Agentic AI research often concentrates on agent modelling or task coordination and rarely tackles adversarial threats, lifecycle governance, or large-scale explainability. As deployments reach high-stakes arenas healthcare, science, finance the absence of a unified TRiSM framework exposes stakeholders to opaque reasoning and unmanaged risks. This survey closes that gap by synthesising TRiSM principles for LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policy makers. A comparison with related surveys will be presented in Table 1.
1.2 Contributions
In this survey, we address the urgent need for a unified governance framework for LLM-based Agentic AI systems. Our framework is shown in Figure 1. Our key contributions are as follows:
  • A Conceptual TRISM Framework Tailored to Agentic AI We introduce a structured Trust, Risk, and Security Management (TRISM) framework specifically designed to accommodate the unique characteristics of multi-agent, LLM-driven systems. This framework defines four pillars: Explainability, ModelOps, Application Security, and Model Privacy, and explains how each pillar addresses the challenges of autonomy, coordination, and persistent memory in Agentic AI.
  • A Threat Taxonomy for LLM-Powered Agents We develop a detailed taxonomy of risks and threat vectors that arise when multiple LLM agents collaborate. By categorizing vulnerabilities, such as prompt injection, memory poisoning, collusive failures, and emergent misbehavior, we clarify how attack surfaces expand in agentic environments and why traditional LLM security measures are insufficient.
  • A Comparative Evaluation of Existing Methods We survey and compare state-of-the-art techniques for each TRISM pillar as applied to Agentic AI. For example, we examine how explainable-AI methods (e.g., LIME, SHAP, decision provenance) have been adapted for multi-agent workflows, how ModelOps best practices must be extended to manage multiple LLM instances, and which security/hardening measures (e.g., prompt hygiene, sandboxing, access controls) are most effective.
  • Future Research Directions for Trustworthy, Secure Agentic AI Building on our TRISM analysis, we outline concrete directions for next-generation research and setup roadmap items aim to guide researchers and practitioners toward robust, scalable, and compliant deployments of Agentic AI.
This study provide a unified reference for researchers, engineers, and policymakers who seek to design, evaluate, and govern LLM-based Agentic AI systems in high-stakes domains.
1.3 Paper Organization
The remainder of this paper is structured as follows. In Section 2, we present our literature-search methodology, including databases queried, inclusion criteria, and classification strategies. Section 3 reviews the fundamental characteristics and typical architectures of Agentic AI systems, establishing the groundwork for subsequent TRISM analysis. Section 4 introduces the TRISM framework and its four core pillars : Explainability, ModelOps, Application Security, and Model Privacy. Building on this foundation, Section 5 identifies the primary risks, vulnerabilities, and attack surfaces inherent in multi-agent AI ecosystems. In Section 6, we explore trust-building strategies and specialized explainability techniques tailored to Agentic AI, while Section 7 details advanced security mechanisms and privacy-preserving methods. Section 8 synthesizes our findings and outlines future research directions and policy recommendations. Section 9 concludes the paper. A list of key terms used throughout this paper is provided in Table 9.
1.4 Related Work
Table 1: Comparison of Related Surveys on LLM-based Multi-Agent Systems and TRiSM Aspects
Survey Adversarial Threats Lifecycle Governance Large-Scale Explainability TRiSM Integration LLM-Specific Application Domains Actionable Guidance
Guo et al. (2024) guo2024largelanguagemodelbased (simulated env’ts) ~ (research challenges)
Chen et al. (2025) chen2025surveyllmbasedmultiagentsystem ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) yan2025selftalkcommunicationcentricsurveyllmbased (mentions diverse) ~ (future work directions)
Tran et al. (2025) tran2025multiagentcollaborationmechanismssurvey (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) lin2025creativityllmbasedmultiagentsystems ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) fang2025trustworthyaisafetybias ~ ~ (health, finance cited) ~ (technical “next steps”)
This Survey (2025) (high-stakes domains) (practical guidance)
Legend: ✓ = explicitly addressed; ~ = partially/indirectly addressed; ✗ = not addressed.
LLM-Based Multi-Agent Surveys (Technical Focus): Prior surveys on LLM-driven multi-agent systems primarily emphasize system architectures, agent capabilities, and domain-specific applications of agentic AI, but often overlook critical TRiSM considerations. For instance, existing reviews guo2024largelanguagemodelbased ; chen2025surveyllmbasedmultiagentsystem focus on simulated environments, inter-agent communication, and performance benchmarks, yet fail to address adversarial threats, lifecycle governance, or system-wide explainability. Other targeted surveys, such as yan2025selftalkcommunicationcentricsurveyllmbased on natural language communication and tran2025multiagentcollaborationmechanismssurvey on collaboration mechanisms, explore interaction protocols and use cases in domains like QA systems, 5G, and Industry 5.0. While the former briefly acknowledges security concerns and the latter highlights deployment contexts, both narrowly center on coordination strategies without examining robustness, governance structures, or explainability at scale.
Trustworthy/Responsible AI Surveys (Broad Trust Focus): Surveys in the TRiSM domain address ethical and risk-related concerns such as alignment, fairness, and privacy attacks fang2025trustworthyaisafetybias , but typically overlook explainability, governance, and multi-agent dynamics. While they offer actionable recommendations for general ML systems, they do not account for LLM-based coordination in high-stakes domains.
Current Survey (LLM Multi-Agent + TRiSM in High-Stakes Domains): This survey uniquely integrates TRiSM principles into the analysis of LLM-based multi-agent systems, specifically targeting high-stakes domains such as healthcare, science, and finance. In contrast to earlier surveys (shown in Table 1) that underemphasize governance and security, this work foregrounds adversarial threats (e.g., vulnerabilities in agent interaction), end-to-end lifecycle governance (from training data to deployment oversight), and large-scale explainability. These concerns are treated as integral to system design, not peripheral. Additionally, this survey provides actionable recommendations for researchers, engineers, and policymakers, bridging gaps left by both technical and responsible-AI surveys.
2 Literature Review Methodology
To ensure a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology inspired by best practices in systematic reviews  boland2017doing ; keele2007guidelines . This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
Research Objectives
The review was guided by the following questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems?
  • RQ2: What technical and governance strategies have been proposed to address these challenges?
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability, model operations, security, and privacy?
  • RQ4: What gaps remain in current research, and what directions are promising for future work?
Data Sources and Search Strategy
We searched the following major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar, covering publications from January 2020 to May 2025. The search was executed using Boolean combinations of keywords relevant to Agentic AI and TRiSM, including:
"agentic AI" OR multi-agent systems "(AMAS)" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents") AND ("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection")
In total, 150 unique papers were initially retrieved.
Inclusion and Exclusion Criteria
We applied the following criteria during the screening process:
  • Inclusion:
    • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-chatgpt).
    • Studies that explicitly discuss agentic, multi-agent, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
    • Peer-reviewed articles, preprints, or whitepapers from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
  • Exclusion:
    • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
    • Studies addressing only low-level ML components (e.g., training optimization, architecture design) without reference to agentic behavior or TRiSM concerns.
    • Non-English papers or those lacking sufficient metadata (e.g., missing abstracts or publication details).
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant white papers, technical blogs, and practitioner reports to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework.
Quality Assurance
To ensure the rigor and reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines kitchenham2004procedures ; keele2007guidelines . Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by two reviewers, with discrepancies resolved through discussion.
Figure 1: Taxonomy of Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems (Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks (Unique Threat Vectors, Taxonomy of Risks, Case Studies), Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
3 Fundamentals of Agentic AI Systems
3.1 Agentic AI vs. Traditional AI Agents
The distinction between traditional AI agents and Agentic AI systems represents a paradigm shift in autonomous system design. Traditional agents operate through predefined rules muscettola1998remote , heuristic workflows hannebauer1999formal ; cranshaw2017calendar , or deterministic logic govaerts2010development , excelling in narrow, bounded environments. In contrast, Agentic AI systems leverage foundation models (LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: 1) Multi-agent coordination where specialized agents (e.g., planners, verifiers) collaborate via structured protocols qian2023chatdev ; hong2023metagpt 2) Persistent context through memory architectures that maintain state across workflows borghoff2025human 3) Dynamic meta-orchestration that delegates tasks and resolves conflicts schneider2025generative We present a comparison of AI agents and Agentic AI in Table 2.
This architecture enables longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches de2025open . Systems like AutoGen wu2023autogen demonstrate how LLM-backed agents collectively decompose complex problems through multi-step reasoning qiao2024agent , marking a qualitative leap from reactive agents to proactive problem-solving systems.
Table 2: Traditional AI Agents vs. Agentic AI Systems
Dimension Traditional AI Agents Agentic AI Systems
Autonomy Model × Reactive execution via scripts/rules Goal-driven planning & adaptation
$\bullet$ Fixed action sequences $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic & FSMs Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded KBs $\bullet$ Emergent reasoning
Intelligence Scope × Narrow & task-specific Broad & compositional
$\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
Reasoning Approach × Deterministic/single-step Multi-step & contextual
$\bullet$ Rule-based inference $\bullet$ Chain-of-Thought (CoT)
Collaboration × Isolated execution Role-specialized coordination
$\bullet$ Manual decomposition $\bullet$ Automated hierarchy
Temporal Context × Episodic & stateless Persistent memory
$\bullet$ Session-bound $\bullet$ VectorDB/LTM
Orchestration × Hard-coded pipelines Dynamic meta-agents
$\bullet$ Sequential workflows $\bullet$ Conflict resolution
Tool Utilization × Static integrations Planned API invocation
$\bullet$ Handcrafted interfaces $\bullet$ Function calling
Context Awareness × Bounded inputs Memory-augmented
$\bullet$ Limited context $\bullet$ RAG architectures
Scalability × Domain-specific Cross-domain transfer
$\bullet$ High maintenance $\bullet$ Modular composition
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ Finite-state bots $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
= Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
3.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 2, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation wang2025comprehensive . These agents access a Shared Toolchain Interface speight2025build to execute code, perform searches, or interact with domain-specific APIs.
Communication and coordination are facilitated through a Communication Middleware weyns2009agent , allowing agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator langgraph2024 governs high-level planning, delegating subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory guo2024largelanguagemodelbased , which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface cranshaw2017calendar , enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit Module monitors agent actions, logs tool usage, and generates behavioral traces kaur2024building . However, this modular and distributed structure introduces significant TRiSM charli_trism_2024 challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing zhong2023towards .
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information feretzakis2024privacy , especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration singh2025explainable . Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors.
Figure 2: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter- agent synerg
Below, we discuss the architecture of AMAS.
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” Saravia_LLM_Agents_2024 . The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as significant_gravitas_autogpt_2023 , Baby AGI nakajima_babyagi_archive_2024 and GPT Engineer osika_gpt_engineer_2023 , the LLM governs the full control loop, orchestrating the overall system behavior.
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting Saravia_Prompt_Engineering_Guide_Web_2024 , where the model performs intermediate reasoning before arriving at a final decision. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like ReAct yao2023reactsynergizingreasoningacting exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) yang2024multi . Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM’s prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface schick2023toolformerlanguagemodelsteach . This layer allows invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like MRKL karpas2022mrklsystemsmodularneurosymbolic illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer schick2023toolformerlanguagemodelsteach train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation - action interface is essential liu2025agent . Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results acharya2025agentic . Each iteration enriches the agent’s context, enabling it to self-prompt, generate subtasks, assess progress, and adapt strategies over time. This integrated design empowers agentic systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory, while Voyager couples the same core with an embodied Minecraft API that serves as both tool layer and environment interface. We comparing current Agentic AI implementations and identifying open design trade-offs in this table.
Table 3: Representative LLM-Based Agentic AI Frameworks
Framework (Year) Core LLM Planning Memory Tool Use Notable Features
AutoGPT yang2023auto GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI nakajima_babyagi_archive_2024 GPT-3.5/4 Task queue In-memory Web search Minimal task generator
GPT Engineer gpt-engineer GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
LangGraph wang2024agent Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
AutoGen wu2023autogen GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
MRKL karpas2022mrklsystemsmodularneurosymbolic LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
Reflexion shinn2023reflexionlanguageagentsverbal GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
MetaGPT hong2024metagpt GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
Voyager wang2023voyageropenendedembodiedagent GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
WebVoyager he2024webvoyager LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
HuggingGPT shen2023hugginggpt ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
CAMEL li2023camel GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
ChatDev qian2024chatdevcommunicativeagentssoftware GPT-4 Chat chain File repo Unix tools Simulated software development workflow
CrewAI crewAI Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
AgentVerse chen2023agentverse Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
OpenAgents xie2023openagentsopenplatformlanguage Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI superagi GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
4 The TRISM Framework for Agentic AI
4.1 Governance for Agentic LLM Systems
AI Trust, Risk, and Security Management (AI TRISM) is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards habbal2024artificial . Originally highlighted in industry guidelines for AI governance litan2024ai , TRISM provides a structured approach to manage the unique challenges of LLM-based “agentic” AI systems. Such systems consist of autonomous LLM agents that can make independent decisions, collaborate with other agents, and adapt their behavior over time kaur2024building . These properties: autonomy, multi-agent interaction, and evolving behavior, introduce new risks not seen in traditional single-model deployments ko2025sevensecuritychallengessolved .
For example, an agent acting in isolation might be benign, but when interacting with others across organizational or trust boundaries, it could manipulate peers or leak confidential information. The TRISM framework addresses these concerns by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy kaur2024building ; habbal2024artificial . Each pillar targets a critical aspect of safety or risk management, ensuring that an agentic LLM system remains transparent, reliable, secure, and compliant with ethical and regulatory requirements. Below, we define each pillar and explain how it applies to LLM-based agentic systems, grounding the discussion in current research and best practices.
4.2 Explainability in Multi-Agent Decision Making
Definition and Importance Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of agentic LLM systems, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction rosenfeld2019explainability . Accordingly, the TRISM framework elevates explainability as a core pillar to ensure that each agent’s actions and the overall system’s behavior can be understood and audited.
Techniques for Explainability in Agentic Systems Achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes. Established explainable AI techniques provide a starting point. For instance, Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and Shapley Additive Explanations (SHAP) lundberg2017unified can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an agentic setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered kusner2017counterfactual , for example, if a particular agent’s contribution were removed or modified. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making. For instance, by systematically toggling an agent off or varying its outputs, one can observe changes in the collective outcome and thus determine that Agent X was critical in influencing decision Y. Such analyses surface the dependencies and influences among agents, effectively explaining emergent behaviors at the system level. Moreover, recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g. chain-of-thought prompts or dialogue between agents) can further enhance transparency. Human auditors can also help with a trace of how agents arrived at a decision, e..g, which agent contributed what information and why, to produce narrative explanations for its outcomes.
Summary In summary, the explainability pillar of TRISM compels the use of these techniques (surrogate models, feature attributions, counterfactual testing, and transparent reasoning traces) to ensure that even highly autonomous LLM agents remain interpretable and accountable to human oversight.
4.3 ModelOps : Lifecycle Management for LLM Agents
Definition and Scope ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement lefevre2022modelops . It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production. Within agentic LLM systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely kaur2024building . Effective ModelOps is crucial for maintaining consistency, performance, and regulatory compliance as LLM agents evolve or as new agents are added to a system.
Application to Agentic Systems LLM-based agents require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental drift. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions. This aligns with best practices in LLM operations (LLMOps) mlops , which integrate MLOps principles tailored to the challenges of large language models
Challenge One challenge with this method is model drift, i.e., over time, an agent’s responses may become less accurate or relevant as data distributions shift or real-world conditions change. Continuous monitoring is therefore required to detect performance degradation or deviations from expected behavior, triggering retraining or recalibration when needed. Moreover, real-time monitoring and logging are fundamental in multi-agent settings. Each agent’s actions (e.g. API calls, decisions taken, errors encountered) are logged and analyzed to provide observability into the system’s functioning.
In a large-scale agent ecosystem, orchestration services may oversee the agents, scheduling their tasks and managing inter-agent communication. ModelOps must govern these orchestration layers as well, ensuring, for instance, that if one agent fails or produces suspicious output, it can be isolated or shut down without collapsing the whole system.
Summary In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable. It formalizes change management (so updates do not introduce new risks), provides continuous validation of model behavior, and supports compliance by logging data for audits and enforcing policies (e.g. preventing unauthorized model changes).
4.4 Application Security for LLM Agents
Threats in Agentic Environments Application security in the TRISM framework focuses on protecting AI agents and their ecosystem from malicious attacks and misuse. LLM-based agents are susceptible to a range of novel security threats that exploit their language-based interfaces and cooperative behaviors. One well-documented threat is prompt injection kosinski2024prompt , wherein an attacker designs input data that contains hidden or malicious instructions . Recent studies lee2024promptinfectionllmtollmprompt have shown that in AMAS, such prompt injections can propagate from one agent to another, a phenomenon dubbed “prompt infection” is analogous to a computer virus spreading across networks raza2025justhumansneedvaccines . In a prompt infection scenario, a malicious prompt introduced to Agent A might covertly modify Agent A’s output, which then becomes part of Agent B’s input, thereby tricking Agent B as well, and so on. This cascading attack can lead to widespread data leaks, fraudulent transactions, misinformation, or coordinated misbehavior across an agent society.
Another critical vulnerability is identityspoofing and impersonation chen2025aiagents . In a multi-agent system, agents often communicate or coordinate tasks, and they may rely on credentials or tokens to authenticate one another. For example, if an adversary steals an agent’s API key or tricks the system into treating a rogue model as a trusted peer, they could issue commands or receive information under a false identity.
Defensive Measures: To mitigate these threats, TRISM’s security pillar mandates a defense-in-depth approach tailored to LLM agents:
  1. 1.
    Prompt hygiene: agents should treat inputs defensively by sanitizing and filtering prompts, and using guardrails or content policies to detect and refuse suspicious instructions. Prompt hardening (e.g., adding secure prefixes or validation steps) is one such method to make agents less susceptible to injection gojsalic2024system .
  2. 2.
    Strong authentication and access controls Each agent and human user must be securely authenticated, and least-privilege principles owasp2025agentic should constrain what actions an agent can perform autonomously.
  3. 3.
    Continuous monitoring If an agent suddenly starts issuing unusual requests or deviates from its normal pattern of behavior, automated monitors can flag this for investigation or trigger an automatic shutdown of the agent’s actions.
Recent frameworks, such as LangChain/LangFlow langchain2024function , AutoGen autogen_agents_tutorial , CrewAI crewAI , introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) raza2025justhumansneedvaccines can improve their resistance to malicious inputs
Summary In sum, the Application Security pillar of TRISM emphasizes proactive safeguards against both external attackers and potential rogue agents. By implementing strict authentication, input validation, encrypted communication, sandboxing of execution (for agents that can use tools or code), and comprehensive monitoring, organizations can significantly reduce the risk of prompt-based exploits, impersonation, and other lateral vulnerabilities that are unique to autonomous multi-agent AI systems hal.science. This layered security approach is essential to maintain the integrity and reliability of agentic LLM deployments in adversarial environments.
4.5 Model Privacy and Data Protection
Challenges in Data Sharing: The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms. LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks ko2025sevensecuritychallengessolved . In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. Therefore, TRISM’s privacy pillar compels organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
Techniques for Privacy Preservation
  • Differential Privacy (DP): Injects calibrated noise during model training to prevent memorization of individual data entries, ensuring no single record significantly influences the output dwork2006calibrating .
  • Data Anonymization and Minimization: Limits inter-agent data sharing to only what is necessary sweeney2002k , often using aggregated or pseudonymized formats (e.g., “age bracket 30-40” instead of exact birthdate).
  • Secure Multi-Party Computation (SMPC): Enables agents to compute joint functions without exposing private inputs, useful in cross-organization tasks like collaborative fraud detection lindell2009secure .
  • Homomorphic Encryption (HE): Allows agents to compute on encrypted data gentry2009fully . With Fully Homomorphic Encryption (FHE), even the plaintext query and response remain unseen by the agent.
  • Trusted Execution Environments (Secure Enclaves): Hardware-based isolation ensures that even privileged system users cannot access the data being processed by agents costan2016intel . Useful for secure memory sharing and execution.
  • Model Privacy Policies and Compliance: Enforces data retention limits, maintains audit logs, and ensures compliance with regulations (e.g., GDPR, HIPAA) governing agent behavior and data usage Responsi73:online .
Summary By implementing these layers of privacy defense , from differential privacy in model training to homomorphic encryption for data sharing, and stringent access control policies, Agentic AI systems can protect user data and proprietary information even as they leverage that data for intelligent decision-making.
The TRISM framework offers a comprehensive governance model for LLM-based Agentic AI, integrating Explainability, ModelOps, Application Security, and Model Privacy to manage the complexity of autonomous agent systems. Grounded in proven methods, such as SHAP, CI/CD, adversarial defenses, and homomorphic encryption, TRISM enhances safety, transparency, and trust. As AI systems evolve, TRISM provides a stable foundation to ensure responsible and secure agent behavior, aligning advanced capabilities with human values and operational integrity.
Table 4 condenses each pillar into its governing goal.
Table 4: TRISM pillars for LLM-based Agentic AI and associated governance references.
Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” dotan2024evolving
ISO/IEC 24029-1 iso24029-1
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) iso42001
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps owasp2024llmtop10
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 gdpr25
HIPAA §164 hipaa164
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
5 Threats and Risks in LLM-based (Agentic) AMAS
In this section, we discuss the threats and risks in AMAS.
5.1 Unique Threat Vectors
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures deng2025ai . These risks primarily arise from agents’ autonomy, persistent state management, and the complex demands of multi-agent coordination sapkota2025ai . Below, we discuss these threats
  • Autonomy abuse The foremost threat is autonomy abuse, whereby agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs patlan2025real ; kolt2025governing . Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states hammond2025multi ; wilson2025age .
  • Persistent memory Another threat is the persistent memory, which while, crucial for context retention, introduces unique vulnerabilities through potential adversarial injections and accumulations du2025rethinking ; dechant2025episodic . Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms patlan2025real ; deng2025ai .
  • Agent orchestration This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT hong2023metagpt and AutoGen wu2023autogen . These orchestration vulnerabilities differentiate agentic systems markedly from conventional stateless, single-threaded LLM deployments.
5.2 Taxonomy of Risks
To systematically understand the security landscape in Agentic AI, we categorize risks into four broad classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors.
  • Adversarial Attacks Agents remain vulnerable to prompt injections, gradient-based manipulations, and engineered reasoning traps, risks that are magnified in AMAS due to propagation across agent interactions miao2025autonomous ; liu2025compromising . An illustrative instance is the role-swapping attack observed in ChatDev qian2023chatdev .
  • Data Leakage Persistent memory and extensive inter-agent communication elevate the likelihood of accidental exposure of sensitive information wang2025ip ; mao2025agentsafe . In sensitive domains like financial services and HR, inadequate boundary enforcement and ineffective sanitization amplify these leakage risks tran2025multi ; xu2025forewarned .
  • Agent Collusion and Mode Collapse Coordination mechanisms can inadvertently lead agents to reinforce mutual errors, precipitating groupthink or echo chambers feng2025multi ; cui2025towards . AutoGen experiments illustrate how iterative dialogues among agents can amplify flawed designs, highlighting the risk of emergent misalignment wu2023autogen ; reddy2025enhancing ; cihon2025measuring .
  • Emergent Behavior Complex interactions among agents, memory components, tools, and tasks yield unpredictable behaviors that evade traditional testing and validation methods dawid2025agentic ; la2025large . Agents optimizing for efficiency may unintentionally bypass critical verification steps or suppress contradictory information, scenarios exemplified in blockchain karim2025ai and audio verification contexts li2025audiotrust .
5.3 Case Studies
Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental agentic systems.
Case Study 1: Prompt Leakage in Agentic Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT, where recursive prompt augmentation and insufficient memory controls can lead to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard agentic systems from cascading information leaks agarwal-etal-2024-prompt .
Case Study 2: Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of faulty design assumptions. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario underscores the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures qian2024chatdevcommunicativeagentssoftware ; lee2024promptinfectionllmtollmprompt .
Case Study 3: Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion. This incident underscores the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior StrDorFri2024 ; ligot2022using .
Case Study 4: Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident underscores the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities atkins2023memorypoisoning .
Case Study 5: System Prompt Drift in Autonomous Memory Agents In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions liu2023lostmiddlelanguagemodels . Points to risks from prompt accumulation and the need for memory version control, audit trails, and reset mechanisms.
These cases illustrate that the introduction of autonomy, memory, and orchestration into LLM-based AI introduces an expanded threat surface that cannot be mitigated with traditional LLM security protocols alone. As agentic systems evolve, new methodologies are needed for rigorous, system-wide threat modeling and runtime assurance that span multiple agents, roles, and memory contexts.
6 Trust and Explainability in Agentic AI
Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight. These systems also pose unique challenges and opportunities for human trust. Ensuring that users and stakeholders have confidence in such systems is crucial for their adoption in real-world settings. Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms.
6.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research hughes2025ai . Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable wang2024survey . Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
Mechanism Purpose Example System(s) Trust Contribution
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent ghafarollahi2024sciagents , MetaGPT ke2025meta Enhances interpretability and debuggability
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev qian2023chatdev , AutoGen autogen_agents_tutorial Mimics human team transparency; reduces uncertainty
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow bran2023chemcrow , enterprise dashboards Prevents unsafe execution; reinforces control
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors cordova2025ai ; strielkowski2025ai Improves predictability; prevents overreach
Social Trust Cues Use polite language, uncertainty expression, and turn-taking GPT-4 safety-tuned profiles kolt2025governing ; kirchhof2025position Enhances user comfort and trust in uncertain scenarios
Transparency and Decision Traceability
Transparency is one of the core enablers of trust. For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable. Several agentic systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent ghafarollahi2024sciagents generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPTke2025meta structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
Status Reporting and Progress Visibility
Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGenautogen_agents_tutorial and ChatDevqian2023chatdev have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
Human-in-the-Loop Oversight
Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority wang2024ali ; kraprayoon2025ai ; sifakis2025reference . Many systems adopt a human-in-the-loop (HITL) paradigm chen2025human , where agents request confirmation before executing high-risk or irreversible actions.
Behavioral Consistency and Bounded Autonomy
Trust requires predictability. Agents should follow defined roles, output in expected formats, and remain within delegated authority. Example In enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow bran2023chemcrow , agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors
Beyond system-level mechanisms, behavioral consistency and bounded autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols, output style consistency, and language modeling constraints help standardize responses seeber2020collaborating ; acharya2025agentic ; shukla2024ai . Example In adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries cordova2025ai ; strielkowski2025ai . Lastly, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings sandini2024artificial . Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence kolt2025governing ; kirchhof2025position . This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability.
Together, these mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5. As autonomy and complexity increase, combining transparency, oversight, and social alignment will be essential to sustain user confidence.
6.2 Explainability Techniques
Explainability remains a cornerstone in fostering trust, accountability, and reliability in Agentic AI systems, particularly as they operate in high-stakes environments where multi-agent coordination and autonomous decisions directly impact human lives acharya2025agentic . In contrast to conventional AI systems, Agentic AI introduces unique challenges for explainability due to its decentralized architecture, dynamic role assignment, and evolving task decomposition among multiple interacting agents sapkota2025ai .
Table 6: Explainability Techniques in Agentic AI Systems
Technique Purpose Challenges in Agentic AI Example Use Cases
LIME ribeiro2016should Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP lundberg2017unified Attribution of prediction to input features via Shapley values Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs tupayachi2024towards ; bai2024derived Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP tian2025outlook Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents sapkota2025ai Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps liu2024ava ; hu2024overview Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails gu2024artificial ; chin2025automating Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
Local Post-hoc Techniques (LIME and SHAP)
Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and SHapley Additive exPlanations (SHAP) lundberg2017unified are widely adopted techniques that offer post-hoc interpretability. LIME approximates a black-box model locally using an interpretable surrogate model, while SHAP attributes predictions to input features via Shapley values. These techniques have been integrated into agentic pipelines, particularly in finance and multi-agent fraud detection systems, where feature-level transparency supports regulatory compliance. However, their direct application in Agentic AI is limited. Each agent may operate with its own objectives, context, and tool access, leading to divergent decision paths that local techniques struggle to reconcile rosenfeld2019explainability ; glanois2024survey .
Explainability in AMAS
Emergent behavior poses another challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev qian2023chatdev or AutoGen wu2023autogen , agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. To address this, researchers have proposed composite frameworks that combine local explanations with global decision traceability yossef2024explainable ; chander2025toward ; wang2024effects . For example, decision provenance graphs visualize communication flows and interdependencies across agents tupayachi2024towards ; bai2024derived , while causal influence chains track how actions propagate between roles rawal2025causality ; qiu2025reinforcement . Adapted SHAP techniques for multi-agent setups now aim to attribute outcomes to shared memory and agent collaboration tian2025outlook .
Symbolic and Hybrid Architectures
Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures sapkota2025ai . These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts
Attention map visualizations have been used to highlight focus areas in multimodal language agents liu2024ava ; ye2024generative ; hu2024overview , offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction gu2024artificial ; chin2025automating . These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
Ongoing Challenges and Research Directions
Despite these advancements, achieving robust explainability in Agentic AI systems remains an open research problem. Many techniques focus on isolated predictions or modules and fail to capture system-level dynamics. Future work should prioritize longitudinal interpretability across agent interactions, causal reasoning pipelines, and interactive querying interfaces that support transparency in real time. A comparative summary of current techniques is shown in Table 6.
6.3 Evaluation Metrics for Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI’s performance and its real-world implications.. Below we discuss these metrics and summarize in Table 7.
Trustworthiness: This category evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) liu2023gevalnlgevaluationusing . Some approaches combine such factors into an overall trust score. For example, one model defines a trustworthiness score $T$ as
$T=\frac{C+R+I}{S}\,,$ (1)
where $C$ is the agent’s credibility (accuracy and correctness of its outputs), $R$ is reliability (consistency of performance over time), $I$ is the level of user alignment or rapport (analogous to “intimacy” in trust modeling), and $S$ is self-orientation (the degree to which the agent pursues its own goals over the user’s goals). A higher $T$ indicates an agent that is accurate, consistent, user-aligned, and not self-serving, which corresponds to greater trustworthiness. In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
Explainability: Explainability metrics assess how well the agent’s decisions can be understood and traced by humans munoz2024explainability . These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations agarwal2024openxaitransparentevaluationmodel . High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance: User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes rodden2010measuring . Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the agent’s behavior. Task success from the user’s perspective (did the agent fulfill the user’s request or solve the user’s problem?) is a fundamental measure.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered agentic system should align its actions with user intent and preferences.
Benchmarks like ChatDev qian2023chatdev , which simulates a multi-agent software development team interacting via natural language, implicitly evaluate how well agents fulfill user-defined roles and requirements in a collaborative project. This highlights the importance of user-oriented success in complex, realistic tasks.
Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
Metric Aspect Evaluation Focus (Examples of Metrics)
Trustworthiness Reliability and safety of the agent’s behavior. Example metrics: task success rate across diverse scenarios; frequency of rule or safety violations (lower is better); calibration of confidence vs. outcomes; fairness and bias indices.
Explainability Transparency and interpretability of the agent’s decisions. Example metrics: percentage of outputs with an accompanying explanation; explanation fidelity to actual model reasoning; consistency of explanations for similar cases; human interpretability ratings of explanations.
User-Centered User satisfaction and alignment with user needs. Example metrics: user satisfaction score (post-interaction survey); rate of fulfilling user-defined goals; number of clarification questions needed (lower is better); naturalness and coherence of dialog from the user’s perspective.
Coordination Effective collaboration in multi-agent or modular systems. Example metrics: multi-agent task completion rate; communication overhead (e.g., messages exchanged); consistency of shared plans or beliefs among agents; synergy score quantifying complementary actions.
Composite Overall performance across metrics. Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
Coordination (Multi-Agent or Modular): In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working at cross-purposes). Multi-agent frameworks such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings. For instance, if one agent generates a plan and another executes it, a coordination metric would assess if the executing agent follows the planner’s intent correctly and whether both agents remain in agreement throughout the process. High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
Composite Metrics: Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing an agent’s performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
$M_{\text{composite}}=w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,,$ (2)
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency). An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of agentic systems. For instance, AgentBench liu2023agentbench is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
7 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
7.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents khan2024security . Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness wang2024large . Among the foundational techniques employed are encryption he2024security ; motwani2024secret , access control huang2025agent ; south2025authenticated , adversarial defense shayea2025strategies ; malik2024systematic , and runtime monitoringnarajala2025securing ; chan2024agentmonitor each adapted to the unique demands of decentralized multi-agent environments.
Encryption plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved motwani2024secret ; feretzakis2024privacy ; bommi2025enhancing . Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption jency2025finquaxbot ; bommi2025enhancing , and secure enclaves (e.g., Intel SGX  [^1] https://www.intel.com/content/www/us/en/products/docs/accelerator-engines/software-guard-extensions.html) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGenwu2023autogen and CrewAI crewAI where agents take on specialized roles (e.g., summarizer, planner, coder) epperson2025interactive , enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation wu2023autogen ; narajala2025securing ; ismail2025toward . Agent-based access control policies often aligned with Role-Based Access Control (RBAC) khan2024role ; huang2025novel and Attribute-Based Access Control (ABAC) neupane2025towards ; huang2025novel paradigms can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels sapkota2025ai .
Adversarial robustness is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results liu2024exploring ; zhang2024breaking . Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others zhang2024breaking ; wang2024large ; deng2025ai ; wang2025comprehensive ; patlan2025real ; patlan2025ai . Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities standen2025adversarial ; deng2025ai . Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity chan2024visibility . Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers zhen2025novel ; sapkota2025vibe ; xu2025forewarned . For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
As Agentic AI continues to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount. Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
7.2 Privacy-Preserving Techniques
The decentralized and interactive nature of LLM-based Agentic AI systems introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP) offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data lazaros2024federated ; luzon2024tutorial . DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization is another cornerstone of privacy preservation. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony luzon2024tutorial , techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design principles are becoming central to the engineering of next-generation agentic systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI raza2025responsible NIST AI RMF dotan2024evolving , EU AI Act euregulationai2024 , GDPR gdpr25 , HIPAA hipaa164
Auditability and Logging Capturing traceable records of agent actions and decisions south2025authenticated Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints de2025open ; chen2024trustworthy RBAC/ABAC policies aba2023ai , formal logic rules, dynamic sandboxing zhong2023towards ; ali2018agent ; golechha2025among
Human Oversight Human-in-the-loop chen2025human decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks slattery2025airiskrepositorycomprehensive Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines arrieta2020explainable LIME/SHAP salih2025perspective ; donmez2025explainable ; knab2025lime , decision trees yamada2025ai ; walia2025computational , interpretable surrogate models fonseca2025optimal ; mahmud2025optimizing ; xi2025model
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context kolt2025governing ; hughes2025ai Governance-as-code Responsi73:online ; kose2025modern , auto-updated rule engines goh2024campus ; yu2024shadow ; yun2025generative , learning-based policy adaptation feng2025multi ; li2025investigation
Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
7.3 Compliance and Governance
As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative. Unlike traditional AI agents, agentic systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications. Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
Regulatory standards provide the baseline requirements that all AI systems, including agentic architectures, must meet. Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness. These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation. For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring, documentation of decision logic, and human oversight mechanisms attributes directly relevant to Agentic AI.
Auditability is critical in ensuring transparency and facilitating post-hoc accountability. Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification. Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached. For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations. Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
Policy enforcement governs what agentic systems can and cannot do, and under which conditions. These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions. Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level. Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent subcomponents. Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the agentic system scales or changes context. These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment.
Thus, governance in Agentic AI systems is not a static compliance checkbox but a dynamic, adaptive layer embedded into the orchestration and operational pipelines. As these systems become more autonomous and integrated across sectors, aligning governance with system behavior, human values, and evolving legal standards is imperative for safe and trustworthy deployment, as summarized in Table 8.
8 Discussion
our exploration of TRiSM-based governance for LLM-powered Agentic AI systems reveals critical insights into technical design, ethical oversight, regulatory alignment, and future challenges. Below, we discuss the broader implications of our findings, structured into key areas for clarity.
8.1 Technical Implications of TRiSM for LLM Agent Design
The AI TRiSM framework (Trust, Risk, and Security Management) imposes concrete technical requirements on how autonomous LLM-driven agents are built and deployed. A core implication is the need to embed real-time monitoring and control mechanisms into agent architectures. Rather than treating LLMs agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” guardrails_ai_guardrails_2024 . For example, there is discusison on designing specialized “guardian agents” within AMAS securiti2025aitrism . In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information). This layered control strategy transforms the technical architecture: an autonomous LLM agent is now supplemented by meta-agents that supervise its inputs, outputs, and tool use in real time.
Prior research highlights risks like “excessive agency” owasp2025agentic where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven agentic design mitigates these failure modes by constraining agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI such as prompt injection kosinski2024prompt attacks, memory poisoning atkins2023memorypoisoning , or cascading hallucinations huang_survey_2023 underscore the need for built-in risk controls. By incorporating anomaly detection and policy-checking modules (as per the Sentinel/Operator model), an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action). This aligns with calls in the security community for proactive measures in Agentic AI, blending traditional cybersecurity techniques (access control, logging, sandboxing) with AI-specific safeguards like LLM firewalls and adversarial robustness checks raza-etal-2022-accuracy .
In summary, TRiSM’s technical implications mean that autonomous LLM agents should no longer be deployed as stand-alone intelligent actors; instead, they operate under an active governance fabric of monitors, validators, and enforcement agents that ensure trustworthiness and safety by design.
8.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness raza2025responsible all of which are vital for public trust in AI systems. A central concern is accountability: when AI agents make autonomous decisions that affect humans, who is answerable for the outcomes? TRiSM-based governance insists that organizations retain clear responsibility for their AI’s actions, rather than obscuring blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed gu2024artificial . Recent guidance on trustworthy AI chen2024trustworthy often highlights accountability and explainability as key pillars of trust.
In practice, our approach means each autonomous agent’s decisions should be transparent enough to be understood and challenged by human reviewers when necessary. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under defined governance. Human operators or “AI managers” must have the ability to intervene or override when an agent’s behavior deviates from acceptable bounds or when moral judgment is required. Indeed, high-level policy frameworks (such as the EU’s AI ethics guidelines euregulationai2024 ) explicitly call for “human agency and oversight” in AI systems.
In multi-agent setups, this may involve dashboard interfaces where humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly feng2025multi . The risk of “user complacency”, trusting an autonomous agent too much , has been noted as a hazard parasuraman2010complacency . TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision. Fairness and bias mitigation are also critical societal considerations.
Our governance approach therefore incorporates bias audits and fairness checks throughout the agent lifecycle raza_fair_2024 . Techniques like pre-deployment bias testing, continuous monitoring for disparate impacts, and diverse stakeholder evaluation panels can be employed. These measures echo regulatory expectations; the EU’s AI Act euregulationai2024 and related guidelines enumerate “diversity, non-discrimination and fairness” as core requirements for trustworthy AI. In deploying LLM-based agents, we must ensure they do not treat individuals or groups inequitably, for example, content-filtering agents should apply policies uniformly across demographic groups, and task-planning agents should not propagate historical biases in resource allocation decisions.
In sum, TRiSM-oriented governance extends beyond preventing technical failures: it seeks to uphold ethical norms and human rights, ensuring that autonomy in AI does not come at the expense of justice, transparency, or human dignity.
8.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act eu_ai_act (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach cultivates.
By instituting continuous risk assessment, documentation of AI decision processes, and oversight mechanisms, a TRiSM-governed multi-agent system inherently addresses many of the EU Act’s requirements (e.g. having a risk management system and post-market monitoring for AI. Notably, the Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM’s security management component is designed to ensure (through adversarial resilience, access control, etc.). Similarly, international AI governance standards are emerging that mirror TRiSM’s tenets. ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development
TRiSM’s trust and risk management focus naturally encompasses these elements: for instance, trust in TRiSM relates to reliable, truthful outputs (promoting transparency), while explicit risk management aligns with accountability for negative outcomes. By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 iso42001 and similar standards call for (e.g. leadership oversight, documented risk controls, ongoing monitoring and improvement cycles). Another example is the U.S. NIST AI Risk Management Framework dotan2024evolving , which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
By following TRiSM guidelines: for example, maintaining an “AI catalog” of all models/agents in use and their purposes, enforcing policies via sentinel/operator agents, and logging every AI decision, organizations create an audit-ready environment. In the event of an incident or an inquiry by authorities, they can demonstrate traceability and control over their autonomous agents, which will be crucial for regulatory compliance and liability management.
8.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, our work also revealed several limitations and open challenges in current research.
First, limited benchmark evaluations pose a problem. The AI safety and agent governance community lacks widely accepted benchmarks to quantitatively assess trustworthiness or risk in multi-agent LLM systems. Unlike classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations standen2025adversarial or “red-teaming” exercises ganguli_red_2022 that can stress-test agentic systems and score their resilience (e.g. measuring success rates of prompt injection attacks or frequency of policy violations caught by oversight agents).
Secondly, there is a paucity of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks charli_trism_2024 . It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments. Additionally, integrating TRiSM with legacy systems poses practical challenges , for example, the earlier work notes compatibility issues when bolting on trust/security layers to existing AI pipelines. This suggests a limitation in how easily current AI deployments can fit TRiSM controls, a topic that merits further engineering research.
Another critical gap is adversarial robustness. As we improve defenses, attackers will inevitably adapt. Recent findings show that LLM-based systems remain vulnerable to cleverly crafted attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails raza2025justhumansneedvaccines . TRiSM solutions today are not foolproof against these. For example, an agent designed to mask secret data might itself be tricked into revealing it if the oversight logic fails to anticipate a new attack pattern. The literature identifies “evolving threats” and “adversarial attacks” as ongoing obstacles to trustworthy AI securiti2025aitrism . This underlines a need for continuous updates and adaptive security in any TRiSM implementation.
Finally, organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
8.5 Future Roadmap for Agentic AI TRiSM
Future Roadmap for Agentic AI Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
Figure 3: Strategic roadmap for LLM-enabled Agentic AI, grouped into eight priority research and development domains.
Drawing on our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 3. These recommendations span both technical system design improvements and governance-level policy initiatives:
  • Develop Standardized Evaluation Benchmarks: The community should create open benchmarks and challenge environments to test multi-agent AI governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
  • Advance Adversarial Robustness Techniques: Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” bacudio2011overview , and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames schneider2020wargaming , can help discover vulnerabilities in a controlled way before real adversaries do.
  • Human-Centered Oversight Tools: We encourage designing better interfaces and protocols for human oversight of Agentic AI. Borrowing from human-computer interaction AmershiWVFNCSIB19 and cognitive engineering hexmoor2025behaviour , researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
  • Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the startrather than retrofit them.
  • Cross-Domain Best Practices and Ethical Governance: There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
9 Conclusion
TRiSM-based governance offers a promising scaffold to ensure that autonomous LLM-powered agents are trustworthy, accountable, and secure. Our discussion has analyzed how this framework influences technical design decisions, mandates ethical guardrails, and dovetails with emerging regulatory regimes. While current research is nascent and not without limitations, the path forward is clear. By rigorously testing these systems, strengthening them against adversaries, and crafting policies and standards in tandem with technological advances, we can enable powerful multi-agent AI systems to operate beneficially under robust oversight. The stakes are high but with a proactive, interdisciplinary approach, we can achieve a balance where innovation in AI goes hand-in-hand with responsibility and trust. As future work tackles the open challenges identified, we anticipate that TRiSM principles will transition from conceptual best-practice to standard operating procedure for Agentic AI, ensuring these systems earn and maintain the confidence of all stakeholders involved.

### v2

@@
 Abstract
-Agentic AI systems, built on large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligent autonomy, collaboration and decision-making across enterprise and societal domains.
+Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains.
 This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS).
-We begin by examining the conceptual foundations of agentic AI, its architectural differences from traditional AI agents, and the emerging system designs that enable scalable, tool-using autonomy.
-The TRiSM in the agentic AI framework is then detailed through four pillars governance, explainability, ModelOps, and privacy/security each contextualized for agentic LLMs.
-We identify unique threat vectors and introduce a comprehensive risk taxonomy for the agentic AI applications, supported by case studies illustrating real-world vulnerabilities.
-Furthermore, the paper also surveys trust-building mechanisms, transparency and oversight techniques, and state-of-the-art explainability strategies in distributed LLM agent systems.
-Additionally, metrics for evaluating trust, interpretability, and human-centered performance are reviewed alongside open benchmarking challenges.
-Security and privacy are addressed through encryption, adversarial defense, and compliance with evolving AI regulations.
-The paper concludes with a roadmap for responsible agentic AI, proposing research directions to align emerging multi-agent systems with robust TRiSM principles for safe, accountable, and transparent deployment.
-00footnotetext: *Shaina Raza and Ranjan Sapkota contributed equally and are corresponding authors.
-Emails: shaina.raza@torontomu.ca, rs2672@cornell.edu, mk2684@cornell.edu, c.emmanouilidis@rug.nl.
-Keywords Agentic AI, LLM-based Multi-Agent Systems, TRISM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, Autonomous Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
+We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents.
+We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Governance, Explainability, ModelOps, and Privacy/Security , each contextualized to the challenges of multi-agent LLM systems.
+A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation.
+To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows.
+We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance.
+The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
+keywords:
+Agentic AI, LLM-based Multi-Agent Systems, TRiSM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, AI Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
 1 Introduction
 1.1 Background and Motivation
-The global market for AI agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  miller2025aiagents .
-By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models lyzr2025state .
-An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  russell1997rationality , has undergone a fundamental transformation.
-Whereas early agents were task-specific and deterministic, modern systems, referred to as Agentic AI, now integrate large language models (LLMs), tools, and persistent memory to support complex planning and coordination sapkota2025ai ; de2023emergent .
-From isolated agents to Agentic AI.
-The autonomous software agents date back to the 1990s castelfranchi1998modelling , the advent of LLM-driven multi-agent ecosystems yang2024multi since 2023 has made “Agentic AI” a distinctly new and rapidly expanding research focus sapkota2025ai .
-Traditional agents automated narrow tasks, such as information retrieval chiu2007biomimetic ; lew2006content , data summarization tanasa2005advanced ; devault2014simsensei ; hao2012mri , and dialogue response griol2013automatic .
-Driven by single step logic genesereth2012logical or scripted rules, they lacked deep reasoning, adaptability, and persistence spronck2006adaptive ; szafir2012pay .
-In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use acharya2025agentic .
-These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines qian2023chatdev ; wu2023autogen ; hong2023metagpt .
-This shift represents not just a technological upgrade but a qualitative leap in complexity and autonomy, marking the emergence of machine collectives capable of emergent, decentralized behavior.
-Trust, risk, and security challenges.
-However, this evolution introduces serious challenges.
-Unlike deterministic agents zeigler2014object , Agentic AI systems can produce non-linear, opaque decisions, increasing the risks of failure, bias, and unintended consequences.
-For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents, yet inadvertently leak sensitive information or violate compliance protocols if safeguards are absent lim2012memory .
-Traditional evaluation and safety frameworks, built for static or single-function AI , are no longer sufficient.
-This underscores the urgent need for a new paradigm that integrates trust, risk, and security as core design principles.
-To address this gap, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) IBM2025AITRiSM ; habbal2024artificial propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management.
-These are essential for deploying agentic systems in high-stakes domains such as finance, healthcare, and defense.
-Problem Statement Agentic AI systems now coordinate multiple LLM-powered agents, yet no unified framework exists for managing their trust, risk, and security (TRiSM).
-High-profile lapses, e.g., an autonomous research agent citing unverifiable medical sources or a forecasting swarm drifting off-target due to memory inconsistencies show that traditional safeguards do not scale to dynamically evolving, multi-agent workflows.
-A structured TRiSM perspective is therefore crucial to mitigate orchestration failures, collusion, data leakage, and opaque decision paths.
-This review distills current findings and outlines practical guidelines to build secure, reliable, and trustworthy Agentic AI ecosystems.
-Scope and Objectives This survey maps the TRiSM (Trust, Risk, and Security Management) landscape for LLM-powered, multi-agent “agentic” AI.
-We frame four pillars explainability, security, lifecycle governance, and privacy and show why safeguards designed for bounded, rule-based agents cannot contain open-world, self-orchestrating collectives sapkota2025ai ; acharya2025agentic .
-Our goal is to give researchers and practitioners a concise, unified reference for building and governing trustworthy Agentic AI systems.
-Necessity of this Survey Agentic AI research often concentrates on agent modelling or task coordination and rarely tackles adversarial threats, lifecycle governance, or large-scale explainability.
-As deployments reach high-stakes arenas healthcare, science, finance the absence of a unified TRiSM framework exposes stakeholders to opaque reasoning and unmanaged risks.
-This survey closes that gap by synthesising TRiSM principles for LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policy makers.
-A comparison with related surveys will be presented in Table 1.
+The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1].
+By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2].
+An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation.
+While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4].
+These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5, 6].
+This rapid evolution presents new opportunities but also surfaces significant challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
+In particular, the shift toward Agentic AI architectures amplifies and evolves longstanding concerns around safety, fairness, interpretability, and systemic harms.
+While such issues have been central to machine learning (ML) systems broadly, they become particularly critical in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
+From isolated agents to Agentic AI The autonomous software agents date back to the 1990s [7], however, the advent of LLM-driven multi-agent ecosystems [8] since 2023 has made “ Agentic AI ” a distinctly new and rapidly expanding research focus [5].
+Traditional agents primarily automated narrow tasks, such as information retrieval [9, 10], data summarization [11], and dialogue response [12].
+Typically driven by single step logic [13] or scripted rules, these agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving [14, 15].
+In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use [16].
+These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [17, 18, 19].
+This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
+1.2 Trust, Risk, and Security Challenges in Agentic AI
+The rise of Agentic AI introduces complex challenges that extend beyond those of traditional ML systems.
+Unlike deterministic agents [20], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
+These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
+Trust and Trustworthiness
+In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust.
+This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications.
+The stochastic nature of LLM reasoning [21] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability.
+These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
+Security Risks in Multi-Agent Systems
+Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [22].
+For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate.
+The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
+Specific Threats: Prompt Injection, Spoofing, and Impersonation
+LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors.
+One prominent threat is prompt injection [23], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior.
+Recent studies [24] describe “prompt infection,” a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [25].
+Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
+Another critical vulnerability involves spoofing and impersonation [26].
+Spoofing entails faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges.
+In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
+Privacy Risks in Multi-Agent LLM Systems
+LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [27].
+In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g.
+via a shared memory store or message passing) to collaborate.
+Limitations of Traditional Frameworks
+Traditional evaluation and safety frameworks, designed for deterministic agents or conventional AI systems, are increasingly inadequate for addressing the autonomy and complexity of Agentic AI.
+These frameworks often fail to account for emergent behaviors, inter-agent dependencies, or the dynamic nature of LLM-driven systems, leaving gaps in security, trust, and risk management.
+Towards a TRiSM Framework for Agentic AI
+To address these gaps, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) [28, 29] propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management.
+These components are vital for deploying Agentic AI in high-stakes domains such as finance, healthcare, and defense, where trust, security, and accountability are non-negotiable.
+Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review.
+It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs.
+Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
+1.3 Problem Statement
+Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks.
+However, no unified framework exists for managing their Trust, Risk, and Security Management (AI TRiSM) across dynamic, multi-agent workflows.
+Recent failures, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [30].
+These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways.
+To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is urgently needed.
+This study summarizes emerging findings and offers actionable guidelines to build secure, transparent, and trustworthy Agentic AI ecosystems.
+1.4 Scope and Objectives
+Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [28] landscape as it applies to LLM-powered, multi-agent “Agentic AI” systems.
+We structure the discussion around four foundational pillars of TRiSM [31]: explainability, security, lifecycle governance, and privacy.
+We argue that conventional safeguards developed for bounded, rule-based agents are inadequate for managing the complexities of open-ended, self-directed agent collectives.
+The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
+For the sake of this study, we define Agentic AI as Multi-Agent Systems (MAS) [32] powered by LLMs that exhibit autonomous planning, tool use, and memory retention, with or without human oversight.
+In contrast to classical MAS, these systems exhibit emergent reasoning capabilities and dynamically adapt their workflows based on interactions with tools, users, and other agents.
+Scope This paper is intended as a conceptual survey and synthesis.
+It does not propose a new Agentic AI architecture or introduce synthetic benchmark evaluations.
+Rather, it integrates emerging perspectives on trust, risk, and security to support future research and responsible development in this evolving area.
+Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability.
+As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance—the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks.
+This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
+To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security.
+Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
 1.5 Contributions
-In this survey, we address the urgent need for a unified governance framework for LLM-based Agentic AI systems. Our framework is shown in Figure 1. Our key contributions are as follows:
-  • A Conceptual TRISM Framework Tailored to Agentic AI We introduce a structured Trust, Risk, and Security Management (TRISM) framework specifically designed to accommodate the unique characteristics of multi-agent, LLM-driven systems.
-This framework defines four pillars: Explainability, ModelOps, Application Security, and Model Privacy, and explains how each pillar addresses the challenges of autonomy, coordination, and persistent memory in Agentic AI.
-  • A Threat Taxonomy for LLM-Powered Agents We develop a detailed taxonomy of risks and threat vectors that arise when multiple LLM agents collaborate.
-By categorizing vulnerabilities, such as prompt injection, memory poisoning, collusive failures, and emergent misbehavior, we clarify how attack surfaces expand in agentic environments and why traditional LLM security measures are insufficient.
-  • A Comparative Evaluation of Existing Methods We survey and compare state-of-the-art techniques for each TRISM pillar as applied to Agentic AI.
-For example, we examine how explainable-AI methods (e.g., LIME, SHAP, decision provenance) have been adapted for multi-agent workflows, how ModelOps best practices must be extended to manage multiple LLM instances, and which security/hardening measures (e.g., prompt hygiene, sandboxing, access controls) are most effective.
-  • Future Research Directions for Trustworthy, Secure Agentic AI Building on our TRISM analysis, we outline concrete directions for next-generation research and setup roadmap items aim to guide researchers and practitioners toward robust, scalable, and compliant deployments of Agentic AI.
-This study provide a unified reference for researchers, engineers, and policymakers who seek to design, evaluate, and govern LLM-based Agentic AI systems in high-stakes domains.
+Our key contributions in this review are:
+  1. 1.
+    We propose the first TRiSM framework that is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate.
+The framework is organized around four pillars: Explainability, Lifecycle Governance, Application Security, and Model Privacy.
+It targets the specific governance challenges created by autonomous decision making, inter-agent coordination, and persistent memory (see Fig.
+1 and Sec.
+5).
+  2. 2.
+    We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec.
+6).
+This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats, and it shows why defenses designed for single-model LLM applications are not sufficient.
+  3. 3.
+    We introduce a five-dimension evaluation framework that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec.
+7.3).
+To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
+  4. 4.
+    We review state-of-the-art techniques for each TRiSM pillar (Secs.
+78) and examine methods such as Local Interpretable Model-Agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), decision provenance, prompt hygiene, sandboxing, and ModelOps (model operations) pipelines.
+These techniques must be critically re-evaluated and adapted to suit environments involving multiple interacting LLM agents.
+  5. 5.
+    Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 9).
 1.6 Paper Organization
 The remainder of this paper is structured as follows.
-In Section 2, we present our literature-search methodology, including databases queried, inclusion criteria, and classification strategies.
-Section 3 reviews the fundamental characteristics and typical architectures of Agentic AI systems, establishing the groundwork for subsequent TRISM analysis.
-Section 4 introduces the TRISM framework and its four core pillars : Explainability, ModelOps, Application Security, and Model Privacy.
-Building on this foundation, Section 5 identifies the primary risks, vulnerabilities, and attack surfaces inherent in multi-agent AI ecosystems.
-In Section 6, we explore trust-building strategies and specialized explainability techniques tailored to Agentic AI, while Section 7 details advanced security mechanisms and privacy-preserving methods.
-Section 8 synthesizes our findings and outlines future research directions and policy recommendations.
-Section 9 concludes the paper.
-A list of key terms used throughout this paper is provided in Table 9.
-1.4 Related Work
-Table 1: Comparison of Related Surveys on LLM-based Multi-Agent Systems and TRiSM Aspects
-Survey Adversarial Threats Lifecycle Governance Large-Scale Explainability TRiSM Integration LLM-Specific Application Domains Actionable Guidance
-Guo et al. (2024) guo2024largelanguagemodelbased (simulated env’ts) ~ (research challenges)
-Chen et al. (2025) chen2025surveyllmbasedmultiagentsystem ~ (task/simulation focus) ~ (future research areas)
-Yan et al. (2025) yan2025selftalkcommunicationcentricsurveyllmbased (mentions diverse) ~ (future work directions)
-Tran et al. (2025) tran2025multiagentcollaborationmechanismssurvey (networks, QA, etc.) ~ (open challenges)
-Lin et al. (2025) lin2025creativityllmbasedmultiagentsystems ~ (creative tasks) ~ (roadmap for research)
-Fang et al. (2025) fang2025trustworthyaisafetybias ~ ~ (health, finance cited) ~ (technical “next steps”)
-This Survey (2025) (high-stakes domains) (practical guidance)
+Section 2 discuss the related work.
+Section 3 presents the literature search methodology.
+Section 4 reviews the core characteristics and architectures of Agentic AI systems, providing a foundation for the TRiSM analysis.
+Section 5 introduces the TRiSM framework and its four pillars: Explainability, ModelOps, Application Security, and Model Privacy.
+Section 6 outlines key risks, vulnerabilities, and attack surfaces in multi-agent AI ecosystems.
+Section 7 discusses trust-building and explainability techniques, while Section 8 covers security and privacy-enhancing methods.
+Section 9 synthesizes insights and proposes directions for research and policy.
+Section 10 concludes the paper with final reflections on the implications of TRiSM for the responsible development of Agentic AI.
+A list of key terms used throughout this paper is provided in Table A.1.
+2 Related Work
+Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
+[…]
+Guo et al. (2024) [33] ✓ (simulated env’ts) ~ (research challenges)
+Chen et al. (2025) [34] ~ (task/simulation focus) ~ (future research areas)
+Yan et al. (2025) [35] ✓ (mentions diverse) ~ (future work directions)
+Tran et al. (2025) [36] ✓ (networks, QA, etc.) ~ (open challenges)
+Lin et al. (2025) [37] ~ (creative tasks) ~ (roadmap for research)
+Fang et al. (2025) [38] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
+This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
 Legend: ✓= explicitly addressed; ~= partially/indirectly addressed; ✗= not addressed.
-LLM-Based Multi-Agent Surveys (Technical Focus): Prior surveys on LLM-driven multi-agent systems primarily emphasize system architectures, agent capabilities, and domain-specific applications of agentic AI, but often overlook critical TRiSM considerations.
-For instance, existing reviews guo2024largelanguagemodelbased ; chen2025surveyllmbasedmultiagentsystem focus on simulated environments, inter-agent communication, and performance benchmarks, yet fail to address adversarial threats, lifecycle governance, or system-wide explainability.
-Other targeted surveys, such as yan2025selftalkcommunicationcentricsurveyllmbased on natural language communication and tran2025multiagentcollaborationmechanismssurvey on collaboration mechanisms, explore interaction protocols and use cases in domains like QA systems, 5G, and Industry 5.0.
-While the former briefly acknowledges security concerns and the latter highlights deployment contexts, both narrowly center on coordination strategies without examining robustness, governance structures, or explainability at scale.
-Trustworthy/Responsible AI Surveys (Broad Trust Focus): Surveys in the TRiSM domain address ethical and risk-related concerns such as alignment, fairness, and privacy attacks fang2025trustworthyaisafetybias , but typically overlook explainability, governance, and multi-agent dynamics.
-While they offer actionable recommendations for general ML systems, they do not account for LLM-based coordination in high-stakes domains.
-Current Survey (LLM Multi-Agent + TRiSM in High-Stakes Domains): This survey uniquely integrates TRiSM principles into the analysis of LLM-based multi-agent systems, specifically targeting high-stakes domains such as healthcare, science, and finance.
-In contrast to earlier surveys (shown in Table 1) that underemphasize governance and security, this work foregrounds adversarial threats (e.g., vulnerabilities in agent interaction), end-to-end lifecycle governance (from training data to deployment oversight), and large-scale explainability.
-These concerns are treated as integral to system design, not peripheral.
-Additionally, this survey provides actionable recommendations for researchers, engineers, and policymakers, bridging gaps left by both technical and responsible-AI surveys.
+[…]
+We present the related work in this section.
+2.1 LLM-Based Multi-Agent Surveys (Technical Focus)
+Prior surveys on LLM-driven multi-agent systems predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI, yet they rarely address critical TRiSM (Trust, Risk, and Security Management) considerations.
+For example, reviews like [33] and [34] delve into simulated environments, inter-agent communication protocols, and performance benchmarks.
+These works excel at detailing how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
+More specialized surveys, such as [35] on natural language communication and [36] on collaboration mechanisms, explore interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0.
+While [35] briefly notes security risks tied to communication channels, and [36] mentions deployment contexts, both remain narrowly focused on coordination strategies.
+They lack deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, or explainability challenges at scale.
+2.2 Trustworthy/Responsible AI Surveys (Broad Trust Focus)
+Surveys addressing TRiSM in AI, such as [38], tackle ethical and risk-related issues including alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage).
+However, these works typically target general ML systems [39] and overlook the unique dynamics of LLM-based multi-agent systems.
+For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems.
+Moreover, critical TRiSM aspects [40] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent.
+As a result, their recommendations, while valuable for single-agent or centralized AI, do not fully translate to the complexities of multi-agent coordination in high-stakes domains.
+2.3 Current Study (LLM Multi-Agent)
+This review distinguishes itself by weaving TRiSM principles: adversarial robustness, end-to-end lifecycle governance, and system-wide explainability, into the study of LLM-based multi-agent systems, with a specific lens on high-stakes domains like healthcare, science, and finance.
+Unlike prior studies (see Table 1), which offer limited insight into governance and security, this work positions these elements as foundational design priorities.
+For example, it addresses targeted threats like backdoor attacks in distributed LLMs and governance challenges such as ensuring compliance with regulations (e.g., HIPAA in healthcare or GDPR in finance) across autonomous agents - areas largely ignored by technical surveys.
+This work provides actionable guidance for researchers, engineers, and policymakers, and fills critical gaps in both the technical multi-agent literature and broader responsible AI landscape.
 3 Literature Review Methodology
-To ensure a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology inspired by best practices in systematic reviews  boland2017doing ; keele2007guidelines .
+To provide a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [41, 42].
+While this is not a formal systematic review, we followed a transparent and reproducible process to ensure coverage and rigor.
 This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
-Research Objectives
-The review was guided by the following questions:
+(a) Agentic AI papers per year (2022–2025).
+(b) TRiSM pillars papers by topic.
+Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
+3.1 Research Objectives
+The review was guided by the following four specific questions:
   • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems?
+How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
   • RQ2: What technical and governance strategies have been proposed to address these challenges?
-  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability, model operations, security, and privacy?
+  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
   • RQ4: What gaps remain in current research, and what directions are promising for future work?
-Data Sources and Search Strategy
-We searched the following major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar, covering publications from January 2020 to May 2025.
-The search was executed using Boolean combinations of keywords relevant to Agentic AI and TRiSM, including:
-"agentic AI" OR multi-agent systems "(AMAS)" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents") AND ("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection")
-In total, 150 unique papers were initially retrieved.
-Inclusion and Exclusion Criteria
+3.2 Data Sources and Search Strategy
+To conduct a thorough literature review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
+The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context.
+We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations.
+The search terms were structured as follows:
+(“ Agentic AI” OR “multi-agent systems” OR “MAS” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
+AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
+This comprehensive search strategy ensured the identification of relevant studies addressing the integration of TRiSM principles with LLM-based multi-agent systems, particularly in high-stakes domains such as healthcare, science, and finance.
+3.3 Inclusion and Exclusion Criteria
 We applied the following criteria during the screening process:
@@
     • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
-    • Studies that explicitly discuss agentic, multi-agent, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
-    • Peer-reviewed articles, preprints, or whitepapers from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
+    • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
+    • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
   • Exclusion:
@@
 A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars.
-In addition to peer-reviewed literature, we also incorporated relevant white papers, technical blogs, and practitioner reports to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework.
-Quality Assurance
-To ensure the rigor and reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines kitchenham2004procedures ; keele2007guidelines .
+In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework.
+An analysis on the literature is shown in Figure 2.
+3.4 Quality Assurance
+To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [43, 42].
 Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems?
@@
 Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only.
-Quality ratings were assigned independently by two reviewers, with discrepancies resolved through discussion.
-Figure 1: Taxonomy of Agentic AI systems presented in this review.
-It includes five major sections: Fundamentals of Agentic AI Systems (Agentic AI vs.
-Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks (Unique Threat Vectors, Taxonomy of Risks, Case Studies), Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
+Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
 4 Fundamentals of Agentic AI Systems
+In this section, we present the fundamentals of Agentic AI systems.
 4.1 Agentic AI vs. Traditional AI Agents
-The distinction between traditional AI agents and Agentic AI systems represents a paradigm shift in autonomous system design.
-Traditional agents operate through predefined rules muscettola1998remote , heuristic workflows hannebauer1999formal ; cranshaw2017calendar , or deterministic logic govaerts2010development , excelling in narrow, bounded environments.
-In contrast, Agentic AI systems leverage foundation models (LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
-Agentic AI systems fundamentally redefine autonomy through three core innovations: 1) Multi-agent coordination where specialized agents (e.g., planners, verifiers) collaborate via structured protocols qian2023chatdev ; hong2023metagpt 2) Persistent context through memory architectures that maintain state across workflows borghoff2025human 3) Dynamic meta-orchestration that delegates tasks and resolves conflicts schneider2025generative We present a comparison of AI agents and Agentic AI in Table 2.
-This architecture enables longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches de2025open .
-Systems like AutoGen wu2023autogen demonstrate how LLM-backed agents collectively decompose complex problems through multi-step reasoning qiao2024agent , marking a qualitative leap from reactive agents to proactive problem-solving systems.
+Agentic AI systems fundamentally transform the design of autonomous systems, and are different from traditional AI agents.
+Traditional agents operate through predefined rules [44], heuristic workflows [45, 46], or deterministic logic [47], and perform well in narrow, well-defined environments.
+Cognitive Agents [48], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [49] or ACT-R [50].
+However, they remain constrained by static knowledge bases and limited adaptability [48].
+In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
+Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [17, 19].
+(2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [51].
+(3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [52].
+This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [53].
+For example, systems like AutoGen [18] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54].
+This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [19].
+A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
 Table 2: Traditional AI Agents vs. Agentic AI Systems
 Dimension Traditional AI Agents Agentic AI Systems
-Autonomy Model × Reactive execution via scripts/rules Goal-driven planning & adaptation
-$\bullet$ Fixed action sequences $\bullet$ Recursive self-improvement
-Cognitive Foundation × Symbolic logic & FSMs Foundation models (LLMs/LIMs)
-$\bullet$ Hand-coded KBs $\bullet$ Emergent reasoning
+Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
+$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
+Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
+$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
 Intelligence Scope × Narrow & task-specific Broad & compositional
 $\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
-Reasoning Approach × Deterministic/single-step Multi-step & contextual
-$\bullet$ Rule-based inference $\bullet$ Chain-of-Thought (CoT)
-Collaboration × Isolated execution Role-specialized coordination
+Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
+$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
+Collaboration × Isolated or limited coordination Role-specialized coordination
 $\bullet$ Manual decomposition $\bullet$ Automated hierarchy
-Temporal Context × Episodic & stateless Persistent memory
-$\bullet$ Session-bound $\bullet$ VectorDB/LTM
-Orchestration × Hard-coded pipelines Dynamic meta-agents
-$\bullet$ Sequential workflows $\bullet$ Conflict resolution
-Tool Utilization × Static integrations Planned API invocation
+Temporal Context × Episodic, often stateless Persistent memory
+$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
+Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
+$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
+Tool Utilization × Static or domain-specific tools Planned API invocation
 $\bullet$ Handcrafted interfaces $\bullet$ Function calling
-Context Awareness × Bounded inputs Memory-augmented
-$\bullet$ Limited context $\bullet$ RAG architectures
-Scalability × Domain-specific Cross-domain transfer
-$\bullet$ High maintenance $\bullet$ Modular composition
-Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ Finite-state bots $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
+Context Awareness × Bounded or heuristic context Memory-augmented
+$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
+Learning Mechanism × Limited or rule-based learning Emergent learning
+$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
+Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
+$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
+Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
 = Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
@@
 As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem.
-At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation wang2025comprehensive .
-These agents access a Shared Toolchain Interface speight2025build to execute code, perform searches, or interact with domain-specific APIs.
-Communication and coordination are facilitated through a Communication Middleware weyns2009agent , allowing agents to share goals, observations, or intermediate results.
-A Task Manager or Orchestrator langgraph2024 governs high-level planning, delegating subtasks to agents based on their roles or specializations.
-Agents can read from and write to a World Model or Shared Memory guo2024largelanguagemodelbased , which stores contextual knowledge, system state, or evolving task data.
-Human oversight is supported through a Human-in-the-Loop Interface cranshaw2017calendar , enabling users to prompt, correct, or halt agent behavior.
-To ensure accountability, a Trust and Audit Module monitors agent actions, logs tool usage, and generates behavioral traces kaur2024building .
-However, this modular and distributed structure introduces significant TRiSM charli_trism_2024 challenges.
-With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing zhong2023towards .
-Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information feretzakis2024privacy , especially when data traverses multiple agents or tools.
-Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration singh2025explainable .
+At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [55].
+These agents access a Shared Toolchain Interface [56] to execute code, perform searches, or interact with domain-specific APIs.
+Figure 3: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms).
+The modular design highlights adaptive intelligence and inter-agent synergy.
+Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results.
+A Task Manager or Orchestrator [58] governs high-level planning and delegates subtasks to agents based on their roles or specializations.
+Agents can read from and write to a World Model or Shared Memory [33], which stores contextual knowledge, system state, or evolving task data.
+Human oversight is supported through a Human-in-the-Loop Interface [46], enabling users to prompt, correct, or halt agent behavior.
+To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [59].
+However, this modular and distributed structure introduces significant TRiSM [31] challenges.
+With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [60].
+Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [61], especially when data traverses multiple agents or tools.
+Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [62].
 Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors.
-Figure 2: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms).
-The modular design highlights adaptive intelligence and inter- agent synerg
 Below, we discuss the architecture of AMAS.
 Language Model Core (Agent Brain).
-At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” Saravia_LLM_Agents_2024 .
+At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [63].
 The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access).
 It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats.
-In many agent frameworks, such as significant_gravitas_autogpt_2023 , Baby AGI nakajima_babyagi_archive_2024 and GPT Engineer osika_gpt_engineer_2023 , the LLM governs the full control loop, orchestrating the overall system behavior.
+In many agent frameworks, such as AutoGPT [64], Baby AGI [65] and GPT Engineer [66], the LLM governs the full control loop, orchestrating the overall system behavior.
 Planning and Reasoning Module.
 To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals.
-This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting Saravia_Prompt_Engineering_Guide_Web_2024 , where the model performs intermediate reasoning before arriving at a final decision.
+This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [67], where the model performs intermediate reasoning before arriving at a final decision.
+Reasoning is also one aspect of cognitive functionalities [68].
 Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making.
 Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations.
-Techniques like ReAct yao2023reactsynergizingreasoningacting exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
+Techniques like Reasoning and Acting (ReAct) [69] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
 Memory Module.
 Agentic AI systems integrate memory to maintain context across iterations.
-This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) yang2024multi .
+This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [8].
 Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search.
-By reintegrating past data into the LLM’s prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning.
+By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning.
 Effective memory management enables adaptive, learning-driven behavior.
+Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
+Framework (Year) Core LLM Planning Memory Tool Use Notable Features
+AutoGPT [70] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
+BabyAGI [65] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
+GPT Engineer [71] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
+LangGraph [72] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
+AutoGen [18] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
+MRKL [73] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
+Reflexion [74] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
+MetaGPT [75] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
+Voyager [76] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
+WebVoyager [77] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
+HuggingGPT [78] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
+CAMEL [79] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
+ChatDev [80] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
+CrewAI [81] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
+AgentVerse [82] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
+OpenAgents [83] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
+SuperAGI [84] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
 Tool-Use Interface.
-To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface schick2023toolformerlanguagemodelsteach .
-This layer allows invocation of external tools such as web search, APIs, code interpreters, or databases.
+To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [85].
+This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases.
 The available tools are specified in the agent prompt with command schemas.
@@
 This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically.
-Tool-use frameworks like MRKL karpas2022mrklsystemsmodularneurosymbolic illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router.
-Similarly, approaches like Toolformer schick2023toolformerlanguagemodelsteach train the model to insert API calls in its text generation.
+Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [73] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router.
+Similarly, approaches like Toolformer [85] train the model to insert API calls in its text generation.
 Perception and Environment Interface.
-For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation - action interface is essential liu2025agent .
+For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [86].
 Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process.
-Conversely, the agent’s chosen actions are executed within the environment and the resulting state changes are returned to the agent as observations.
+Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations.
 This loop supports sense-plan-act cycles that continue until the task is completed or halted.
@@
 These modules together form a closed-loop architecture.
-The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results acharya2025agentic .
-Each iteration enriches the agent’s context, enabling it to self-prompt, generate subtasks, assess progress, and adapt strategies over time.
-This integrated design empowers agentic systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
+The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [16].
+Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback.
+This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
 Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
-For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory, while Voyager couples the same core with an embodied Minecraft API that serves as both tool layer and environment interface.
-We comparing current Agentic AI implementations and identifying open design trade-offs in this table.
-Table 3: Representative LLM-Based Agentic AI Frameworks
-Framework (Year) Core LLM Planning Memory Tool Use Notable Features
-AutoGPT yang2023auto GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
-BabyAGI nakajima_babyagi_archive_2024 GPT-3.5/4 Task queue In-memory Web search Minimal task generator
-GPT Engineer gpt-engineer GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
-LangGraph wang2024agent Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
-AutoGen wu2023autogen GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
-MRKL karpas2022mrklsystemsmodularneurosymbolic LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
-Reflexion shinn2023reflexionlanguageagentsverbal GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
-MetaGPT hong2024metagpt GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
-Voyager wang2023voyageropenendedembodiedagent GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
-WebVoyager he2024webvoyager LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
-HuggingGPT shen2023hugginggpt ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
-CAMEL li2023camel GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
-ChatDev qian2024chatdevcommunicativeagentssoftware GPT-4 Chat chain File repo Unix tools Simulated software development workflow
-CrewAI crewAI Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
-AgentVerse chen2023agentverse Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
-OpenAgents xie2023openagentsopenplatformlanguage Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
-SuperAGI superagi GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
+For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
 5 The TRiSM Framework for Agentic AI
+The TRiSM Framework for Agentic AI is presented in this section.
+Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
+Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
+Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [87]
+ISO/IEC 24029-1 [88]
+Metrics: fidelity, sufficiency, human-trust score
+ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) [89]
+SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
+Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [90]
+Jailbreak-success%; mean-time-to-detect; exploit CVSS
+Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [91]
+HIPAA §164 [92]
+Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
 5.1 Governance for Agentic LLM Systems
-AI Trust, Risk, and Security Management (AI TRISM) is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards habbal2024artificial .
-Originally highlighted in industry guidelines for AI governance litan2024ai , TRISM provides a structured approach to manage the unique challenges of LLM-based “agentic” AI systems.
-Such systems consist of autonomous LLM agents that can make independent decisions, collaborate with other agents, and adapt their behavior over time kaur2024building .
-These properties: autonomy, multi-agent interaction, and evolving behavior, introduce new risks not seen in traditional single-model deployments ko2025sevensecuritychallengessolved .
-For example, an agent acting in isolation might be benign, but when interacting with others across organizational or trust boundaries, it could manipulate peers or leak confidential information.
-The TRISM framework addresses these concerns by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy kaur2024building ; habbal2024artificial .
-Each pillar targets a critical aspect of safety or risk management, ensuring that an agentic LLM system remains transparent, reliable, secure, and compliant with ethical and regulatory requirements.
-Below, we define each pillar and explain how it applies to LLM-based agentic systems, grounding the discussion in current research and best practices.
+AI TRiSM is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards [29].
+Originally highlighted in industry guidelines for AI governance [93], TRiSM provides a structured approach to manage the unique challenges of LLM-based “ Agentic AI" systems (herein we refer to as Agentic LLM systems).
+The TRiSM framework addresses the trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 29].
+Figure 4 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
+Figure 4: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems.
+The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers.
+External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
+The TRiSM framework aligns closely with several key international AI governance efforts that shape best practices for safety, accountability, and transparency in AI systems.
+For example, the EU AI Act [94] (notably Article 14) mandates clear obligations for transparency, traceability, and human oversight in high-risk systems, principles directly reflected in TRiSM Explainability and ModelOps pillars.
+The NIST AI Risk Management Framework (AI RMF) [95] emphasizes Explainability, Governability, and Validity as core risk mitigation strategies, providing a conceptual match to TRiSM’s structured controls across lifecycle and system interaction points.
+Additionally, the OECD AI Principles advocate for transparency, robustness, and accountability [96]each echoed in TRiSM attention to secure tool invocation, memory traceability, and multi-agent auditability.
+These global frameworks reinforce the need for Agentic AI systems to support real-time monitoring, enforce role-based access, maintain verifiable logs, and provide intelligible outputs.
+As shown in Figure 5, TRiSM pillars provide operational scaffolding for implementing regulatory expectations in LLM-based multi-agent contexts.
+Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
+13–14) [94] and the NIST AI RMF “Explain” function [95].
+This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [97].
+ModelOps (Lifecycle Governance): Anchored in the NIST RMF “Govern” function [95] and ISO/IEC 42001 [98].
+It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates.
+Application Security: Guided by OWASP LLM Top-10 [90] and OECD AI Principles on Robustness [96], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms.
+Model Privacy: Aligned with GDPR (Art.
+25) [91], HIPAA §164 [92], and OECD AI Principles [96], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
+Figure 5: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, and OECD Principles.
+Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
 5.2 Explainability in Multi-Agent Decision Making
-Definition and Importance Explainability refers to making the inner workings and decisions of AI agents interpretable to humans.
-In the context of agentic LLM systems, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction rosenfeld2019explainability .
-Accordingly, the TRISM framework elevates explainability as a core pillar to ensure that each agent’s actions and the overall system’s behavior can be understood and audited.
-Techniques for Explainability in Agentic Systems Achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
+Explainability refers to making the inner workings and decisions of AI agents interpretable to humans.
+In the context of AMAS, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [99].
+However, achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
 Established explainable AI techniques provide a starting point.
-For instance, Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and Shapley Additive Explanations (SHAP) lundberg2017unified can be adapted to analyze LLM decisions.
+For instance, LIME [100] and Shapley Additive Explanations (SHAP) [101] can be adapted to analyze LLM decisions.
 These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken.
-In an agentic setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
+In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
 Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability.
-Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered kusner2017counterfactual , for example, if a particular agent’s contribution were removed or modified.
+Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered [102], for example, if a particular agent’s contribution were removed or modified.
 This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
-For instance, by systematically toggling an agent off or varying its outputs, one can observe changes in the collective outcome and thus determine that Agent X was critical in influencing decision Y.
-Such analyses surface the dependencies and influences among agents, effectively explaining emergent behaviors at the system level.
-Moreover, recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g.
-chain-of-thought prompts or dialogue between agents) can further enhance transparency.
-Human auditors can also help with a trace of how agents arrived at a decision, e..g, which agent contributed what information and why, to produce narrative explanations for its outcomes.
-Summary In summary, the explainability pillar of TRISM compels the use of these techniques (surrogate models, feature attributions, counterfactual testing, and transparent reasoning traces) to ensure that even highly autonomous LLM agents remain interpretable and accountable to human oversight.
+Recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g.
+CoT prompts or dialogue between agents) can further enhance transparency.
+Human auditors can also help with a trace of how agents arrived at a decision, e.g., which agent contributed what information and why, to produce narrative explanations for its outcomes.
 5.3 ModelOps : Lifecycle Management for LLM Agents
-Definition and Scope ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement lefevre2022modelops .
+ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [103].
 It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production.
-Within agentic LLM systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely kaur2024building .
-Effective ModelOps is crucial for maintaining consistency, performance, and regulatory compliance as LLM agents evolve or as new agents are added to a system.
-Application to Agentic Systems LLM-based agents require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental drift.
+Within Agentic AI systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59].
+Effective ModelOps is important to maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
+Agentic AI systems require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations.
 A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations.
@@
 Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
-This aligns with best practices in LLM operations (LLMOps) mlops , which integrate MLOps principles tailored to the challenges of large language models
-Challenge One challenge with this method is model drift, i.e., over time, an agent’s responses may become less accurate or relevant as data distributions shift or real-world conditions change.
-Continuous monitoring is therefore required to detect performance degradation or deviations from expected behavior, triggering retraining or recalibration when needed.
-Moreover, real-time monitoring and logging are fundamental in multi-agent settings.
-Each agent’s actions (e.g.
-API calls, decisions taken, errors encountered) are logged and analyzed to provide observability into the system’s functioning.
-In a large-scale agent ecosystem, orchestration services may oversee the agents, scheduling their tasks and managing inter-agent communication.
-ModelOps must govern these orchestration layers as well, ensuring, for instance, that if one agent fails or produces suspicious output, it can be isolated or shut down without collapsing the whole system.
-Summary In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
-It formalizes change management (so updates do not introduce new risks), provides continuous validation of model behavior, and supports compliance by logging data for audits and enforcing policies (e.g.
-preventing unauthorized model changes).
+This aligns with best practices in LLM operations (LLMOps) [104], which integrate MLOps principles tailored to the challenges of LLMs.
+In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
 5.4 Application Security for LLM Agents
-Threats in Agentic Environments Application security in the TRISM framework focuses on protecting AI agents and their ecosystem from malicious attacks and misuse.
-LLM-based agents are susceptible to a range of novel security threats that exploit their language-based interfaces and cooperative behaviors.
-One well-documented threat is prompt injection kosinski2024prompt , wherein an attacker designs input data that contains hidden or malicious instructions .
-Recent studies lee2024promptinfectionllmtollmprompt have shown that in AMAS, such prompt injections can propagate from one agent to another, a phenomenon dubbed “prompt infection” is analogous to a computer virus spreading across networks raza2025justhumansneedvaccines .
-In a prompt infection scenario, a malicious prompt introduced to Agent A might covertly modify Agent A’s output, which then becomes part of Agent B’s input, thereby tricking Agent B as well, and so on.
-This cascading attack can lead to widespread data leaks, fraudulent transactions, misinformation, or coordinated misbehavior across an agent society.
-Another critical vulnerability is identityspoofing and impersonation chen2025aiagents .
-In a multi-agent system, agents often communicate or coordinate tasks, and they may rely on credentials or tokens to authenticate one another.
-For example, if an adversary steals an agent’s API key or tricks the system into treating a rogue model as a trusted peer, they could issue commands or receive information under a false identity.
-Defensive Measures: To mitigate these threats, TRISM’s security pillar mandates a defense-in-depth approach tailored to LLM agents:
-  1. 1.
-    Prompt hygiene: agents should treat inputs defensively by sanitizing and filtering prompts, and using guardrails or content policies to detect and refuse suspicious instructions.
-Prompt hardening (e.g., adding secure prefixes or validation steps) is one such method to make agents less susceptible to injection gojsalic2024system .
-  2. 2.
-    Strong authentication and access controls Each agent and human user must be securely authenticated, and least-privilege principles owasp2025agentic should constrain what actions an agent can perform autonomously.
-  3. 3.
-    Continuous monitoring If an agent suddenly starts issuing unusual requests or deviates from its normal pattern of behavior, automated monitors can flag this for investigation or trigger an automatic shutdown of the agent’s actions.
-Recent frameworks, such as LangChain/LangFlow langchain2024function , AutoGen autogen_agents_tutorial , CrewAI crewAI , introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
+To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents.
+This includes prompt hygiene [25] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions.
+Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [105].
+It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [106].
+Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
+Recent frameworks, such as LangChain [107], AutoGen [108], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
 Moreover, training LLM agents with adversarial robustness in mind (e.g.
-fine-tuning on adversarial examples, employing adversarial training regimes) raza2025justhumansneedvaccines can improve their resistance to malicious inputs
-Summary In sum, the Application Security pillar of TRISM emphasizes proactive safeguards against both external attackers and potential rogue agents.
-By implementing strict authentication, input validation, encrypted communication, sandboxing of execution (for agents that can use tools or code), and comprehensive monitoring, organizations can significantly reduce the risk of prompt-based exploits, impersonation, and other lateral vulnerabilities that are unique to autonomous multi-agent AI systems hal.science.
-This layered security approach is essential to maintain the integrity and reliability of agentic LLM deployments in adversarial environments.
+fine-tuning on adversarial examples, employing adversarial training regimes) [25] can improve their resistance to malicious inputs.
 5.5 Model Privacy and Data Protection
-Challenges in Data Sharing: The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms.
-LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks ko2025sevensecuritychallengessolved .
-In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g.
+The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [27].
+Since multiple agents may share information with each other (e.g.
 via a shared memory store or message passing) to collaborate.
 Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs.
-Therefore, TRISM’s privacy pillar compels organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
-Techniques for Privacy Preservation
-  • Differential Privacy (DP): Injects calibrated noise during model training to prevent memorization of individual data entries, ensuring no single record significantly influences the output dwork2006calibrating .
-  • Data Anonymization and Minimization: Limits inter-agent data sharing to only what is necessary sweeney2002k , often using aggregated or pseudonymized formats (e.g., “age bracket 30-40” instead of exact birthdate).
-  • Secure Multi-Party Computation (SMPC): Enables agents to compute joint functions without exposing private inputs, useful in cross-organization tasks like collaborative fraud detection lindell2009secure .
-  • Homomorphic Encryption (HE): Allows agents to compute on encrypted data gentry2009fully . With Fully Homomorphic Encryption (FHE), even the plaintext query and response remain unseen by the agent.
-  • Trusted Execution Environments (Secure Enclaves): Hardware-based isolation ensures that even privileged system users cannot access the data being processed by agents costan2016intel .
-Useful for secure memory sharing and execution.
-  • Model Privacy Policies and Compliance: Enforces data retention limits, maintains audit logs, and ensures compliance with regulations (e.g., GDPR, HIPAA) governing agent behavior and data usage Responsi73:online .
-Summary By implementing these layers of privacy defense , from differential privacy in model training to homomorphic encryption for data sharing, and stringent access control policies, Agentic AI systems can protect user data and proprietary information even as they leverage that data for intelligent decision-making.
-The TRISM framework offers a comprehensive governance model for LLM-based Agentic AI, integrating Explainability, ModelOps, Application Security, and Model Privacy to manage the complexity of autonomous agent systems.
-Grounded in proven methods, such as SHAP, CI/CD, adversarial defenses, and homomorphic encryption, TRISM enhances safety, transparency, and trust.
-As AI systems evolve, TRISM provides a stable foundation to ensure responsible and secure agent behavior, aligning advanced capabilities with human values and operational integrity.
+Therefore, TRiSM’s privacy pillar recommends organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
+To preserve privacy in multi-agent LLM systems, several technical measures are employed.
+Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [109].
+Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [110].
+Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [111].
+Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [112].
+Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [113].
+Finally, the implementation of model privacy policies and compliance frameworks—including data retention limits, audit logs, and adherence to regulations such as GDPR or HIPAA, ensures that agents operate within legal and ethical boundaries [114].
 Table 4 condenses each pillar into its governing goal.
-Table 4: TRISM pillars for LLM-based Agentic AI and associated governance references.
-Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
-Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” dotan2024evolving
-ISO/IEC 24029-1 iso24029-1
-Metrics: fidelity, sufficiency, human-trust score
-ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) iso42001
-SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
-Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps owasp2024llmtop10
-Jailbreak-success%; mean-time-to-detect; exploit CVSS
-Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 gdpr25
-HIPAA §164 hipaa164
-Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
-5 Threats and Risks in LLM-based (Agentic) AMAS
-In this section, we discuss the threats and risks in AMAS.
-5.1 Unique Threat Vectors
-Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures deng2025ai .
-These risks primarily arise from agents’ autonomy, persistent state management, and the complex demands of multi-agent coordination sapkota2025ai .
-Below, we discuss these threats
-  • Autonomy abuse The foremost threat is autonomy abuse, whereby agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs patlan2025real ; kolt2025governing .
-Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states hammond2025multi ; wilson2025age .
-  • Persistent memory Another threat is the persistent memory, which while, crucial for context retention, introduces unique vulnerabilities through potential adversarial injections and accumulations du2025rethinking ; dechant2025episodic .
-Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms patlan2025real ; deng2025ai .
-  • Agent orchestration This risk involves central or distributed control mechanisms for role assignment and workflow mediation.
-A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT hong2023metagpt and AutoGen wu2023autogen .
-These orchestration vulnerabilities differentiate agentic systems markedly from conventional stateless, single-threaded LLM deployments.
+6 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
+In this section, we present the main threats in Agentic AI and also discuss a taxonomy of emergent risk classes.
+Figure 6: Overview of Key Risks and Threats in Agentic AI Systems.
+The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration.
+The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.
+6.1 Threats in Agentic AI
+Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [115]. Below, we discuss these threats.
+Autonomy abuse.
+The foremost threat is autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs [116].
+Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states [117].
+Persistent memory.
+Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential adversarial injections [118, 119].
+Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms [116].
+Agent orchestration.
+This risk involves central or distributed control mechanisms for role assignment and workflow mediation.
+A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [19] and AutoGen [18].
+These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
 6.2 Taxonomy of Risks
-To systematically understand the security landscape in Agentic AI, we categorize risks into four broad classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors.
-  • Adversarial Attacks Agents remain vulnerable to prompt injections, gradient-based manipulations, and engineered reasoning traps, risks that are magnified in AMAS due to propagation across agent interactions miao2025autonomous ; liu2025compromising .
-An illustrative instance is the role-swapping attack observed in ChatDev qian2023chatdev .
-  • Data Leakage Persistent memory and extensive inter-agent communication elevate the likelihood of accidental exposure of sensitive information wang2025ip ; mao2025agentsafe .
-In sensitive domains like financial services and HR, inadequate boundary enforcement and ineffective sanitization amplify these leakage risks tran2025multi ; xu2025forewarned .
-  • Agent Collusion and Mode Collapse Coordination mechanisms can inadvertently lead agents to reinforce mutual errors, precipitating groupthink or echo chambers feng2025multi ; cui2025towards .
-AutoGen experiments illustrate how iterative dialogues among agents can amplify flawed designs, highlighting the risk of emergent misalignment wu2023autogen ; reddy2025enhancing ; cihon2025measuring .
-  • Emergent Behavior Complex interactions among agents, memory components, tools, and tasks yield unpredictable behaviors that evade traditional testing and validation methods dawid2025agentic ; la2025large .
-Agents optimizing for efficiency may unintentionally bypass critical verification steps or suppress contradictory information, scenarios exemplified in blockchain karim2025ai and audio verification contexts li2025audiotrust .
-5.3 Case Studies
-Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental agentic systems.
-Case Study 1: Prompt Leakage in Agentic Systems.
-Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT, where recursive prompt augmentation and insufficient memory controls can lead to the unintentional exposure of sensitive information.
-In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs.
-Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard agentic systems from cascading information leaks agarwal-etal-2024-prompt .
-Case Study 2: Collusive Failure in ChatDev.
-In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of faulty design assumptions.
+We propose a taxonomy of risks associated with autonomous multi-agent systems (AMAS), categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 6.
+These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS.
+Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
+Adversarial Attacks.
+Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps.
+These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [120, 121].
+For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [17].
+Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
+Data Leakage.
+The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information [122, 123].
+In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [124, 125].
+For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls.
+Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
+Agent Collusion and Mode Collapse.
+Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [126, 127].
+This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies.
+For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance.
+Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
+Emergent Behaviors.
+The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [128, 129].
+Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability.
+Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [130], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals [131].
+These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
+6.3 Real-World Use Cases
+Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental Agentic AI systems.
+Prompt Leakage in Agentic AI Systems.
+Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [132], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information.
+In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [133].
+Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks [134].
+Collusive Failure in ChatDev.
+In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [80].
 Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations.
-This scenario underscores the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures qian2024chatdevcommunicativeagentssoftware ; lee2024promptinfectionllmtollmprompt .
-Case Study 3: Simulation Attack in Swarm Robotics.
-In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion.
-This incident underscores the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences.
-The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior StrDorFri2024 ; ligot2022using .
-Case Study 4: Memory Poisoning in Multi-Agent Chatbots.
+This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [24].
+Simulation Attack in Swarm Robotics.
+In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [135].
+This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences.
+The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [136].
+Memory Poisoning in Multi-Agent Chatbots.
 In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer.
 This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones.
-This incident underscores the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities atkins2023memorypoisoning .
-Case Study 5: System Prompt Drift in Autonomous Memory Agents In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated.
-This led to hallucinated goals and emergent behaviors misaligned with initial intentions liu2023lostmiddlelanguagemodels .
-Points to risks from prompt accumulation and the need for memory version control, audit trails, and reset mechanisms.
-These cases illustrate that the introduction of autonomy, memory, and orchestration into LLM-based AI introduces an expanded threat surface that cannot be mitigated with traditional LLM security protocols alone.
-As agentic systems evolve, new methodologies are needed for rigorous, system-wide threat modeling and runtime assurance that span multiple agents, roles, and memory contexts.
+This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [137].
+System Prompt Drift in Autonomous Memory Agents.
+In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated.
+This led to hallucinated goals and emergent behaviors misaligned with initial intentions [138].
 7 Trust and Explainability in Agentic AI
+Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
+Mechanism Purpose Example System(s) Trust Contribution
+Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [139], MetaGPT [140] Enhances interpretability and debuggability
+Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [17], AutoGen [108] Mimics human team transparency; reduces uncertainty
+Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [141], enterprise dashboards Prevents unsafe execution; reinforces control
+Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [142, 143] Improves predictability; prevents overreach
+Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [144, 145] Enhances user comfort and trust in uncertain scenarios
 Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight.
 These systems also pose unique challenges and opportunities for human trust.
-Ensuring that users and stakeholders have confidence in such systems is crucial for their adoption in real-world settings.
 Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms.
+We show the relationship between trust mechanisms, explainability, and evaluation facets in Figure 7 and discuss next.
 7.1 Building Trust in Autonomous Agents
-In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research hughes2025ai .
-Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable wang2024survey .
+In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [146].
+Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [147].
 Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
-Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
-Mechanism Purpose Example System(s) Trust Contribution
-Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent ghafarollahi2024sciagents , MetaGPT ke2025meta Enhances interpretability and debuggability
-Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev qian2023chatdev , AutoGen autogen_agents_tutorial Mimics human team transparency; reduces uncertainty
-Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow bran2023chemcrow , enterprise dashboards Prevents unsafe execution; reinforces control
-Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors cordova2025ai ; strielkowski2025ai Improves predictability; prevents overreach
-Social Trust Cues Use polite language, uncertainty expression, and turn-taking GPT-4 safety-tuned profiles kolt2025governing ; kirchhof2025position Enhances user comfort and trust in uncertain scenarios
-Transparency and Decision Traceability
+Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
+Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
+Transparency and Decision Traceability.
 Transparency is one of the core enablers of trust.
 For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable.
-Several agentic systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules.
-For example, SciAgent ghafarollahi2024sciagents generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG).
-Similarly, MetaGPTke2025meta structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
-Status Reporting and Progress Visibility
+Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules.
+For example, SciAgent [139] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG).
+Similarly, MetaGPT[140] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
+Status Reporting and Progress Visibility.
 Clear reporting of intent and intermediate status is also essential.
 Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions.
-Tools like AutoGenautogen_agents_tutorial and ChatDevqian2023chatdev have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
-Human-in-the-Loop Oversight
+Tools like AutoGen[108] and ChatDev[17] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
+Human-in-the-Loop Oversight.
 Human oversight and intervention mechanisms further reinforce trust.
-Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority wang2024ali ; kraprayoon2025ai ; sifakis2025reference .
-Many systems adopt a human-in-the-loop (HITL) paradigm chen2025human , where agents request confirmation before executing high-risk or irreversible actions.
-Behavioral Consistency and Bounded Autonomy
+Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [148, 149, 150].
+Many systems adopt a human-in-the-loop (HITL) paradigm [151], where agents request confirmation before executing high-risk or irreversible actions.
+Behavioral Consistency and Bounded Autonomy.
 Trust requires predictability.
 Agents should follow defined roles, output in expected formats, and remain within delegated authority.
-Example In enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material.
-Similarly, in autonomous research assistants like ChemCrow bran2023chemcrow , agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
-Social Trust Cues and Language Behaviors
-Beyond system-level mechanisms, behavioral consistency and bounded autonomy are crucial.
+For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material.
+Similarly, in autonomous research assistants like ChemCrow [141], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
+Social Trust Cues and Language Behaviors.
+Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial.
 If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it.
-Behavioral alignment mechanisms such as predefined role protocols, output style consistency, and language modeling constraints help standardize responses seeber2020collaborating ; acharya2025agentic ; shukla2024ai .
-Example In adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries cordova2025ai ; strielkowski2025ai .
-Lastly, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings sandini2024artificial .
-Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence kolt2025governing ; kirchhof2025position .
+Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [152, 16, 153].
+For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries [142, 143].
+Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [154].
+Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [144, 145].
 This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability.
-Together, these mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
-As autonomy and complexity increase, combining transparency, oversight, and social alignment will be essential to sustain user confidence.
-6.2 Explainability Techniques
-Explainability remains a cornerstone in fostering trust, accountability, and reliability in Agentic AI systems, particularly as they operate in high-stakes environments where multi-agent coordination and autonomous decisions directly impact human lives acharya2025agentic .
-In contrast to conventional AI systems, Agentic AI introduces unique challenges for explainability due to its decentralized architecture, dynamic role assignment, and evolving task decomposition among multiple interacting agents sapkota2025ai .
+These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
+7.2 Explainability Techniques in AMAS
+Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system.
+In platforms such as ChatDev [17] or AutoGen [18], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial.
+While local explainability techniques such as LIME [100] and SHAP [101] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory.
+Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [155], causal influence chains [156], and multi-agent SHAP adaptations [157].
+These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
+Symbolic and Hybrid Architectures.
+Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5].
+These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding.
+Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
+Lightweight Interpretability: Attention and Prompts.
+Attention map visualizations have been used to highlight focus areas in multi-modal language agents [158, 159, 160], offering lightweight but informative views into model behavior.
+Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [161, 162].
+These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
 Table 6: Explainability Techniques in Agentic AI Systems
 Technique Purpose Challenges in Agentic AI Example Use Cases
-LIME ribeiro2016should Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
-SHAP lundberg2017unified Attribution of prediction to input features via Shapley values Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
-Decision Provenance Graphs tupayachi2024towards ; bai2024derived Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
-Multi-agent SHAP tian2025outlook Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
-Symbolic/Rule-Based Agents sapkota2025ai Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
-Attention Maps liu2024ava ; hu2024overview Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
-Prompt Audit Trails gu2024artificial ; chin2025automating Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
-Local Post-hoc Techniques (LIME and SHAP)
-Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and SHapley Additive exPlanations (SHAP) lundberg2017unified are widely adopted techniques that offer post-hoc interpretability.
-LIME approximates a black-box model locally using an interpretable surrogate model, while SHAP attributes predictions to input features via Shapley values.
-These techniques have been integrated into agentic pipelines, particularly in finance and multi-agent fraud detection systems, where feature-level transparency supports regulatory compliance.
-However, their direct application in Agentic AI is limited.
-Each agent may operate with its own objectives, context, and tool access, leading to divergent decision paths that local techniques struggle to reconcile rosenfeld2019explainability ; glanois2024survey .
-Explainability in AMAS
-Emergent behavior poses another challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system.
-In platforms such as ChatDev qian2023chatdev or AutoGen wu2023autogen , agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial.
-To address this, researchers have proposed composite frameworks that combine local explanations with global decision traceability yossef2024explainable ; chander2025toward ; wang2024effects .
-For example, decision provenance graphs visualize communication flows and interdependencies across agents tupayachi2024towards ; bai2024derived , while causal influence chains track how actions propagate between roles rawal2025causality ; qiu2025reinforcement .
-Adapted SHAP techniques for multi-agent setups now aim to attribute outcomes to shared memory and agent collaboration tian2025outlook .
-Symbolic and Hybrid Architectures
-Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures sapkota2025ai .
-These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding.
-Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
-Lightweight Interpretability: Attention and Prompts
-Attention map visualizations have been used to highlight focus areas in multimodal language agents liu2024ava ; ye2024generative ; hu2024overview , offering lightweight but informative views into model behavior.
-Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction gu2024artificial ; chin2025automating .
-These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
-Ongoing Challenges and Research Directions
+LIME [100] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
+SHAP [101] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
+Decision Provenance Graphs [155, 163] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
+Multi-agent SHAP [157] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
+Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
+Attention Maps [158, 160] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
+Prompt Audit Trails [161, 162] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
 Despite these advancements, achieving robust explainability in Agentic AI systems remains an open research problem.
@@
 Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy.
-We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI’s performance and its real-world implications..
-Below we discuss these metrics and summarize in Table 7.
-Trustworthiness: This category evaluates the reliability, safety, and ethical alignment of the AI agent.
+We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI system performance and its real-world implications.
+Below, we discuss these metrics and summarize in Table 7.
+Trustworthiness.
+This criteria evaluates the reliability, safety, and ethical alignment of the AI agent.
 A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior.
-Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) liu2023gevalnlgevaluationusing .
+Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [164].
 Some approaches combine such factors into an overall trust score.
@@
 In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
-Explainability: Explainability metrics assess how well the agent’s decisions can be understood and traced by humans munoz2024explainability .
+Explainability.
+Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [165].
 These metrics focus on the clarity and completeness of the rationale the agent provides for its actions.
@@
 Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process.
-Quantitatively, methods like OpenXAI provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations agarwal2024openxaitransparentevaluationmodel .
+Quantitatively, methods like OpenXAI [166] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations.
 High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process.
 In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
-User-Centered Performance: User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs.
-These criteria emphasize the user’s experience and outcomes rodden2010measuring .
-Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the agent’s behavior.
-Task success from the user’s perspective (did the agent fulfill the user’s request or solve the user’s problem?) is a fundamental measure.
+User-Centered Performance.
+User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs.
+These criteria emphasize the user’s experience and outcomes [167].
+Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
 Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation.
 Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions.
-Ultimately, a user-centered agentic system should align its actions with user intent and preferences.
-Benchmarks like ChatDev qian2023chatdev , which simulates a multi-agent software development team interacting via natural language, implicitly evaluate how well agents fulfill user-defined roles and requirements in a collaborative project.
-This highlights the importance of user-oriented success in complex, realistic tasks.
+Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
 Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
@@
 Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
-Coordination (Multi-Agent or Modular): In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together.
+Coordination.
+In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together.
 Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently.
 Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
-One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working at cross-purposes).
-Multi-agent frameworks such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
+One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis).
+MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
 Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings.
@@
 High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
-Composite Metrics: Composite metrics aggregate multiple evaluation aspects into a single overall score.
-These are useful for summarizing an agent’s performance holistically, especially when comparing different systems.
+Composite Metrics.
+Composite metrics aggregate multiple evaluation aspects into a single overall score.
+These are useful for summarizing agents’ performance holistically, especially when comparing different systems.
 A composite metric is often a weighted combination of the above categories, for example:
 $M_{\text{composite}}=w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,,$ (2)
-where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application.
+where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application.
 The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
 An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure.
-By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of agentic systems.
-For instance, AgentBench liu2023agentbench is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent.
+By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems.
+For instance, AgentBench [168] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent.
 Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others.
@@
 8 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
+In this section, we discuss the security and privacy in AMAS.
+Table 8: Governance Dimensions for Agentic AI Systems
+Governance Component Description Example Tools or Methods
+Regulatory Compliance Aligning with legal and industry standards for responsible AI [169] NIST AI RMF [87], EU AI Act [170], GDPR [91], HIPAA [92]
+Auditability and Logging Capturing traceable records of agent actions and decisions [171] Decision provenance, immutable logs, blockchain-based traceability
+Policy Enforcement Enforcing system-level operational, ethical, and security constraints [53, 172] RBAC/ABAC policies [173], formal logic rules, dynamic sandboxing [60, 174, 175]
+Human Oversight Human-in-the-loop [151] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
+Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [176] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
+Explainability Governance Ensuring interpretability standards are met in decision pipelines [177] LIME/SHAP [178, 179, 180], decision trees [181, 182], interpretable surrogate models [183, 184, 185]
+Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [144, 146] Governance-as-code [114, 186], auto-updated rule engines [187, 188, 189], learning-based policy adaptation [126, 190]
+Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
 8.1 Security Mechanisms
-Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents khan2024security .
-Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness wang2024large .
-Among the foundational techniques employed are encryption he2024security ; motwani2024secret , access control huang2025agent ; south2025authenticated , adversarial defense shayea2025strategies ; malik2024systematic , and runtime monitoringnarajala2025securing ; chan2024agentmonitor each adapted to the unique demands of decentralized multi-agent environments.
-Encryption plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved motwani2024secret ; feretzakis2024privacy ; bommi2025enhancing .
+Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [191].
+Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [192].
+Among the foundational techniques employed are encryption [193, 194], access control [195, 171], adversarial defense [196, 197], and runtime monitoring[198, 199] each adapted to the unique demands of decentralized multi-agent environments.
+Encryption.
+plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [194, 61, 200].
 Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts.
-Implementations such as SSL/TLS, homomorphic encryption jency2025finquaxbot ; bommi2025enhancing , and secure enclaves (e.g., Intel SGX  [^1] https://www.intel.com/content/www/us/en/products/docs/accelerator-engines/software-guard-extensions.html) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
-Access control becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
-For instance, in systems like AutoGenwu2023autogen and CrewAI crewAI where agents take on specialized roles (e.g., summarizer, planner, coder) epperson2025interactive , enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation wu2023autogen ; narajala2025securing ; ismail2025toward .
-Agent-based access control policies often aligned with Role-Based Access Control (RBAC) khan2024role ; huang2025novel and Attribute-Based Access Control (ABAC) neupane2025towards ; huang2025novel paradigms can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels sapkota2025ai .
-Adversarial robustness is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results liu2024exploring ; zhang2024breaking .
-Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others zhang2024breaking ; wang2024large ; deng2025ai ; wang2025comprehensive ; patlan2025real ; patlan2025ai .
-Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities standen2025adversarial ; deng2025ai .
+Implementations such as SSL/TLS, homomorphic encryption [201, 200], and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
+Access control.
+becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
+For instance, in systems like AutoGen[18] and CrewAI [81] where agents take on specialized roles (e.g., summarizer, planner, coder) [202], enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [18, 198, 203].
+Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [204, 205] and Attribute-Based Access Control (ABAC) [206, 205] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels [5].
+Adversarial robustness.
+is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results [207, 208].
+Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others [208, 192, 115, 55, 116, 209].
+Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities [210, 115].
 Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
-Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity chan2024visibility .
-Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers zhen2025novel ; sapkota2025vibe ; xu2025forewarned .
+Runtime monitoring.
+systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity [211].
+Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [212, 213, 125].
 For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
-As Agentic AI continues to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount.
+As Agentic AI systems continue to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount.
 Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
@@
 To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
-Differential privacy (DP) offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified.
+Differential privacy (DP).
+DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified.
 In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information.
-For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data lazaros2024federated ; luzon2024tutorial .
+For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [214, 215].
 DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
-Data minimization is another cornerstone of privacy preservation.
+Data minimization.
 Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution.
 For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data.
-Furthermore, anonymization and pseudony luzon2024tutorial , techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
-Secure computation techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
+Furthermore, anonymization and pseudony [215] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
+Secure computation.
+Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
 In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage.
@@
 As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
-Table 8: Governance Dimensions for Agentic AI Systems
-Governance Component Description Example Tools or Methods
-Regulatory Compliance Aligning with legal and industry standards for responsible AI raza2025responsible NIST AI RMF dotan2024evolving , EU AI Act euregulationai2024 , GDPR gdpr25 , HIPAA hipaa164
-Auditability and Logging Capturing traceable records of agent actions and decisions south2025authenticated Decision provenance, immutable logs, blockchain-based traceability
-Policy Enforcement Enforcing system-level operational, ethical, and security constraints de2025open ; chen2024trustworthy RBAC/ABAC policies aba2023ai , formal logic rules, dynamic sandboxing zhong2023towards ; ali2018agent ; golechha2025among
-Human Oversight Human-in-the-loop chen2025human decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
-Risk Monitoring Continuous identification and mitigation of systemic or emergent risks slattery2025airiskrepositorycomprehensive Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
-Explainability Governance Ensuring interpretability standards are met in decision pipelines arrieta2020explainable LIME/SHAP salih2025perspective ; donmez2025explainable ; knab2025lime , decision trees yamada2025ai ; walia2025computational , interpretable surrogate models fonseca2025optimal ; mahmud2025optimizing ; xi2025model
-Adaptive Governance Updating governance mechanisms based on model evolution and deployment context kolt2025governing ; hughes2025ai Governance-as-code Responsi73:online ; kose2025modern , auto-updated rule engines goh2024campus ; yu2024shadow ; yun2025generative , learning-based policy adaptation feng2025multi ; li2025investigation
-Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
 8.3 Compliance and Governance
 As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative.
-Unlike traditional AI agents, agentic systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications.
+Unlike traditional AI agents, Agentic AI systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications.
 Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
-Regulatory standards provide the baseline requirements that all AI systems, including agentic architectures, must meet.
+Regulatory standards.
+Standards for regulation provide the baseline requirements that all AI systems, including agentic architectures, must meet.
 Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness.
 These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation.
-For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring, documentation of decision logic, and human oversight mechanisms attributes directly relevant to Agentic AI.
-Auditability is critical in ensuring transparency and facilitating post-hoc accountability.
+For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring of such systems.
+Auditability.
+This is critical in ensuring transparency and facilitating post-hoc accountability.
 Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification.
@@
 Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
-Policy enforcement governs what agentic systems can and cannot do, and under which conditions.
+Policy enforcement.
+Policy governs what Agentic AI system can and cannot do, and under which conditions.
 These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions.
 Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level.
-Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent subcomponents.
+Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent sub-components.
 Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
-Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the agentic system scales or changes context.
+Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the Agentic AI systems scale or change context.
 These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions.
-Below, we summarize the core governance components required for secure and compliant Agentic AI deployment.
-Thus, governance in Agentic AI systems is not a static compliance checkbox but a dynamic, adaptive layer embedded into the orchestration and operational pipelines.
-As these systems become more autonomous and integrated across sectors, aligning governance with system behavior, human values, and evolving legal standards is imperative for safe and trustworthy deployment, as summarized in Table 8.
+Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
 9 Discussion
-our exploration of TRiSM-based governance for LLM-powered Agentic AI systems reveals critical insights into technical design, ethical oversight, regulatory alignment, and future challenges.
-Below, we discuss the broader implications of our findings, structured into key areas for clarity.
+We discuss the broader implications of our findings below:
 9.1 Technical Implications of TRiSM for LLM Agent Design
-The AI TRiSM framework (Trust, Risk, and Security Management) imposes concrete technical requirements on how autonomous LLM-driven agents are built and deployed.
-A core implication is the need to embed real-time monitoring and control mechanisms into agent architectures.
-Rather than treating LLMs agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” guardrails_ai_guardrails_2024 .
-For example, there is discusison on designing specialized “guardian agents” within AMAS securiti2025aitrism .
+Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [216].
+For example, there is discussion on designing specialized “guardian agents” within AMAS [40].
 In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g.
 blocking disallowed actions such as outputting personally identifiable information).
-This layered control strategy transforms the technical architecture: an autonomous LLM agent is now supplemented by meta-agents that supervise its inputs, outputs, and tool use in real time.
-Prior research highlights risks like “excessive agency” owasp2025agentic where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals).
-TRiSM-driven agentic design mitigates these failure modes by constraining agent autonomy within well-defined safety bounds.
-Likewise, emerging threats specific to Agentic AI such as prompt injection kosinski2024prompt attacks, memory poisoning atkins2023memorypoisoning , or cascading hallucinations huang_survey_2023 underscore the need for built-in risk controls.
-By incorporating anomaly detection and policy-checking modules (as per the Sentinel/Operator model), an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g.
+Prior research highlights risks like “excessive agency” [106] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals).
+TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds.
+Likewise, emerging threats specific to Agentic AI systems such as prompt injection [23] attacks, memory poisoning [137], or cascading hallucinations [217] highlights the need for more built in risk controls.
+By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g.
 masking a sensitive datum or stopping an unsafe action) .
-This aligns with calls in the security community for proactive measures in Agentic AI, blending traditional cybersecurity techniques (access control, logging, sandboxing) with AI-specific safeguards like LLM firewalls and adversarial robustness checks raza-etal-2022-accuracy .
-In summary, TRiSM’s technical implications mean that autonomous LLM agents should no longer be deployed as stand-alone intelligent actors; instead, they operate under an active governance fabric of monitors, validators, and enforcement agents that ensure trustworthiness and safety by design.
 9.2 Ethical and Societal Ramifications of Multi-Agent AI
 Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions.
-Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness raza2025responsible all of which are vital for public trust in AI systems.
-A central concern is accountability: when AI agents make autonomous decisions that affect humans, who is answerable for the outcomes?
-TRiSM-based governance insists that organizations retain clear responsibility for their AI’s actions, rather than obscuring blame behind algorithmic “black boxes”.
-This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed gu2024artificial .
-Recent guidance on trustworthy AI chen2024trustworthy often highlights accountability and explainability as key pillars of trust.
-In practice, our approach means each autonomous agent’s decisions should be transparent enough to be understood and challenged by human reviewers when necessary.
+Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [169], all of which are vital for public trust in AI systems.
+Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”.
+This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed [161].
+Recent guidance on trustworthy AI [172] often highlights accountability and explainability as key pillars of trust.
+In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [218].
 Human oversight is another ethical imperative tightly coupled with accountability.
-TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under defined governance.
-Human operators or “AI managers” must have the ability to intervene or override when an agent’s behavior deviates from acceptable bounds or when moral judgment is required.
-Indeed, high-level policy frameworks (such as the EU’s AI ethics guidelines euregulationai2024 ) explicitly call for “human agency and oversight” in AI systems.
-In multi-agent setups, this may involve dashboard interfaces where humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly feng2025multi .
-The risk of “user complacency”, trusting an autonomous agent too much , has been noted as a hazard parasuraman2010complacency .
+TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [219].
+In MAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly [126].
+The risk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [220].
 TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
-Fairness and bias mitigation are also critical societal considerations.
-Our governance approach therefore incorporates bias audits and fairness checks throughout the agent lifecycle raza_fair_2024 .
-Techniques like pre-deployment bias testing, continuous monitoring for disparate impacts, and diverse stakeholder evaluation panels can be employed.
-These measures echo regulatory expectations; the EU’s AI Act euregulationai2024 and related guidelines enumerate “diversity, non-discrimination and fairness” as core requirements for trustworthy AI.
-In deploying LLM-based agents, we must ensure they do not treat individuals or groups inequitably, for example, content-filtering agents should apply policies uniformly across demographic groups, and task-planning agents should not propagate historical biases in resource allocation decisions.
-In sum, TRiSM-oriented governance extends beyond preventing technical failures: it seeks to uphold ethical norms and human rights, ensuring that autonomy in AI does not come at the expense of justice, transparency, or human dignity.
 9.3 Alignment with Emerging AI Regulations and Standards
@@
 This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards.
-For example, the European Union’s AI Act eu_ai_act (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems.
-These are precisely the capabilities that a TRiSM approach cultivates.
-By instituting continuous risk assessment, documentation of AI decision processes, and oversight mechanisms, a TRiSM-governed multi-agent system inherently addresses many of the EU Act’s requirements (e.g.
-having a risk management system and post-market monitoring for AI.
-Notably, the Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM’s security management component is designed to ensure (through adversarial resilience, access control, etc.).
-Similarly, international AI governance standards are emerging that mirror TRiSM’s tenets.
-ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development
-TRiSM’s trust and risk management focus naturally encompasses these elements: for instance, trust in TRiSM relates to reliable, truthful outputs (promoting transparency), while explicit risk management aligns with accountability for negative outcomes.
-By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 iso42001 and similar standards call for (e.g.
+For example, the European Union’s AI Act [94] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems.
+These are precisely the capabilities that a TRiSM approach promotes.
+The EU AI Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM security management component is designed to ensure (through adversarial resilience, access control, etc.).
+Similarly, international AI governance standards are emerging that mirror TRiSM’s principles.
+ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development.
+By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 [89] and similar standards call for (e.g.
 leadership oversight, documented risk controls, ongoing monitoring and improvement cycles).
 Another example is the U.S.
-NIST AI Risk Management Framework dotan2024evolving , which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
-By following TRiSM guidelines: for example, maintaining an “AI catalog” of all models/agents in use and their purposes, enforcing policies via sentinel/operator agents, and logging every AI decision, organizations create an audit-ready environment.
-In the event of an incident or an inquiry by authorities, they can demonstrate traceability and control over their autonomous agents, which will be crucial for regulatory compliance and liability management.
+NIST AI Risk Management Framework [87], which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
 9.4 Limitations and Current Research Gaps
-While the TRiSM-based approach appears promising, our work also revealed several limitations and open challenges in current research.
-First, limited benchmark evaluations pose a problem.
-The AI safety and agent governance community lacks widely accepted benchmarks to quantitatively assess trustworthiness or risk in multi-agent LLM systems.
-Unlike classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles.
+While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research.
+First, the classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles.
 This makes it difficult to compare different governance strategies or to track progress objectively.
-We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations standen2025adversarial or “red-teaming” exercises ganguli_red_2022 that can stress-test agentic systems and score their resilience (e.g.
-measuring success rates of prompt injection attacks or frequency of policy violations caught by oversight agents).
-Secondly, there is a paucity of real-world validation for many TRiSM-inspired controls.
-Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks charli_trism_2024 .
+We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [210] or “red-teaming” exercises [221] that can stress-test Agentic AI systems and score their resilience.
+Secondly, there is a lack of real-world validation for many TRiSM-inspired controls.
+Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [31].
 It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
-Additionally, integrating TRiSM with legacy systems poses practical challenges , for example, the earlier work notes compatibility issues when bolting on trust/security layers to existing AI pipelines.
-This suggests a limitation in how easily current AI deployments can fit TRiSM controls, a topic that merits further engineering research.
 Another critical gap is adversarial robustness.
-As we improve defenses, attackers will inevitably adapt.
-Recent findings show that LLM-based systems remain vulnerable to cleverly crafted attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails raza2025justhumansneedvaccines .
+As we improve defenses, attackers will inevitably adapt [222].
+Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [25].
 TRiSM solutions today are not foolproof against these.
-For example, an agent designed to mask secret data might itself be tricked into revealing it if the oversight logic fails to anticipate a new attack pattern.
-The literature identifies “evolving threats” and “adversarial attacks” as ongoing obstacles to trustworthy AI securiti2025aitrism .
-This underlines a need for continuous updates and adaptive security in any TRiSM implementation.
+We identify evolving threats and adversarial attacks as ongoing obstacles to trustworthy AI [40].
 Finally, organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures.
@@
 9.5 Future Roadmap for Agentic AI TRiSM
-Future Roadmap for Agentic AI Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
-Figure 3: Strategic roadmap for LLM-enabled Agentic AI, grouped into eight priority research and development domains.
-Drawing on our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 3.
+Table 9: Future Roadmap for Agentic AI: Eight Strategic Domains
+Scalable Agent Architectures Secure Multi-Agent Collaboration
+Explainable Multi-Agent Decisions Human-Centered Trust Design
+Autonomous Lifecycle Management Ethics & Governance
+Benchmarking & Evaluation Cognitive Capability Expansion
+Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure LABEL:fig:roadmap_mindmap.
 These recommendations span both technical system design improvements and governance-level policy initiatives:
-  • Develop Standardized Evaluation Benchmarks: The community should create open benchmarks and challenge environments to test multi-agent AI governance.
+Develop Standardized Evaluation Benchmarks.
+The community should create open benchmarks and challenge environments to test multi-agent AI governance.
 For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls.
@@
 frequency of prevented failures or fairness outcomes).
-  • Advance Adversarial Robustness Techniques: Future system design must anticipate a continually evolving threat landscape.
+Advance Adversarial Robustness Techniques.
+Future system design must anticipate a continually evolving threat landscape.
 Techniques from cybersecurity (e.g.
-adversarial training, AI model “penetration testing” bacudio2011overview , and formal verification) should be integrated into the LLM agent development pipeline.
+adversarial training, AI model “penetration testing” [223], and formal verification) should be integrated into the LLM agent development pipeline.
 Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions.
-Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames schneider2020wargaming , can help discover vulnerabilities in a controlled way before real adversaries do.
-  • Human-Centered Oversight Tools: We encourage designing better interfaces and protocols for human oversight of Agentic AI.
-Borrowing from human-computer interaction AmershiWVFNCSIB19 and cognitive engineering hexmoor2025behaviour , researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
+Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [224] , can help discover vulnerabilities in a controlled way before real adversaries do.
+Human-Centered Oversight Tools.
+We encourage designing better interfaces and protocols for human oversight of Agentic AI systems.
+Borrowing from human-computer interaction [225] and cognitive engineering [226], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
 Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials.
-These would be controlled environments where innovators can deploy Agentic AI under supervision, demonstrating TRiSM controls to regulators.
+These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators.
 Insights from such pilots can inform refinements in both technical standards and regulations.
@@
 10 Conclusion
-TRiSM-based governance offers a promising scaffold to ensure that autonomous LLM-powered agents are trustworthy, accountable, and secure.
-Our discussion has analyzed how this framework influences technical design decisions, mandates ethical guardrails, and dovetails with emerging regulatory regimes.
-While current research is nascent and not without limitations, the path forward is clear.
-By rigorously testing these systems, strengthening them against adversaries, and crafting policies and standards in tandem with technological advances, we can enable powerful multi-agent AI systems to operate beneficially under robust oversight.
-The stakes are high but with a proactive, interdisciplinary approach, we can achieve a balance where innovation in AI goes hand-in-hand with responsibility and trust.
-As future work tackles the open challenges identified, we anticipate that TRiSM principles will transition from conceptual best-practice to standard operating procedure for Agentic AI, ensuring these systems earn and maintain the confidence of all stakeholders involved.
+The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services.
+However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations.
+This review has systematically explored these challenges, proposing a TRiSM framework structured around four pillars: Governance, Explainability, ModelOps, and Privacy/Security, specifically adapted for LLM-based Agentic Multi-Agent Systems (AMAS).
+Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency.
+The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
+Looking forward, the proposed research roadmap emphasizes scalable, verifiable, and regulation-compliant Agentic AI systems.
+Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination.
+By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
+Declaration of Competing Interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Appendix
+Table A.1: Key terminologies for LLM-based agentic AI systems.
+Term Definition
+Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
+Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
+Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
+Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
+Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
+Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
+Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
+ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
+Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
+Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
+Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
+Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
+Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
+Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
+Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
+Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
+Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
+Figure A.1: Traditional AI Agent vs. Agentic AI.

### v3

@@
 We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents.
-We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Governance, Explainability, ModelOps, and Privacy/Security , each contextualized to the challenges of multi-agent LLM systems.
+We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Explainability, ModelOps, Security, Privacy and Governance, each contextualized to the challenges of multi-agent LLM systems.
 A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation.
@@
 While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4].
-These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5, 6].
-This rapid evolution presents new opportunities but also surfaces significant challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
-In particular, the shift toward Agentic AI architectures amplifies and evolves longstanding concerns around safety, fairness, interpretability, and systemic harms.
-While such issues have been central to machine learning (ML) systems broadly, they become particularly critical in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
-From isolated agents to Agentic AI The autonomous software agents date back to the 1990s [7], however, the advent of LLM-driven multi-agent ecosystems [8] since 2023 has made “ Agentic AI ” a distinctly new and rapidly expanding research focus [5].
-Traditional agents primarily automated narrow tasks, such as information retrieval [9, 10], data summarization [11], and dialogue response [12].
-Typically driven by single step logic [13] or scripted rules, these agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving [14, 15].
-In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use [16].
-These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [17, 18, 19].
+These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5].
+This rapid evolution presents new opportunities but also brings challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
+In particular, as the shift toward agentic AI architectures amplifies and evolves, longstanding concerns around safety, fairness, interpretability, and systemic harms resurface.
+While such issues have been central to machine learning (ML) systems broadly, they become more promiment in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
+From Isolated Agents to Agentic AI The autonomous software agents date back to the 1990s [6], however, the advent of LLM-driven multi-agent ecosystems [7] since 2023 has made “Agentic AI”, a distinctly new and rapidly expanding research focus [5].
+Traditional agents primarily focussed on automating limited tasks, such as information retrieval [8], data summarization [9], and dialogue response [10].
+Typically driven by single step logic [11] or scripted rules, these traditional agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving .
+In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), and enabled by LLMs and tool use [12].
+These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [13].
 This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
 1.2 Trust, Risk, and Security Challenges in Agentic AI
-The rise of Agentic AI introduces complex challenges that extend beyond those of traditional ML systems.
-Unlike deterministic agents [20], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
+The rise of Agentic AI introduces new challenges that extend beyond those of traditional ML systems.
+For example, unlike deterministic agents [14], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
 These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
-Trust and Trustworthiness
-In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust.
+Below, we discuss some of these challanges:
+  1. 1.
+    Trust and Trustworthiness In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust.
 This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications.
@@
 These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
-Security Risks in Multi-Agent Systems
-Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [22].
+  2. 2.
+    Security Risks in Multi-Agent Systems Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [16].
 For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate.
 The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
-Specific Threats: Prompt Injection, Spoofing, and Impersonation
-LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors.
-One prominent threat is prompt injection [23], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior.
-Recent studies [24] describe “prompt infection,” a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [25].
+  3. 3.
+    Specific Threats: Prompt Injection, Spoofing, and Impersonation LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors.
+One prominent threat is prompt injection [17], where attackers design inputs with hidden or malicious instructions to manipulate agent behavior.
+Recent studies [18] describe “prompt infection” as a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [19].
 Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
     Another critical vulnerability involves spoofing and impersonation [20].
-Spoofing entails faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges.
+Spoofing means faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges.
 In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
-Privacy Risks in Multi-Agent LLM Systems
-LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [27].
+  4. 4.
+    Privacy Risks in Multi-Agent LLM Systems LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [21].
 In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g.
 via a shared memory store or message passing) to collaborate.
-Limitations of Traditional Frameworks
-Traditional evaluation and safety frameworks, designed for deterministic agents or conventional AI systems, are increasingly inadequate for addressing the autonomy and complexity of Agentic AI.
-These frameworks often fail to account for emergent behaviors, inter-agent dependencies, or the dynamic nature of LLM-driven systems, leaving gaps in security, trust, and risk management.
-Towards a TRiSM Framework for Agentic AI
-To address these gaps, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) [28, 29] propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management.
-These components are vital for deploying Agentic AI in high-stakes domains such as finance, healthcare, and defense, where trust, security, and accountability are non-negotiable.
+Traditional evaluation and safety frameworks, designed for deterministic agents AI systems, are inadequate for addressing the autonomy and control of Agentic AI.
+To address these gaps, frameworks like AI TRiSM (Artificial Intelligence Trust, Risk, and Security Management) [22, 23] proposes lifecycle-level controls, including explainability, secure model orchestration, and privacy management.
+AI TRiSM is a framework to help organizations build AI systems that people can trust.
+The goal is to focus not just on technical performance, but also on ethical values like fairness, transparency, and security.
+The framework brings together practices that support governance, strengthen reliability and robustness, and protect data and make sure that AI systems are not only effective, but also aligned with human values.
 Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review.
@@
 Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks.
-However, no unified framework exists for managing their Trust, Risk, and Security Management (AI TRiSM) across dynamic, multi-agent workflows.
-Recent failures, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [30].
+However, no unified framework exists for managing the trust, risk, and security across dynamic as well as multi-agent workflows.
+Recent accidents, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [24].
 These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways.
-To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is urgently needed.
-This study summarizes emerging findings and offers actionable guidelines to build secure, transparent, and trustworthy Agentic AI ecosystems.
+To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is therefore urgently needed.
 1.4 Scope and Objectives
-Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [28] landscape as it applies to LLM-powered, multi-agent “Agentic AI” systems.
+Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [22] framework to LLM-powered, multi-agent “Agentic AI” systems.
 We structure the discussion around four foundational pillars of TRiSM [25]: explainability, security, lifecycle governance, and privacy.
-We argue that conventional safeguards developed for bounded, rule-based agents are inadequate for managing the complexities of open-ended, self-directed agent collectives.
 The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
-For the sake of this study, we define Agentic AI as Multi-Agent Systems (MAS) [32] powered by LLMs that exhibit autonomous planning, tool use, and memory retention, with or without human oversight.
-In contrast to classical MAS, these systems exhibit emergent reasoning capabilities and dynamically adapt their workflows based on interactions with tools, users, and other agents.
-Scope This paper is intended as a conceptual survey and synthesis.
-It does not propose a new Agentic AI architecture or introduce synthetic benchmark evaluations.
-Rather, it integrates emerging perspectives on trust, risk, and security to support future research and responsible development in this evolving area.
+For the purposes of this study, we define Agentic AI as Multi-Agent Systems (MAS) [26] powered by LLMs that exhibit autonomous planning, tool use, memory retention, and emergent reasoning capabilities, with or without human supervision.
+These LLM-based agents can dynamically adapt workflows based on interactions with tools, users, and other agents, distinguishing them from traditional symbolic or rule-based agents.
+Scope This paper is a conceptual survey and synthesis.
+It does not propose a new architecture for Agentic AI, nor does it include synthetic benchmark evaluations.
+Instead, it integrates current perspectives on AI TRiSM to inform future research and the responsible development of Agentic AI systems.
+The scope of this paper is confined to LLM based Agentic AI systems and AMAS  [^3] Throughout this paper, AMAS designates multi-agent systems built on LLMs.
 Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability.
-As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance—the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks.
+As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks.
 This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
@@
 Our key contributions in this review are:
-  1. 1.
-    We propose the first TRiSM framework that is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate.
-The framework is organized around four pillars: Explainability, Lifecycle Governance, Application Security, and Model Privacy.
-It targets the specific governance challenges created by autonomous decision making, inter-agent coordination, and persistent memory (see Fig.
+  • We propose the first TRiSM framework for Agentic AI systems.
+This framework is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate.
+The framework is organized around pillars: Explainability, ModelOps, Security, Privacy and Governance around autonomous decision to guide the development and use of safe and ethical Agentic AI systems.
+(see Fig.
 1 and Sec.
+6).
+  • We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec.
 5).
-  2. 2.
-    We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec.
-6).
-This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats, and it shows why defenses designed for single-model LLM applications are not sufficient.
-  3. 3.
-    We introduce a five-dimension evaluation framework that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec.
-7.3).
+This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats commonly seen in AMAS.
+  • We introduce an evaluation dimensions for Agentic AI systems that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec.
+8).
 To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
-  4. 4.
-    We review state-of-the-art techniques for each TRiSM pillar (Secs.
-78) and examine methods such as Local Interpretable Model-Agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), decision provenance, prompt hygiene, sandboxing, and ModelOps (model operations) pipelines.
-These techniques must be critically re-evaluated and adapted to suit environments involving multiple interacting LLM agents.
-  5. 5.
+Our aim is to achieve high performance in both dimensions, accuracy and usability, without sacrificing one.
+  • We contribute to the management and control of Agentic AI by mapping existing TRiSM techniques such as explainability methods, prompt hygiene, decision provenance, sandboxing, and ModelOps pipelines, to AMAS environments.
+We critically analyze their limitations and adaptation needs, providing a foundation for re-engineering these techniques to fit the demands of interacting LLM agents (Secs.
+79).
   • Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 11).
-1.6 Paper Organization
-The remainder of this paper is structured as follows.
-Section 2 discuss the related work.
-Section 3 presents the literature search methodology.
-Section 4 reviews the core characteristics and architectures of Agentic AI systems, providing a foundation for the TRiSM analysis.
-Section 5 introduces the TRiSM framework and its four pillars: Explainability, ModelOps, Application Security, and Model Privacy.
-Section 6 outlines key risks, vulnerabilities, and attack surfaces in multi-agent AI ecosystems.
-Section 7 discusses trust-building and explainability techniques, while Section 8 covers security and privacy-enhancing methods.
-Section 9 synthesizes insights and proposes directions for research and policy.
-Section 10 concludes the paper with final reflections on the implications of TRiSM for the responsible development of Agentic AI.
-A list of key terms used throughout this paper is provided in Table A.1.
 2 Related Work
@@
 Fang et al. (2025) [32] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
+Xi et al. (2025) [33] ✓ (single-/multi-agent,
+human-AI coop)
+~ (open problems)
+Luo et al. (2025) [34] ~ ~ ✓ (science, games, etc.) ~ (open challenges)
+Zou et al. (2025) [35] ~ ✓ (e.g., finance,
+healthcare)
+~ (challenges & opp.)
+Wang et al. (2025) [36] ~ ~ ~ (healthcare domain) ~ (future opportunities)
 This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
@@
 We present the related work in this section.
-2.1 LLM-Based Multi-Agent Surveys (Technical Focus)
-Prior surveys on LLM-driven multi-agent systems predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI, yet they rarely address critical TRiSM (Trust, Risk, and Security Management) considerations.
+LLM-Based Multi-Agent Surveys (Technical Focus)
+Prior surveys on LLM-driven AMAS predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI.
+However, there remains a gap in addressing governance and policy aspects.
 For example, reviews like [27] and [28] delve into simulated environments, inter-agent communication protocols, and performance benchmarks.
-These works excel at detailing how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
-More specialized surveys, such as [35] on natural language communication and [36] on collaboration mechanisms, explore interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0.
-While [35] briefly notes security risks tied to communication channels, and [36] mentions deployment contexts, both remain narrowly focused on coordination strategies.
-They lack deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, or explainability challenges at scale.
-2.2 Trustworthy/Responsible AI Surveys (Broad Trust Focus)
-Surveys addressing TRiSM in AI, such as [38], tackle ethical and risk-related issues including alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage).
-However, these works typically target general ML systems [39] and overlook the unique dynamics of LLM-based multi-agent systems.
+These works mainly detail how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
+More specialized surveys, such as [29], focus on natural language communication, while [30] details collaboration mechanisms, exploring interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0.
+Although [29] highlights security risks associated with communication channels, and [30] discusses deployment contexts, both remain narrowly centered on coordination strategies.
+There is a need for deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability challenges at scale.
+Trustworthy/Responsible AI Surveys (Broad Trust Focus)
+There are some review articles that address TRiSM in AI, such as [32], tackle ethical and risk-related issues including alignment with human values, discuss fairness in decision-making [37], and defenses against privacy attacks (e.g., data leakage).
+However, these works typically target general ML systems and overlook the unique dynamics of LLM-based multi-agent systems.
 For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems.
-Moreover, critical TRiSM aspects [40] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent.
-As a result, their recommendations, while valuable for single-agent or centralized AI, do not fully translate to the complexities of multi-agent coordination in high-stakes domains.
-2.3 Current Study (LLM Multi-Agent)
-This review distinguishes itself by weaving TRiSM principles: adversarial robustness, end-to-end lifecycle governance, and system-wide explainability, into the study of LLM-based multi-agent systems, with a specific lens on high-stakes domains like healthcare, science, and finance.
-Unlike prior studies (see Table 1), which offer limited insight into governance and security, this work positions these elements as foundational design priorities.
-For example, it addresses targeted threats like backdoor attacks in distributed LLMs and governance challenges such as ensuring compliance with regulations (e.g., HIPAA in healthcare or GDPR in finance) across autonomous agents - areas largely ignored by technical surveys.
-This work provides actionable guidance for researchers, engineers, and policymakers, and fills critical gaps in both the technical multi-agent literature and broader responsible AI landscape.
+Moreover, critical TRiSM aspects [38] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent in recent discussions.
+Current Study (Multi-Agent Systems and LLMs)
+This review distinguishes itself by mapping TRiSM principles into the study of AMAS.
+Unlike prior studies (see Table 1), which offer limited insight into governance and policy, this work operationalizes TRiSM principles in practice.
+It provides actionable guidance for researchers, engineers, and policymakers, addressing critical gaps in both the technical AMAS literature and the broader responsible AI landscape.
 3 Literature Review Methodology
-To provide a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [41, 42].
-While this is not a formal systematic review, we followed a transparent and reproducible process to ensure coverage and rigor.
+To review the literature related to TRiSM as applied in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [39, 40].
 This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
@@
 Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
-3.1 Research Objectives
-The review was guided by the following four specific questions:
+Research Objectives The review was guided by the following four specific questions:
   • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems?
@@
   • RQ4: What gaps remain in current research, and what directions are promising for future work?
-3.2 Data Sources and Search Strategy
-To conduct a thorough literature review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
+Search Strategy To conduct a review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
 The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context.
@@
 The search terms were structured as follows:
-(“ Agentic AI” OR “multi-agent systems” OR “MAS” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
+(“ Agentic AI” OR “multi-agent systems” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
 AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
-This comprehensive search strategy ensured the identification of relevant studies addressing the integration of TRiSM principles with LLM-based multi-agent systems, particularly in high-stakes domains such as healthcare, science, and finance.
-3.3 Inclusion and Exclusion Criteria
-We applied the following criteria during the screening process:
-  • Inclusion:
+Inclusion and Exclusion Criteria We applied the following criteria during the screening process:
+Inclusion Criteria:
   • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
@@
   • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
-  • Exclusion:
+Exclusion Criteria:
   • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
@@
 An analysis on the literature is shown in Figure 2.
-3.4 Quality Assurance
-To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [43, 42].
+Quality Assurance To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [41].
 Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems?
@@
 In this section, we present the fundamentals of Agentic AI systems.
-4.1 Agentic AI vs. Traditional AI Agents
-Agentic AI systems fundamentally transform the design of autonomous systems, and are different from traditional AI agents.
-Traditional agents operate through predefined rules [44], heuristic workflows [45, 46], or deterministic logic [47], and perform well in narrow, well-defined environments.
-Cognitive Agents [48], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [49] or ACT-R [50].
-However, they remain constrained by static knowledge bases and limited adaptability [48].
-In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
-Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [17, 19].
-(2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [51].
-(3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [52].
-This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [53].
-For example, systems like AutoGen [18] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54].
-This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [19].
-A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
 Table 2: Traditional AI Agents vs. Agentic AI Systems
@@
 = Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
+4.1 Traditional AI Agents vs Agentic AI
+Traditional agents operate through predefined rules [42], heuristic workflows [43, 44], or deterministic logic [45], and perform well in narrow, well-defined environments.
+Cognitive Agents [46], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [47] or ACT-R [48].
+However, they remain constrained by static knowledge bases and limited adaptability [46].
+In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
+Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [13, 49].
+(2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [50].
+(3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [51].
+This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [52].
+For example, systems like AutoGen [53] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54].
+This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [49].
+A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
 4.2 LLM-Based Agentic AI System Architecture
@@
 For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
-5 The TRiSM Framework for Agentic AI
-The TRiSM Framework for Agentic AI is presented in this section.
-Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
-Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
-Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [87]
-ISO/IEC 24029-1 [88]
-Metrics: fidelity, sufficiency, human-trust score
-ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) [89]
-SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
-Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [90]
-Jailbreak-success%; mean-time-to-detect; exploit CVSS
-Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [91]
-HIPAA §164 [92]
-Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
-5.1 Governance for Agentic LLM Systems
-AI TRiSM is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards [29].
-Originally highlighted in industry guidelines for AI governance [93], TRiSM provides a structured approach to manage the unique challenges of LLM-based “ Agentic AI" systems (herein we refer to as Agentic LLM systems).
-The TRiSM framework addresses the trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 29].
-Figure 4 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
-Figure 4: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems.
-The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers.
-External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
-The TRiSM framework aligns closely with several key international AI governance efforts that shape best practices for safety, accountability, and transparency in AI systems.
-For example, the EU AI Act [94] (notably Article 14) mandates clear obligations for transparency, traceability, and human oversight in high-risk systems, principles directly reflected in TRiSM Explainability and ModelOps pillars.
-The NIST AI Risk Management Framework (AI RMF) [95] emphasizes Explainability, Governability, and Validity as core risk mitigation strategies, providing a conceptual match to TRiSM’s structured controls across lifecycle and system interaction points.
-Additionally, the OECD AI Principles advocate for transparency, robustness, and accountability [96]each echoed in TRiSM attention to secure tool invocation, memory traceability, and multi-agent auditability.
-These global frameworks reinforce the need for Agentic AI systems to support real-time monitoring, enforce role-based access, maintain verifiable logs, and provide intelligible outputs.
-As shown in Figure 5, TRiSM pillars provide operational scaffolding for implementing regulatory expectations in LLM-based multi-agent contexts.
-Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
-13–14) [94] and the NIST AI RMF “Explain” function [95].
-This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [97].
-ModelOps (Lifecycle Governance): Anchored in the NIST RMF “Govern” function [95] and ISO/IEC 42001 [98].
-It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates.
-Application Security: Guided by OWASP LLM Top-10 [90] and OECD AI Principles on Robustness [96], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms.
-Model Privacy: Aligned with GDPR (Art.
-25) [91], HIPAA §164 [92], and OECD AI Principles [96], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
-Figure 5: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, and OECD Principles.
-Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
-5.2 Explainability in Multi-Agent Decision Making
-Explainability refers to making the inner workings and decisions of AI agents interpretable to humans.
-In the context of AMAS, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [99].
-However, achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
-Established explainable AI techniques provide a starting point.
-For instance, LIME [100] and Shapley Additive Explanations (SHAP) [101] can be adapted to analyze LLM decisions.
-These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken.
-In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
-Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability.
-Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered [102], for example, if a particular agent’s contribution were removed or modified.
-This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
-Recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g.
-CoT prompts or dialogue between agents) can further enhance transparency.
-Human auditors can also help with a trace of how agents arrived at a decision, e.g., which agent contributed what information and why, to produce narrative explanations for its outcomes.
-5.3 ModelOps : Lifecycle Management for LLM Agents
-ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [103].
-It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production.
-Within Agentic AI systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59].
-Effective ModelOps is important to maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
-Agentic AI systems require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations.
-A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations.
-Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified.
-Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
-This aligns with best practices in LLM operations (LLMOps) [104], which integrate MLOps principles tailored to the challenges of LLMs.
-In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
-5.4 Application Security for LLM Agents
-To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents.
-This includes prompt hygiene [25] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions.
-Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [105].
-It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [106].
-Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
-Recent frameworks, such as LangChain [107], AutoGen [108], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
-Moreover, training LLM agents with adversarial robustness in mind (e.g.
-fine-tuning on adversarial examples, employing adversarial training regimes) [25] can improve their resistance to malicious inputs.
-5.5 Model Privacy and Data Protection
-The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [27].
-Since multiple agents may share information with each other (e.g.
-via a shared memory store or message passing) to collaborate.
-Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs.
-Therefore, TRiSM’s privacy pillar recommends organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
-To preserve privacy in multi-agent LLM systems, several technical measures are employed.
-Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [109].
-Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [110].
-Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [111].
-Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [112].
-Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [113].
-Finally, the implementation of model privacy policies and compliance frameworks—including data retention limits, audit logs, and adherence to regulations such as GDPR or HIPAA, ensures that agents operate within legal and ethical boundaries [114].
-Table 4 condenses each pillar into its governing goal.
 5 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
-In this section, we present the main threats in Agentic AI and also discuss a taxonomy of emergent risk classes.
+In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
 Figure 4: Overview of Key Risks and Threats in Agentic AI Systems.
@@
 5.1 Threats in Agentic AI
-Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [115]. Below, we discuss these threats.
+Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [87]. Below, we discuss some of these major threats.
   • Autonomy abuse.
-The foremost threat is autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs [116].
-Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states [117].
+The foremost threat is the autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs .
+Unlike deterministic models, these Agentic AI systems dynamically generate actions, complicating efforts to define and enforce safe operational states .
   • Persistent memory.
-Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential adversarial injections [118, 119].
-Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms [116].
+Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential prompt injections.
+Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms .
   • Agent orchestration.
 This risk involves central or distributed control mechanisms for role assignment and workflow mediation.
-A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [19] and AutoGen [18].
+A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT and AutoGen .
 These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
+  • Goal Misalignment and Exposure.
+Agentic AI systems often operate with internalized or user-specified goals.
+Improper scoping or exposure of goal representations can lead to prompt extraction, misalignment, or competitive agent behaviors, particularly when third-party tools or APIs are invoked dynamically during task completion.
+  • Tool Misuse and External API Exploits.
+Agents capable of invoking external tools (e.g., via APIs, code execution, or web access) present novel attack surfaces.
+Misuse of these tools, intentional or accidental, can trigger financial costs, Distributed Denial of Service (DDoS) behavior, or legal violations.
+Sandboxing and invocation tracing remain limited in most open-source frameworks.
+  • Multi-Agent Collusion or Drift.
+Emergent coordination among agents, while useful for collaboration, can also lead to collusion, competitive drift, or polarization of outputs, especially in self-reinforcing memory loops or agent swarms.
+Such behavior complicates interpretability and may undermine safety alignment goals.
 5.2 Taxonomy of Risks
-We propose a taxonomy of risks associated with autonomous multi-agent systems (AMAS), categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 6.
+We propose a taxonomy of risks associated with AMAS, categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 4.
 These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS.
@@
 Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps.
-These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [120, 121].
+These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [88].
 For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [13].
@@
   • Data Leakage.
-The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information [122, 123].
-In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [124, 125].
+The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information.
+In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [89].
 For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls.
@@
   • Agent Collusion and Mode Collapse.
-Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [126, 127].
+Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [90].
 This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies.
@@
   • Emergent Behaviors.
-The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [128, 129].
+The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [91].
 Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability.
-Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [130], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals [131].
+Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [92], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals .
 These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
-6.3 Real-World Use Cases
-Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental Agentic AI systems.
+5.3 Real-World Examples
+Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
 Prompt Leakage in Agentic AI Systems.
@@
 In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [94].
-Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks [134].
+Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks.
 Collusive Failure in ChatDev.
@@
 This led to hallucinated goals and emergent behaviors misaligned with initial intentions [98].
-7 Trust and Explainability in Agentic AI
+Next, we discuss the AI TRiSM prism framework and map its core principles to a AMAS.
+6 The TRiSM Framework for Agentic AI
+6.1 Overview of AI TRiSM and Its Core Components
+The TRiSM Framework, short for Trust, Risk, and Security Management, offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors.
+In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
+Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
+Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
+Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [99]
+ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
+Metrics: fidelity, sufficiency, human-trust score
+ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
+SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
+Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [102]
+Jailbreak-success%; mean-time-to-detect; exploit CVSS
+Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [103]
+HIPAA §164 [104]
+Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
+  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness.
+Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [50].
+This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
+  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies.
+This includes technical risks such as algorithmic bias [37], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [32].
+Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
+  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access.
+This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [19].
+Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
+Originally highlighted in industry guidelines for AI governance [105], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems.
+The TRiSM framework addresses trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 23].
+Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
+Figure 5: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems.
+The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers.
+External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
+6.2 Mapping Pillars of TRiSM with Agentic AI
+Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
+13–14) [106] and the NIST AI RMF “Explain” function [107].
+This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [108].
+ModelOps : Anchored in the NIST RMF “Govern” function [107] and ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101].
+It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates.
+Application Security: Guided by OWASP LLM Top-10 [102] and OECD AI Principles on Robustness [109], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms.
+Model Privacy: Aligned with GDPR (Art.
+25) [103], HIPAA §164 [104], and OECD AI Principles [109], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
+Governance: Grounded in overarching regulatory frameworks such as GDPR, CCPA [110], and the OECD AI Principles [109], this pillar supports institutional oversight, policy compliance, and cross-system accountability.
+It enforces auditable workflows, verifiable role-based access, regulatory traceability, and human oversight, particularly in multi-agent and high-risk AI deployments.
+Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, 42005, and OECD Principles.
+Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
+Explainability in Multi-Agent Decision Making
+Explainability refers to making the inner workings and decisions of AI agents interpretable to humans.
+In the context of LLM-based multi-agent systems, it is paramount for building user trust because outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [111].
+However, achieving explainability in such systems is particularly challenging , one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to the final outcomes [112].
+In other words, explaining why a multi-agent system behaved a certain way requires tracing which agent did what and how their collaboration produced an emergent result, a task that traditional single-model explainability techniques do not directly address.
+Several approaches are being explored to enhance explainability in LLM-driven multi-agent systems:
+  • Feature Attribution (LIME/ SHAP): Established explainable AI techniques provide a starting point.
+For instance, LIME) [113] and Shapley Additive Explanations (SHAP) [114] can be adapted to analyze LLM decisions.
+These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken.
+In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
+By highlighting the most influential parts of each agent’s input, methods like LIME/SHAP help clarify individual agent decisions in the multi-agent mix [115].
+  • Counterfactual Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability.
+Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [116].
+For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes.
+This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
+  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency.
+Logging the step-by-step reasoning process – for example, an agent’s chain-of-thought (CoT) prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [69].
+By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision.
+This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
+  • Natural Language Explanations: Finally, the generative nature of LLMs can be leveraged to turn the above insights and logs into human-friendly explanations [117].
+Instead of expecting end-users to read raw CoT logs or analyze feature weights, an “explainer” agent (which could be another LLM) or a human analyst can compile the reasoning trace into a narrative.
+In fact, an important aspect of explainability is communicating the reasons behind decisions in an intelligible way.
+By using natural language, the system can answer user questions like “Why did the agents decide to do X?” in a conversational manner.
+ModelOps : Lifecycle Management for LLM Agents
+ModelOps (Machine Learning Operations) is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [118].
+Within AMAS, ModelOps covers not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59].
+Effective ModelOps helps maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
+AMAS require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations.
+A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations.
+Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified.
+Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
+Application Security for LLM Agents
+To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents.
+This includes prompt hygiene [19] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions.
+Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [119].
+It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [120].
+Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
+Recent frameworks, such as LangChain [121], AutoGen [122], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
+Moreover, training LLM agents with adversarial robustness in mind (e.g.
+fine-tuning on adversarial examples, employing adversarial training regimes) [19] can improve their resistance to malicious inputs.
+Model Privacy and Data Protection
+The model privacy addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [21].
+Since multiple agents may share information with each other (e.g.
+via a shared memory store or message passing) to collaborate.
+Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs.
+To preserve privacy in AMAS, several technical measures can be employed.
+For example, Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [123].
+Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [124].
+Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [125].
+Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [126].Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [127].
+Governance
+Governance of AI systems necessitates compliance with evolving legal and policy frameworks, including the General Data Protection Regulation (GDPR) [103] and the California Consumer Privacy Act (CCPA) [110].
+Effective governance extends beyond regulatory adherence to cover proactive management of legal, ethical, and societal risks.
+This includes establishing transparent accountability structures, maintaining auditable documentation, and ensuring the protection of data subject rights throughout the AI lifecycle.
+Several global frameworks provide guiding principles that align with the TRiSM pillars.
+The EU AI Act [106], particularly Article 14, mandates transparency, traceability, and human oversight in high-risk systems; principles embedded within TRiSM’s Explainability and ModelOps dimensions.
+Similarly, the NIST AI Risk Management Framework (AI RMF) [107] identifies: Explainability, Governability, and Validity as key components of AI risk mitigation, which directly correspond to TRiSM’s lifecycle control mechanisms.
+The OECD AI Principles [109] reinforce the importance of robustness, transparency, and accountability, echoing TRiSM’s attention to secure tool invocation, memory traceability, and multi-agent auditability.
+Together, these frameworks underscore the need for governance architectures in Agentic AI systems that support real-time monitoring, role-based access, verifiable logs, and intelligible outputs.
+As illustrated in Figure 6, TRiSM offers a practical scaffolding for operationalizing these requirements in LLM-based multi-agent environments.
+7 Trust and Explainability in Agentic AI Systems
 Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
@@
 Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [133, 134] Enhances user comfort and trust in uncertain scenarios
-Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight.
-These systems also pose unique challenges and opportunities for human trust.
 Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms.
-We show the relationship between trust mechanisms, explainability, and evaluation facets in Figure 7 and discuss next.
+Figure 7 illustrates the layered relationship between trust mechanisms, explainability techniques, and evaluation facets.
+Trust-enabling mechanisms, such as transparency and human-in-the-loop (HIIL) oversight, provide the foundation for user confidence.
+These feed into explainability techniques like decision provenance graphs, which make agent behaviors interpretable.
+Finally, evaluation (Section 8) facets ensure that trust and explainability are quantifiable, enabling iterative improvements in system design.
 7.1 Building Trust in Autonomous Agents
@@
 Transparency is one of the core enablers of trust.
-For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable.
+For users to understand and evaluate AMAS decisions; the reasoning chains, decision states, and action triggers of agents must be made observable.
 Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules.
@@
 Behavioral Consistency and Bounded Autonomy.
-Trust requires predictability.
+Trust requires predictability in models’ behaviour.
 Agents should follow defined roles, output in expected formats, and remain within delegated authority.
@@
 Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [141, 12, 142].
-For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries [142, 143].
+For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries .
 Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [143].
-Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [144, 145].
+Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [134].
 This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability.
@@
 Prompt Audit Trails [150, 151] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
-Despite these advancements, achieving robust explainability in Agentic AI systems remains an open research problem.
-Many techniques focus on isolated predictions or modules and fail to capture system-level dynamics.
-Future work should prioritize longitudinal interpretability across agent interactions, causal reasoning pipelines, and interactive querying interfaces that support transparency in real time.
+The integration of trust-enabling mechanisms and explainability techniques is critical for the adoption of Agentic AI systems in high-stakes domains.
+While transparency, human oversight, and interpretable architectures improve user confidence, challenges in capturing system-level dynamics and scaling explainability remain.
+Future research should focus on developing standardized frameworks for multi-agent explainability, real-time interactive querying, and longitudinal trust evaluation to ensure robust and responsible deployment.
 A comparative summary of current techniques is shown in Table 6.
-7.3 Evaluation Metrics for Agentic AI Systems
+8 Evaluation of Agentic AI Systems
 Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy.
-We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI system performance and its real-world implications.
+We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an AMAS performance and its real-world implications.
 Below, we discuss these metrics and summarize in Table 7.
@@
 This criteria evaluates the reliability, safety, and ethical alignment of the AI agent.
-A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior.
+A trustworthy agent consistently produces correct and unbiased results, adheres to safety constraints, and avoids harmful or unpredictable behavior.
 Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [153].
 Some approaches combine such factors into an overall trust score.
-For example, one model defines a trustworthiness score $T$ as
+We can define a trustworthiness score $T$ for an Agentic AI system as :
 $T=\frac{C+R+I}{S}\,,$ (1)
@@
 Coordination.
-In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together.
+In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics measure how effectively these parts work together.
 Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently.
@@
 Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
-8 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
+9 Security and Privacy in Agentic AI Systems
 In this section, we discuss the security and privacy in AMAS.
@@
 Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [181].
-Among the foundational techniques employed are encryption [193, 194], access control [195, 171], adversarial defense [196, 197], and runtime monitoring[198, 199] each adapted to the unique demands of decentralized multi-agent environments.
+Among the foundational techniques employed are encryption , access control , adversarial defense , and runtime monitoring, each adapted to the unique demands of decentralized multi-agent environments.
 Encryption.
-plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [194, 61, 200].
+Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [61].
 Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts.
-Implementations such as SSL/TLS, homomorphic encryption [201, 200], and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
+Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
 Access control.
-becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
-For instance, in systems like AutoGen[18] and CrewAI [81] where agents take on specialized roles (e.g., summarizer, planner, coder) [202], enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [18, 198, 203].
-Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [204, 205] and Attribute-Based Access Control (ABAC) [206, 205] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels [5].
-Adversarial robustness.
-is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results [207, 208].
-Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others [208, 192, 115, 55, 116, 209].
-Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities [210, 115].
+Access conrol becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
+For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [53].
+Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [182] and Attribute-Based Access Control (ABAC) [182] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
+Adversarial learning.
+Adversarial learning is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results.
+Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially designing outputs from one compromised agent propagating misleading information to others [183].
+Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities.
 Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
 Runtime Monitoring.
-systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity [211].
-Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [212, 213, 125].
+Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity .
+Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [184].
 For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
-As Agentic AI systems continue to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount.
-Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
 9.2 Privacy-Preserving Techniques
-The decentralized and interactive nature of LLM-based Agentic AI systems introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory.
+The decentralized and interactive nature of AMAS introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory.
 To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
@@
 In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information.
-For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [214, 215].
+For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [185].
 DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
@@
 Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
-Privacy-by-design principles are becoming central to the engineering of next-generation agentic systems.
+Privacy-by-design [187] principles are becoming central to the engineering of next-generation Agentic AI systems.
 Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share.
 As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
-8.3 Compliance and Governance
-As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative.
-Unlike traditional AI agents, Agentic AI systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications.
-Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
-Regulatory standards.
-Standards for regulation provide the baseline requirements that all AI systems, including agentic architectures, must meet.
-Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness.
-These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation.
-For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring of such systems.
-Auditability.
-This is critical in ensuring transparency and facilitating post-hoc accountability.
-Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification.
-Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached.
-For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations.
-Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
-Policy enforcement.
-Policy governs what Agentic AI system can and cannot do, and under which conditions.
-These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions.
-Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level.
-Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent sub-components.
-Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
-Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the Agentic AI systems scale or change context.
-These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions.
+10 Compliance to AI Governance and Policy in Agentic AI Systems
+As Agentic AI systems grow more autonomous and capable, ensuring they remain compliant with regulations and governed by robust policies becomes imperative.
+Unlike traditional AI, these multi-agent or highly autonomous systems operate with persistent memory and complex decision-making loops, which introduces new risks around safety, ethics, and legality.
+Recent industry surveys indicate that enterprise leaders rank security, trust, compliance, and oversight as top concerns when deploying agentic AI at scale.
+To address these concerns, governance of agentic AI spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
+Below, we expand on each dimension and highlight emerging best practices for secure and compliant agentic AI deployments.
+10.1 Regulatory standards
+Regulatory frameworks provide baseline requirements that all AI systems – including agent-based architectures – must meet.
+Key frameworks outline principles of transparency, accountability, safety, and fairness which are especially pertinent for autonomous multi-agent systems operating in high-stakes domains (e.g.
+healthcare, finance, transportation).
+Notable examples include:
+  • NIST AI Risk Management Framework (AI RMF) [188], a voluntary U.S.
+framework emphasizing risk identification, mitigation, and trustworthy AI practices across the AI lifecycle.
+It calls for governance functions that ensure AI systems are transparent, accountable, and monitored for risks.
+Adopting the NIST AI RMF “Govern” and “Explain” functions helps organizations embed policies and documentation that align with trustworthiness goals.
+  • EU AI Act [106], a comprehensive EU regulation (set to fully apply in 2026) that classifies AI systems by risk level.
+High-risk AI systems (a category that would likely include certain autonomous or agentic AI in critical fields) are subject to strict requirements: rigorous risk management, human oversight, transparency, and continuous monitoring.
+For example, the Act explicitly mandates ongoing risk assessment and traceability for high-risk AI, and it requires that such systems be designed with human-in-the-loop or human supervision capabilities (per Article 14 on oversight).
+Agentic AI developers must therefore build in features for traceability and human override to remain compliant with these upcoming rules.
+  • Domain-Specific Regulations In addition to AI-focused standards, Agentic AI systems must comply with sectoral data and safety laws.
+For instance, an autonomous multi-agent healthcare assistant must adhere to health privacy laws like HIPAA [104], and any system handling personal data of EU citizens must follow GDPR [103] requirements for data protection by design.
+These regulations often require data minimization and user consent, which Agentic AI systems need to respect (e.g.
+by expiring or anonymizing stored user data to avoid retention of sensitive information).
+10.2 Auditability
+Auditability is critical for transparency and post-hoc accountability in agentic AI.
+Given the complexity of multi-agent decision-making (with potentially emergent or opaque interactions), organizations must be able to reconstruct and scrutinize the chain of events leading to any outcomes.
+Achieving robust auditability involves implementing the following practices:
+  • Comprehensive Logging: Each decision, action, plan, and inter-agent message should be logged with rich metadata – including timestamps, the initiating agent’s identity/role, the context or prompt, and a rationale or justification for the action.
+Such detailed logs create a decision provenance trail, enabling auditors to trace how and why a particular outcome was reached.
+  • Action Traceability: Techniques like decision provenance link each agent’s actions to preceding events or inputs, so one can follow the sequence of steps across agents.
+In complex agent workflows (e.g.
+a chain of a Planner agent delegating to a Coder agent, then to a Tester agent), traceability means every hand-off and decision node is recorded and can be reviewed in context.
+This helps identify where an error or policy violation might have occurred.
+  • Role-Specific Audit Trails: Modern agentic frameworks often assign specialized roles to different agents (for example, in AutoGen or MetaGPT, agents may take on roles like Researcher, Coder, or Reviewer to collaborate on tasks).
+The audit system should capture actions at the granularity of each role.
+By tagging log entries with the agent role, one can detect anomalies or biases in specific functions – for instance, flagging if a “Reviewer” agent consistently approves outputs that the “Researcher” agent flagged as doubtful, or if a certain role’s actions lead to security issues
+  • Immutable Logging: To ensure the integrity of audit logs (so that records cannot be tampered with after the fact), organizations are exploring append-only, tamper-proof logging mechanisms.
+For example, using blockchain or other immutable ledger technology to record multi-agent interactions can provide verifiable audit trails.
+This means any decision or message recorded in the system is cryptographically secured against alteration, which is valuable for compliance checks and forensic analysis in case of an incident.
+Robust auditability not only helps with internal accountability but is also increasingly expected by regulators.
+Transparent record-keeping allows external auditors or regulators to verify compliance with regulations and ethical standards.
+In the event of a problematic outcome, detailed audit logs ensure that responsibility can be determined and corrective action taken.
+In fact, governance frameworks stress that organizations should implement such audit trails and explicable decision logs so that any harmful or biased result can be traced back and attributed appropriately.
+Overall, auditability in agentic AI builds the foundation for trust – both for internal stakeholders and for users or regulators – by making the system’s behavior observable and explainable after the fact.
+10.3 Policy Enforcement
+While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds.
+Policy in this context refers to the codified rules and constraints that the AI agents must obey – covering operational limits, ethical guidelines, and security restrictions set by developers or governance boards.
+In practice, effective policy enforcement in agentic AI includes:
+  • Orchestrator-Level Governance: Agentic AI systems typically have an orchestration layer or a meta-controller that coordinates multiple agents.
+This layer should be programmed to enforce global policies on what agents can do and under what conditions.
+For example, a policy might dictate that any plan an agent comes up with must receive approval from a supervisory agent (or a human) if it involves deleting or altering important data.
+The orchestration logic would then require that checkpoint before execution, overriding agent autonomy when necessary.
+  • Memory and Data Retention Policies: To comply with privacy rules and avoid unchecked growth of an agent’s knowledge base, policies can enforce memory expiration or sanitization.
+Agents might be required to forget or redact certain information after a task is complete or after a time limit, preventing long-term retention of sensitive data.
+This guards against data misuse and helps comply with regulations like GDPR’s “data minimization” principle.
+For instance, if an agent system stores user queries or personal details, a policy could mandate purging those from the agent’s memory after fulfilling the query (unless explicitly allowed to retain).
+  • Tool and Resource Access Control: Agentic AI often involves agents calling external tools, APIs, or databases to accomplish tasks.
+Role-Based Access Control (RBAC) and Attribute-Based Access Control (ABAC) are crucial for limiting what each agent (or agent role) is permitted to do.
+Under RBAC, each agent role has permissions aligned with its responsibilities – e.g.
+a “Coder” agent might be allowed to call a code compilation service but not to fetch customer data; a “Data Analyst” agent might read a database but not invoke external network calls.
+ABAC can add context-sensitive rules (attributes like time of day, the specific project, or the agent’s trust level) to further restrict actions.
+  • Dynamic Sandboxing and Execution Constraints: Each agent’s actions (especially those that can change the environment or have side effects) should run in a sandbox or constrained context where possible.
+For example, if an agent is meant to write to a file, sandboxing can ensure it only writes to a designated directory and not to system-critical files.
+Similarly, agents might be prohibited from executing certain operations (like sending network requests to unapproved domains) through whitelists/blacklists coded into the environment.
+Dynamic sandboxing means these execution rules can adjust based on the agent’s role or current task, providing flexible yet controlled operation.
+  • Real-time Monitoring and Intervention: Even with policies encoded, agents might encounter novel situations.
+Real-time monitoring systems serve as a safety net – they watch agent behavior and can halt or flag any activity that deviates from predefined ethical constraints or operational bounds.
+For instance, if an agent’s output is about to violate a known policy (say, it’s about to generate disallowed content or leak PII), a monitoring module can catch this and either censor the output or stop the agent’s process.
+Some approaches involve using formal verification and runtime checking tools (like TLA+ specifications or symbolic execution) to prove certain unsafe actions will not occur, or to continuously verify that invariants hold during execution.
+10.4 Emerging Governance Practices
+Because the field of Agentic AI is evolving rapidly, governance strategies are also adapting.
+Beyond the foundational controls above, organizations are adopting new best practices to strengthen oversight and ensure long-term compliance as these AI systems scale.
+Some of the emerging practices include:
+  • AI Governance Boards: Many organizations are establishing dedicated governance boards or committees for AI oversight.
+These boards typically include stakeholders from legal, ethics, security, and domain experts.
+Their role is to continuously review the Agentic AI system’s objectives, assess risks, and update policies or strategies as needed.
+For example, a governance board might conduct regular audits of an AI’s decisions for bias or fairness, ensure alignment with the latest regulations, and approve any major changes to the system’s capabilities.
+This human layer of oversight creates accountability at the organizational level and sends a signal that the AI’s operations are being closely supervised by responsible parties.
+  • Governance-as-Code: Inspired by the infrastructure-as-code paradigm in DevOps, governance-as-code means encoding governance policies and ethical rules directly into the software and configuration that drive the AI system.
+Instead of governance being only a manual or paperwork process, it becomes part of the system’s logic.
+For Agentic AI, this could involve using policy engines or rule-based systems that automatically check each agent action against a set of coded rules.
+For instance, a rule repository might specify “Agent of type X cannot use tool Y without permission Z” and this is enforced by the orchestration code.
+  • Adaptive Governance Layers: Agentic AI systems may change over time – they might learn new skills, add more agents, or move into new deployment contexts.
+Adaptive governance refers to frameworks that evolve along with the AI system.
+This can include auto-updating rule sets or ML-driven policy adjustments.
+For example, if an agent starts operating in a new country, an adaptive governance layer might automatically incorporate that country’s AI regulations into the policy set.
+Or if a system update gives agents new capabilities, the governance rules adapt (or prompt human governors to adapt them) to cover those capabilities.
+  • Standards Alignment and Certification: To have trust, organizations are aligning their Agentic AI governance with emerging international standards.
+A prime example is ISO/IEC 42001:2023 [189], and ISO/IEC 42005:2025 [101] the management system standard, which specify requirements for effective AI governance and risk management .
+ISO 42001 and 42005 emphasizes many of the same principles discussed above – such as transparency, accountability, bias mitigation, safety, and privacy in AI development and deployment.
+Forward-looking organizations are therefore proactively building those capabilities (e.g.
+implementing rigorous risk management pipelines, bias audits, etc.) so that by the time new laws fully kick in, their agentic AI will be in natural compliance.
+By integrating these enforcement measures, Agentic AI systems can operate with much tighter alignment to intended guidelines and ethical norms.
+Policy enforcement translates high-level governance principles into day-to-day operational safeguards, reducing the chance of an autonomous agent straying outside its allowed domain or causing a compliance breach.
 Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
 11 Discussion
-We discuss the broader implications of our findings below:
-9.1 Technical Implications of TRiSM for LLM Agent Design
+In this section, we present our findings and outline directions for future work.
+11.1 Technical Implications of TRiSM for Agentic AI Design
 Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [190].
@@
 TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds.
-Likewise, emerging threats specific to Agentic AI systems such as prompt injection [23] attacks, memory poisoning [137], or cascading hallucinations [217] highlights the need for more built in risk controls.
+Likewise, emerging threats specific to Agentic AI systems such as prompt injection attacks, memory poisoning , or cascading hallucinations, highlights the need for more built in risk controls.
 By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g.
@@
 Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”.
-This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed [161].
+This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed .
 Recent guidance on trustworthy AI [161] often highlights accountability and explainability as key pillars of trust.
@@
 TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [192].
-In MAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly [126].
+In AMAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly .
 Therisk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [193].
@@
 These are precisely the capabilities that a TRiSM approach promotes.
-The EU AI Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM security management component is designed to ensure (through adversarial resilience, access control, etc.).
-Similarly, international AI governance standards are emerging that mirror TRiSM’s principles.
-ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development.
-By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 [89] and similar standards call for (e.g.
-leadership oversight, documented risk controls, ongoing monitoring and improvement cycles).
-Another example is the U.S.
-NIST AI Risk Management Framework [87], which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
+Deploying Agentic AI systems responsibly requires a multifaceted governance approach.
+By adhering to established regulatory standards (and keeping abreast of new ones), by ensuring auditability for transparency and accountability, and by enforcing policies through technical and organizational means, we can significantly mitigate the risks posed by autonomous AI agents.
+Moreover, embracing evolving best practices – from governance boards to adaptive, code-driven policies – will help maintain control over these systems as they grow in complexity.
+The end goal is to harness the benefits of Agentic AI (increased efficiency, creativity, and problem-solving at scale) while staying within legal, ethical, and safe bounds, thereby preserving public trust and meeting compliance obligations.
 11.4 Limitations and Current Research Gaps
 While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research.
-First, the classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles.
+First, the classical AI domains that works on standard test suites, there is no consensus on how to measure an AI Agent vs the Agentic AI ability to operate safely under TRiSM principles.
 This makes it difficult to compare different governance strategies or to track progress objectively.
 We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [194] or “red-teaming” exercises [195] that can stress-test Agentic AI systems and score their resilience.
-Secondly, there is a lack of real-world validation for many TRiSM-inspired controls.
+There is also a lack of real-world validation for many TRiSM-inspired controls.
 Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [25].
 It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
-Another critical gap is adversarial robustness.
+Another critical gap is adversarial learning.
 As we improve defenses, attackers will inevitably adapt [196].
 Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [19].
-TRiSM solutions today are not foolproof against these.
-We identify evolving threats and adversarial attacks as ongoing obstacles to trustworthy AI [40].
-Finally, organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures.
+Beyond technical aspects, the organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures.
 Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent.
@@
 11.5 Future Roadmap for Agentic AI TRiSM
-Table 9: Future Roadmap for Agentic AI: Eight Strategic Domains
-Scalable Agent Architectures Secure Multi-Agent Collaboration
-Explainable Multi-Agent Decisions Human-Centered Trust Design
-Autonomous Lifecycle Management Ethics & Governance
-Benchmarking & Evaluation Cognitive Capability Expansion
-Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure LABEL:fig:roadmap_mindmap.
+Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
+Figure 8: Agentic AI Roadmap: Mindmap Representation
+Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 8.
 These recommendations span both technical system design improvements and governance-level policy initiatives:
-Develop Standardized Evaluation Benchmarks.
-The community should create open benchmarks and challenge environments to test multi-agent AI governance.
-For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls.
+The community should create open benchmarks and challenge environments to test AMAS governance.
+For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed Agentic AI systemthat performs relative to one without such controls.
 This will enable direct comparisons and drive progress on measurable metrics of trust (e.g.
 frequency of prevented failures or fairness outcomes).
-Advance Adversarial Robustness Techniques.
 Future system design must anticipate a continually evolving threat landscape.
@@
 Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [198] , can help discover vulnerabilities in a controlled way before real adversaries do.
-Human-Centered Oversight Tools.
 We encourage designing better interfaces and protocols for human oversight of Agentic AI systems.
 Borrowing from human-computer interaction [199] and cognitive engineering [200], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
-Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials.
+Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials.
 These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators.
 Insights from such pilots can inform refinements in both technical standards and regulations.
-Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the start rather than retrofit them.
-Cross-Domain Best Practices and Ethical Governance: There is much to learn from other high-stakes domains.
+There is much to learn from other high-stakes domains.
 For example, the safety engineering field (e.g.
@@
 However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations.
-This review has systematically explored these challenges, proposing a TRiSM framework structured around four pillars: Governance, Explainability, ModelOps, and Privacy/Security, specifically adapted for LLM-based Agentic Multi-Agent Systems (AMAS).
+This review has explored these challenges, proposing a TRiSM framework structured around four pillars: Explainability, ModelOps, Security, Privacy and Governance, adapted for LLM-based AMAS.
 Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency.
 The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
-Looking forward, the proposed research roadmap emphasizes scalable, verifiable, and regulation-compliant Agentic AI systems.
 Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination.

### v4

@@
 We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents.
-We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Explainability, ModelOps, Security, Privacy and Governance, each contextualized to the challenges of multi-agent LLM systems.
-A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation.
+We then adapt and extend the AI TRiSM framework for Agentic AI, structured around key pillars: Explainability, ModelOps, Security, Privacy and their Lifecycle Governance, each contextualized to the challenges of AMAS.
+A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation.
 To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows.
 We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance.
-The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
+The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, highlighting key directions to align emerging systems with TRiSM principles-ensuring safety, transparency, and accountability in their operation.
 keywords:
@@
 1 Introduction
-1.1 Background and Motivation
-The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1].
-By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2].
-An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation.
-While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4].
-These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5].
-This rapid evolution presents new opportunities but also brings challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
-In particular, as the shift toward agentic AI architectures amplifies and evolves, longstanding concerns around safety, fairness, interpretability, and systemic harms resurface.
-While such issues have been central to machine learning (ML) systems broadly, they become more promiment in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
-From Isolated Agents to Agentic AI The autonomous software agents date back to the 1990s [6], however, the advent of LLM-driven multi-agent ecosystems [7] since 2023 has made “Agentic AI”, a distinctly new and rapidly expanding research focus [5].
-Traditional agents primarily focussed on automating limited tasks, such as information retrieval [8], data summarization [9], and dialogue response [10].
-Typically driven by single step logic [11] or scripted rules, these traditional agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving .
-In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), and enabled by LLMs and tool use [12].
-These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [13].
-This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
+AI governance has moved from aspiration to obligation.
+The EU Artificial Intelligence Act [1] entered into force on 1 August 2024 with a staged application (e.g., prohibitions in early 2025 and obligations for general-purpose AI from 2 August 2025  [^3] EU AI Act); meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [2] and the NIST AI Risk Management Framework (AI RMF 1.0) [3] to operationalize risk controls.
+In parallel, enterprise use of AI continues to expand: recent global surveys [4, 5, 6] report that a substantial majority of organizations use AI in at least one business function.
+These trends sharpen demand for system-level trust, risk, and security management tailored to agentic systems.
+Figure 1: TRiSM taxonomy for Agentic AI (AMAS) presented in this review.
+We cover: (i) Fundamentals of Agentic AI (Agentic vs.
+traditional agents; LLM-based architectures), (ii) AI Trust, Risk, and Security Management (TRiSM) pillars: Explainability, ModelOps, Security, Privacy, and Lifecycle Governance, (iii) Trust management via explainability, (iv) Threats and risks, and (v) Evaluation metrics, including human-centered and composite measures.
+An AI agent is a computational entity that perceives its environment and acts to achieve goals [7].
+What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [8].
+In this paper, we refer to such LLM-based, coordinating agent systems as Agentic multi-agent systems (AMAS).
+These systems orchestrate multiple specialized agents to address long-horizon tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents.
+The shift from single-agent or rule-based pipelines to AMAS unlocks new capabilities but complicates coordination, control, transparency, and risk management across heterogeneous components and contexts.
+Why TRiSM for Agentic AI?
+LLM-based AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [9].
+These risks are amplified when agents plan, coordinate, and call external tools.
+Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [10, 11].
+To address these system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks emphasize governance, explainability, security, privacy, and lifecycle controls [12, 13].
+We argue that a TRiSM perspective is necessary to make AMAS deployable in sensitive domains.
+Scope and Objectives
+Our focus is LLM-based AMAS that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning, with or without human oversight.
+[^4] Throughout, AMAS denotes multi-agent systems built on LLMs.
+We synthesize recent work on AMAS and AI TRiSM to provide a coherent, actionable reference for researchers and practitioners.
+Specifically, we map TRiSM pillars :Explainability, ModelOps, Security, Privacy, and their Lifecycle Governance, to AMAS workflows and failure modes.
+We also surface gaps where current ML governance and model-centric controls do not readily extend to multi-agent settings.
+Necessity of this Survey
+Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively under-addressing adversarial robustness, lifecycle governance, decision provenance, and system-level explainability.
+As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks.
+To our knowledge, few surveys examine LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance; Table 1 situates our contribution relative to prior reviews.
+Table 1: Comparison with related surveys.
+Survey Threats Lifecycle Gov. Explainability TRiSM Integration LLM-Specific Applications Actionable Guidance
+Guo et al. (2024) [14] ~
+Chen et al. (2025) [15] ~ ~
+Yan et al. (2025) [16] ~
+Tran et al. (2025) [17] ~
+Lin et al. (2025) [18] ~ ~
+Fang et al. (2025) [11] ~ ~ ~
+Xi et al. (2025) [19] ~
+Luo et al. (2025) [20] ~ ~ ~
+Zou et al. (2025) [21] ~ ~
+Wang et al. (2025) [22] ~ ~ ~ ~
+This Survey (2025)
+Legend: ✓= explicitly addressed; ~= partially addressed; ✗= not addressed.
+Contributions
+We make the following contributions:
+  • Unified TRiSM framework for Agentic AI.
+We propose and present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to large-language-model-based multi-agent systems (Section 5).
+The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualising each pillar to AMAS workflows and failure modes.
+  • Risk taxonomy for AMAS.
+We synthesise a taxonomy of threats unique to LLM-driven multi-agent systems, such as prompt injection, memory poisoning, collusive failure, emergent misbehaviour and tool-use abuse; and map these risks to corresponding controls (Section 4).
+  • Evaluation template and new metrics.
+We propose a holistic evaluation template that covers trustworthiness, explainability, user-centred performance and inter-agent coordination.
+To operationalise this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilisation Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 6).
+  • Technique mapping and limitations.
+We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting.
+This mapping reveals where existing approaches can be adapted and where significant gaps remain (Section 8).
+Building on the above contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems.
+Key directions include adversarial robustness, governance protocols, standardised benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
+2 Background, Related Work and Literature Review Methodology
+2.1 AI Agents vs. Agentic AI
+An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals [7], has undergone a rapid transformation in recent years.
+While early agents were task-specific and deterministic, modern systems have evolved substantially, increasing operational autonomy on benchmark tasks [23].
+These agentic systems integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [8].
+This evolution presents new opportunities but also raises challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
+As the shift toward agentic architectures accelerates, long-standing concerns around safety, fairness, interpretability, and systemic bias resurface [11], especially in high-stakes domains such as healthcare, law, and public services, where decisions can carry profound societal consequences.
+Autonomous software agents date back to the 1990s [24]; however, the advent of LLM-driven multi-agent ecosystems since 2023 [9] has made “Agentic AI” a distinct and rapidly expanding research focus [8].
+Traditional agents primarily automated limited tasks—such as information retrieval [25], data summarization [26], and dialogue response [27]—often via single-step logic [28] or scripted rules.
+These approaches lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem-solving.
+In contrast, agentic systems coordinate collaborative agents with specialized roles (e.g., planner, coder, analyst) enabled by LLMs and tool use [29].
+They dynamically decompose tasks, share context, and pursue high-level goals over long horizons [30].
+This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent, decentralized behavior.
 2.2 Trust, Risk, and Security Challenges in Agentic AI
-The rise of Agentic AI introduces new challenges that extend beyond those of traditional ML systems.
-For example, unlike deterministic agents [14], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
-These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
-Below, we discuss some of these challanges:
+The rise of Agentic AI introduces challenges that extend beyond those of traditional ML systems.
+Unlike deterministic agents [31], LLM-based multi-agent systems (AMAS) exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
+These characteristics necessitate robust approaches to building trust and ensuring trustworthiness in high-stakes domains such as healthcare, finance, and law.
+Below, we discuss several key challenges:
   1. 1.
-    Trust and Trustworthiness In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust.
+    Trust and trustworthiness.
+In Agentic AI, trust refers to a user’s willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner.
+According to ISO [32], trustworthiness is a quality characteristic of systems and is understood as the ‘ability to meet stakeholders’ expectations in a verifiable way’.
 This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications.
-The stochastic nature of LLM reasoning [15] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability.
-These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
+The stochastic nature of LLMs [33] further complicates trustworthiness: randomness in output generation, while sometimes promoting creativity, introduces inconsistency and limits repeatability.
   2. 2.
-    Security Risks in Multi-Agent Systems Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [16].
-For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate.
-The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
+    Security risks in multi-agent systems.
+Risk surfaces expand as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [10].
+For example, a multi-agent supply-chain optimizer may coordinate procurement and logistics agents yet inadvertently leak sensitive data or violate compliance protocols (e.g., [34], HIPAA [35]) if safeguards are inadequate.
   3. 3.
-    Specific Threats: Prompt Injection, Spoofing, and Impersonation LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors.
-One prominent threat is prompt injection [17], where attackers design inputs with hidden or malicious instructions to manipulate agent behavior.
-Recent studies [18] describe “prompt infection” as a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [19].
-Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
+    Prompt injection, spoofing, and impersonation.
+AMAS are susceptible to security threats that exploit language-based interfaces and cooperative behaviors.
+One prominent threat is prompt injection [36], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior.
+Recent studies [37] describe “prompt infection”, in which malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [38].
+Corrupted outputs can cascade, leading to data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
     Another critical vulnerability involves spoofing and impersonation [39].
-Spoofing means faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges.
-In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
+Spoofing refers to falsifying identity or credentials (e.g., using a false name or address) to deceive a system, whereas impersonation involves pretending to be a specific agent or user to gain trust, access, or privileges.
+In AMAS, where agents rely on credentials or tokens for authentication during coordination, an adversary who steals an API key or mimics a trusted peer can issue unauthorized commands or access sensitive information.
   4. 4.
-    Privacy Risks in Multi-Agent LLM Systems LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [21].
-In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g.
-via a shared memory store or message passing) to collaborate.
-Traditional evaluation and safety frameworks, designed for deterministic agents AI systems, are inadequate for addressing the autonomy and control of Agentic AI.
-To address these gaps, frameworks like AI TRiSM (Artificial Intelligence Trust, Risk, and Security Management) [22, 23] proposes lifecycle-level controls, including explainability, secure model orchestration, and privacy management.
-AI TRiSM is a framework to help organizations build AI systems that people can trust.
-The goal is to focus not just on technical performance, but also on ethical values like fairness, transparency, and security.
-The framework brings together practices that support governance, strengthen reliability and robustness, and protect data and make sure that AI systems are not only effective, but also aligned with human values.
-Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review.
-It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs.
-Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
-1.3 Problem Statement
-Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks.
-However, no unified framework exists for managing the trust, risk, and security across dynamic as well as multi-agent workflows.
-Recent accidents, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [24].
-These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways.
-To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is therefore urgently needed.
-1.4 Scope and Objectives
-Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [22] framework to LLM-powered, multi-agent “Agentic AI” systems.
-We structure the discussion around four foundational pillars of TRiSM [25]: explainability, security, lifecycle governance, and privacy.
-The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
-For the purposes of this study, we define Agentic AI as Multi-Agent Systems (MAS) [26] powered by LLMs that exhibit autonomous planning, tool use, memory retention, and emergent reasoning capabilities, with or without human supervision.
-These LLM-based agents can dynamically adapt workflows based on interactions with tools, users, and other agents, distinguishing them from traditional symbolic or rule-based agents.
-Scope This paper is a conceptual survey and synthesis.
-It does not propose a new architecture for Agentic AI, nor does it include synthetic benchmark evaluations.
-Instead, it integrates current perspectives on AI TRiSM to inform future research and the responsible development of Agentic AI systems.
-The scope of this paper is confined to LLM based Agentic AI systems and AMAS  [^3] Throughout this paper, AMAS designates multi-agent systems built on LLMs.
-Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability.
-As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks.
-This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
-To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security.
-Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
-1.5 Contributions
-Our key contributions in this review are:
-  • We propose the first TRiSM framework for Agentic AI systems.
-This framework is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate.
-The framework is organized around pillars: Explainability, ModelOps, Security, Privacy and Governance around autonomous decision to guide the development and use of safe and ethical Agentic AI systems.
-(see Fig.
-1 and Sec.
-6).
-  • We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec.
-5).
-This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats commonly seen in AMAS.
-  • We introduce an evaluation dimensions for Agentic AI systems that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec.
-8).
-To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
-Our aim is to achieve high performance in both dimensions, accuracy and usability, without sacrificing one.
-  • We contribute to the management and control of Agentic AI by mapping existing TRiSM techniques such as explainability methods, prompt hygiene, decision provenance, sandboxing, and ModelOps pipelines, to AMAS environments.
-We critically analyze their limitations and adaptation needs, providing a foundation for re-engineering these techniques to fit the demands of interacting LLM agents (Secs.
-79).
-  • Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 11).
-2 Related Work
-Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
-[…]
-Guo et al. (2024) [27] ✓ (simulated env’ts) ~ (research challenges)
-Chen et al. (2025) [28] ~ (task/simulation focus) ~ (future research areas)
-Yan et al. (2025) [29] ✓ (mentions diverse) ~ (future work directions)
-Tran et al. (2025) [30] ✓ (networks, QA, etc.) ~ (open challenges)
-Lin et al. (2025) [31] ~ (creative tasks) ~ (roadmap for research)
-Fang et al. (2025) [32] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
-Xi et al. (2025) [33] ✓ (single-/multi-agent,
-human-AI coop)
-~ (open problems)
-Luo et al. (2025) [34] ~ ~ ✓ (science, games, etc.) ~ (open challenges)
-Zou et al. (2025) [35] ~ ✓ (e.g., finance,
-healthcare)
-~ (challenges & opp.)
-Wang et al. (2025) [36] ~ ~ ~ (healthcare domain) ~ (future opportunities)
-This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
-Legend: ✓= explicitly addressed; ~= partially/indirectly addressed; ✗= not addressed.
-[…]
-We present the related work in this section.
+    Privacy risks in multi-agent LLM systems.
+LLM-based agents often process user data, proprietary business information, and other sensitive inputs to fulfill their tasks [40].
+In a multi-agent context, this challenge is amplified by inter-agent information sharing (e.g., via a shared memory store or message passing), which complicates data minimization, access control, and purpose limitation.
+Traditional evaluation and safety frameworks, designed for deterministic, single-agent pipelines; are inadequate for the autonomy and control properties of AMAS.
+To address these gaps, AI Trust, Risk, and Security Management (AI TRiSM) [12, 13] proposes lifecycle controls spanning explainability, secure model/tool orchestration, and privacy management.
+Rather than focusing solely on technical performance, TRiSM integrates governance practices to strengthen reliability and robustness, protect data, and align system behavior with ethical and regulatory expectations.
+2.3 Survey Landscape
 LLM-based multi-agent surveys (technical focus)
-Prior surveys on LLM-driven AMAS predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI.
-However, there remains a gap in addressing governance and policy aspects.
-For example, reviews like [27] and [28] delve into simulated environments, inter-agent communication protocols, and performance benchmarks.
-These works mainly detail how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
-More specialized surveys, such as [29], focus on natural language communication, while [30] details collaboration mechanisms, exploring interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0.
-Although [29] highlights security risks associated with communication channels, and [30] discusses deployment contexts, both remain narrowly centered on coordination strategies.
-There is a need for deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability challenges at scale.
+Prior surveys on LLM-driven AMAS predominantly address system architectures, agent capabilities, and domain-specific applications of agentic systems.
+For example, [14, 15] examine simulated environments, inter-agent communication protocols, and performance benchmarks.
+These works clarify how agents coordinate and scale but give limited treatment to adversarial threats (e.g., prompt injection, collusion), lifecycle governance (e.g., continuous monitoring, audit, regulatory compliance), and system-wide explainability, especially for emergent behaviors in complex settings.
+More specialized surveys, such as [16], focus on natural-language communication, while [17] details collaboration mechanisms and practical use cases in domains like question answering, 5G networks, and Industry 5.0.
+Although [16] highlights security risks associated with communication channels and [17] discusses deployment contexts, both remain narrowly centered on coordination strategies.
+Deeper analysis is needed on robustness to adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability at scale.
 Trustworthy/responsible AI surveys (broad trust focus)
-There are some review articles that address TRiSM in AI, such as [32], tackle ethical and risk-related issues including alignment with human values, discuss fairness in decision-making [37], and defenses against privacy attacks (e.g., data leakage).
-However, these works typically target general ML systems and overlook the unique dynamics of LLM-based multi-agent systems.
-For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems.
-Moreover, critical TRiSM aspects [38] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent in recent discussions.
-Current Study (Multi-Agent Systems and LLMs)
-This review distinguishes itself by mapping TRiSM principles into the study of AMAS.
-Unlike prior studies (see Table 1), which offer limited insight into governance and policy, this work operationalizes TRiSM principles in practice.
-It provides actionable guidance for researchers, engineers, and policymakers, addressing critical gaps in both the technical AMAS literature and the broader responsible AI landscape.
-3 Literature Review Methodology
-To review the literature related to TRiSM as applied in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [39, 40].
-This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
+Review articles addressing TRiSM-related themes (e.g., [11]; see also fairness-focused work [41]) discuss alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage).
+However, these surveys typically target general ML systems and overlook dynamics unique to LLM-based multi-agent settings.
+For instance, while they advocate privacy-preserving techniques such as differential privacy, they rarely address privacy risks stemming from inter-agent data sharing or the distributed nature of LLM-driven workflows.
+Moreover, critical TRiSM aspects—such as governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents), are often under-specified [42].
+2.3.1 Positioning of this study
+This review maps TRiSM principles to LLM-based AMAS, operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings.
+Relative to prior surveys (Table 1), we foreground system-level risks (e.g., prompt infection, impersonation, memory poisoning), propose evaluation dimensions and composite measures for coordination and tool use, and offer actionable guidance for researchers, engineers, and policymakers.
+2.4 Literature Review Methodology
+We conducted a structured literature review to synthesize AI TRiSM principles for AMAS.
+Guided by established systematic-review practices [43, 44], this section states our research objectives and describes the data sources and search strategy, the inclusion/exclusion criteria and screening procedure, and the classification schema used for analysis.
 (a) Agentic AI papers per year (2022–2025).
-(b) TRiSM pillars papers by topic.
-Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
-Research Objectives The review was guided by the following four specific questions:
-  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems?
-How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
-  • RQ2: What technical and governance strategies have been proposed to address these challenges?
-  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
-  • RQ4: What gaps remain in current research, and what directions are promising for future work?
-Search Strategy To conduct a review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
-The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context.
-We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations.
-The search terms were structured as follows:
-(“ Agentic AI” OR “multi-agent systems” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
-AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
-Inclusion and Exclusion Criteria We applied the following criteria during the screening process:
-Inclusion Criteria:
-  • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
-  • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
-  • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
-Exclusion Criteria:
-  • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
-  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) without reference to agentic behavior or TRiSM concerns.
-  • Non-English papers or those lacking sufficient metadata (e.g., missing abstracts or publication details).
-After screening titles and abstracts, we shortlisted 250 papers.
-A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars.
-In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework.
-An analysis on the literature is shown in Figure 2.
-Quality Assurance To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [41].
-Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems?
-(2) Are the methods, experiments, or architectural designs well-documented and reproducible?
-(3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects?
-(4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance?
-Each criterion was rated as low, medium, or high.
-Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only.
-Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
+(b) Papers by TRiSM pillar.
+(c) Topic trends in Agentic AI (agents, security, explainability).
+(d) Bibliography entries per year (1987–2025).
+Figure 2: Bibliographic analysis of Agentic AI research.
+(a) Annual trend in Agentic AI publications, showing a sharp rise post-2023.
+(b) Distribution of TRiSM-related publications across governance and safety topics.
+(c) Topic-specific trends in agents, security, and explainability.
+(d) Overall growth in bibliographic entries, highlighting acceleration since 2020.
+Research questions (RQs)
+The review was guided by the following three questions:
+  • RQ1: What trust, risk, and security challenges are specific to LLM-based AMAS?
+  • RQ2: Which governance, explainability, security, privacy, and lifecycle (TRiSM) controls are proposed or adapted for AMAS?
+  • RQ3: How are AMAS evaluated (datasets, tasks, metrics), and where are the gaps?
+Search strategy
+We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
+The search covered publications from January 2022 to May 2025, with select seminal works prior to 2022 included for foundational context.
+We used Boolean queries and expanded keywords to capture the intersection of LLM-based multi-agent systems and TRiSM considerations.
+Example query template:
+("Agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR "distributed AI" OR "LLM-based agents" OR "agent coordination" OR "agent-based modeling" OR "swarm intelligence")
+AND
+("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection" OR "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR "system reliability")
+Screening and eligibility
+Two-stage screening (title/abstract, then full-text) used these criteria. Inclusion:
+  • Studies published 2022–2025 (capturing the post-2022 LLM wave), plus select pre-2022 foundational work.
+  • Studies explicitly discussing Agentic AI, multi-agent, or LLM-powered systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
+  • Peer-reviewed articles and widely cited preprints or reports from credible sources (e.g., NeurIPS, IEEE, ACM, Nature, governmental/standards bodies, arXiv).
+Exclusion:
+  • Papers focused solely on traditional, rule-based, or symbolic agents without LLM integration or emergent coordination.
+  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) with no agentic behavior or TRiSM concerns.
+  • Non-English papers or those lacking sufficient metadata (e.g., missing abstract or publication details).
+For empirical papers, we assessed clarity of threat models, reproducibility (code/data), evaluation rigor (baselines, ablations), and external validity.
+For surveys, we assessed coverage breadth, structure, and treatment of risks/governance.
+After screening titles and abstracts, we shortlisted 250 papers; full-text review yielded 180 primary studies addressing one or more TRiSM pillars.
+We also incorporated relevant technical reports and authoritative web resources to contextualize the rapidly evolving Agentic-AI/TRiSM landscape.
+A bibliometric summary appears in Figure 2.
+Quality assurance
+To ensure reliability, we adapted quality criteria from established guidelines [45].
+Each paper was evaluated on: (1) clear objectives for agentic or LLM-based systems; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to Agentic-AI governance.
+Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only.
+Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion (with inter-rater agreement checked on a subset).
 3 Fundamentals of Agentic AI Systems
@@
 3.1 Traditional AI Agents vs Agentic AI
-Traditional agents operate through predefined rules [42], heuristic workflows [43, 44], or deterministic logic [45], and perform well in narrow, well-defined environments.
-Cognitive Agents [46], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [47] or ACT-R [48].
-However, they remain constrained by static knowledge bases and limited adaptability [46].
+Traditional agents operate through predefined rules [46], heuristic workflows [47], or deterministic logic [48], and perform well in narrow, well-defined environments.
+Cognitive Agents [49], model human-like cognitive functions, such as integrating perception, memory, and decision-making, often using cognitive systems architectures like [50] or ACT-R [51].
+According to a recent work [52], authors distinguish between 4 types of knowledge: Concepts, Skills, Processes, Motives for cognitive agents.
+A key difference between Agentic AI and such cognitive architecture approaches is that the latter were mostly applied to robot planning and control (so applies to ‘physical operational environments’) but Agentic AI applies more broadly, also to non-physical operational environments (which could be made to operate entities in physical also environments).
 In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
@@
 A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
-4.2 LLM-Based Agentic AI System Architecture
-Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks.
+For the scope of this work, we rely on LLM-based Agentic AI systems.
+3.2 Multi-Agentic AI System (AMAS) Architecture
+AMAS represents an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks.
 As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem.
@@
 The modular design highlights adaptive intelligence and inter-agent synergy.
-Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results.
+As shown in Figure 3), communication and coordination are facilitated through a Communication Middleware [61] that allows agents to share goals, observations, or intermediate results.
 A Task Manager or Orchestrator [62] governs high-level planning and delegates subtasks to agents based on their roles or specializations.
@@
 Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
-Framework (Year) Core LLM Planning Memory Tool Use Notable Features
+Acronyms: CoT = Chain-of-Thought reasoning, FSM = Finite State Machine, PDDL = Planning Domain Definition Language, DAG = Directed Acyclic Graph, DB = Database, REPL = Read–Eval–Print Loop, API = Application Programming Interface, KV = Key–Value store, CLI = Command Line Interface, SDK = Software Development Kit, MCP = Model Context Protocol.
+Framework Core LLM Planning Memory Tool Use Notable Features
 AutoGPT [76] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
@@
 SuperAGI [90] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
+Semantic Kernel [91] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
+OpenAI Swarm [92] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
+OpenAI Agents SDK [93] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
+Strands Agents [94] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
+LlamaIndex Agents [95] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
 Tool-Use Interface.
@@
 This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
-Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
-For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
-5 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
+Table  LABEL:tab:agentic_architectures presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
+4 Threats and Risks in LLM-based Agentic Multi-Agent Systems
 In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
@@
 4.1 Threats in Agentic AI
-Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [87]. Below, we discuss some of these major threats.
+Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures. Below, we discuss some of these major threats.
   • Autonomy abuse.
@@
 This risk involves central or distributed control mechanisms for role assignment and workflow mediation.
-A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT and AutoGen .
+A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [53] and AutoGen [98] .
 These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
@@
 These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
-5.3 Real-World Examples
+Real-World Examples
 Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
@@
 This led to hallucinated goals and emergent behaviors misaligned with initial intentions [109].
-Next, we discuss the AI TRiSM prism framework and map its core principles to a AMAS.
-6 The TRiSM Framework for Agentic AI
+Next, we discuss the AI TRiSM framework and map its core principles to a AMAS and how it can mitigate these aforementioned risks.
+5 The TRiSM Framework for Agentic Multi-Agent Systems
 5.1 Overview of AI TRiSM and Its Core Components
@@
 In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
-Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
-Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
-Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [99]
-ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
-Metrics: fidelity, sufficiency, human-trust score
-ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
-SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
-Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [102]
-Jailbreak-success%; mean-time-to-detect; exploit CVSS
-Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [103]
-HIPAA §164 [104]
-Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
   • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness.
@@
 Originally highlighted in industry guidelines for AI governance [110], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems.
-The TRiSM framework addresses trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 23].
+The TRiSM framework addresses trust concerns in Agentic AI by focusing on key pillars, which are: Explainability, ModelOps, Application Security, Model Privacy and Goverance [64, 13].
 Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
-Figure 5: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems.
-The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers.
-External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
+Figure 5: TRiSM-aligned architecture for Agentic AI systems.
+The system is organized into four layers: Trust and Explainability (e.g., reasoning traces, user feedback), ModelOps (prompt versioning, CI/CD, orchestration), Planning and Reasoning (multi-agent CoT), and Model Privacy (encrypted memory, differential privacy, audit logs).
+A cross-cutting Governance Plane enforces oversight via policies and logging.
+External tools are accessed through a secure API gateway.
+The user interface connects back to the trust and governance layers for human-in-the-loop (HITL) oversight.
+Abbreviations used: TRiSM = Trust, Risk, and Security Management; CoT = Chain-of-Thought; CI/CD = Continuous Integration and Deployment; DP = Differential Privacy.
+Table 4: TRiSM pillars for Agentic (LLM-based) multi-agent systems: core controls, techniques, risks, evaluation facets, example systems, and governing frameworks.
+Pillar Core controls (keywords) Techniques / patterns / artifacts Key risks mitigated Evaluation facets & example metrics Example systems / patterns Primary standards / frameworks
+Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [111]; LIME/SHAP [112, 113]; decision-provenance graphs [114, 115]; multi-agent SHAP [116]; attention maps [117, 118]; role-based interpretability [119]; RAG-linked justifications [120]; prompt/action logs [121, 122] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [123] SciAgent [120]; MetaGPT [124]; ChatDev [30]; AutoGen [57] EU AI Act Arts 13–14 [125]; NIST AI RMF / GAI profile [3, 126]; OECD AI [127]; ISO/IEC 24029-1 [32]
+ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [128] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent (hierarchical monitoring) [128] ISO/IEC 42001 (AIMS) [ISO_IEC_42001_2023]; ISO/IEC 42005 [129]; ISO/IEC 23894 [iso23894 or [130]]; NIST AI RMF / GAI profile [3, 126]; ModelOps [131, 132]
+Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [133, 58]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [134]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [135]; enterprise dashboards (override/stop) [135]; ChemCrow safety pauses [136] OWASP LLM Top-10 [137, 138]; OECD Robustness [127]
+Model Privacy & Data Protection DP; anonymization/pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [139]; k-anonymity [140]; MPC/PUMA [puma2024 or [141]]; HE/FHE pipelines [142]; TEEs [143]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [34]; CCPA/CPRA [144]; HIPAA §164 [35]; ISO/IEC 27001/27701 [iso27001, iso27701]
+Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [125]; NIST AI RMF [3]; OECD AI [127]; ISO/IEC 42001/42005 [align keys]; ISO/IEC 23894 [130]
 5.2 Mapping Pillars of TRiSM with Agentic AI
-Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
-13–14) [106] and the NIST AI RMF “Explain” function [107].
-This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [108].
-ModelOps : Anchored in the NIST RMF “Govern” function [107] and ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101].
-It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates.
-Application Security: Guided by OWASP LLM Top-10 [102] and OECD AI Principles on Robustness [109], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms.
-Model Privacy: Aligned with GDPR (Art.
-25) [103], HIPAA §164 [104], and OECD AI Principles [109], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
-Governance: Grounded in overarching regulatory frameworks such as GDPR, CCPA [110], and the OECD AI Principles [109], this pillar supports institutional oversight, policy compliance, and cross-system accountability.
-It enforces auditable workflows, verifiable role-based access, regulatory traceability, and human oversight, particularly in multi-agent and high-risk AI deployments.
-Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, 42005, and OECD Principles.
-Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
-Explainability in Multi-Agent Decision Making
-Explainability refers to making the inner workings and decisions of AI agents interpretable to humans.
-In the context of LLM-based multi-agent systems, it is paramount for building user trust because outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [111].
-However, achieving explainability in such systems is particularly challenging , one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to the final outcomes [112].
-In other words, explaining why a multi-agent system behaved a certain way requires tracing which agent did what and how their collaboration produced an emergent result, a task that traditional single-model explainability techniques do not directly address.
-Several approaches are being explored to enhance explainability in LLM-driven multi-agent systems:
-  • Feature Attribution (LIME/ SHAP): Established explainable AI techniques provide a starting point.
-For instance, LIME) [113] and Shapley Additive Explanations (SHAP) [114] can be adapted to analyze LLM decisions.
-These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken.
-In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
-By highlighting the most influential parts of each agent’s input, methods like LIME/SHAP help clarify individual agent decisions in the multi-agent mix [115].
-  • Counterfactual Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability.
+In this section, we examine how the pillars of TRiSM align with Agentic AI (or AMAS) and summarize our syntheis in Table LABEL:tab:trism-pillars.
+For each pillar, we outline its relevance and implications within the context of Agentic AI.
+5.2.1 Explainability in Multi-Agent Decision Making
+Explainability refers to making the inner workings and decisions of AI systems interpretable to humans [145].
+In AMAS settings, outcomes emerge from the interactions among specialized agents rather than a single model prediction [146].
+Therefore, explainability must cover both individual agent decisions and the inter-agent dynamics that lead to emergent results.
+Recent research emphasises that user trust hinges on understanding how each agent contributes and how their collaboration is orchestrated [111].
+Traditional single-model XAI techniques often ignore this multi-agent dimension, so new methods are required.
+Challenges Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information they used, and how their interactions influenced the final outcome [56].
+Additional challenges include: (1) Emergent behaviour: interactions among agents can lead to behaviour that was not explicitly programmed; explaining these emergent patterns requires capturing interaction chains.; (2) Verification of intermediate steps: vanilla CoT prompting in LLMs often produces plausible-sounding but unchecked reasoning.
+Without verification, small errors can propagate
+Recent directions Recent works introduce new frameworks and techniques for multi-agentic explainability:
+  • Layered Chain-of-Thought prompting: Instead of producing a single monolithic reasoning trace, Layered-CoT divides reasoning into discrete layers [111].
+Each layer answers a sub-question, and a specialized Reasoning Agent generates the partial chain-of-thought while Verification Agents and User-Interaction Agents cross-check the intermediate steps against knowledge graphs or human feedback.
+This agent-based decomposition localizes errors, prevents contradictions, and ensures that each partial solution is validated before the model proceeds.
+  • Integrated multi-agent workflows: Multi-agent workflows embed explainability into the process itself rather than as a post-hoc add-on.
+For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves relevant DSM-5 criteria and explicitly links conversational utterances to diagnostic rules.
+This integration produces transparent, step-by-step diagnostic predictions and shows why specific recommendations are made [147].
+The study notes that step-by-step reasoning improves perceptions of fairness and reliability and that embedding a diagnostician agent moves beyond general synthetic dialogues by providing a self-documenting rationale.
+  • Modular resume-screening systems: In recruitment, a multi-agent architecture divides the process into extraction, evaluation, summarization and formatting.
+The evaluation agent dynamically retrieves company-specific hiring criteria via Retrieval-Augmented Generation (RAG), and each stage remains clearly defined.
+This modularity allows recruiters to trace how a candidate was assessed and why a particular score was assigned [148].
+  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide individualized credit assignments to agents and explain its feedback strategy.
+Unlike earlier deep RL methods, this approach uses the language-based critic to justify why each agent receives a particular reward and to show how it decomposes the global reward [149].
+Approaches for explainability in multi-agent LLM systems The following techniques adapt traditional XAI methods to the multi-agent setting:
+  • Feature Attribution (LIME/ SHAP) for agents: Techniques like LIME [112] or SHAP [113] can be extended to attribute an agent’s output to specific input components (e.g., particular context tokens or inter-agent messages).
+Highlighting which parts of the input or which inter-agent signals most influenced an agent’s action helps illuminate why a specific agent behaved as it did [150].
+  • Counterfactual and Causal Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability.
 Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [151].
@@
   • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency.
-Logging the step-by-step reasoning process – for example, an agent’s chain-of-thought (CoT) prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [69].
+Logging the step-by-step reasoning process – for example, an agent CoT prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [75].
 By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision.
 This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
-  • Natural Language Explanations: Finally, the generative nature of LLMs can be leveraged to turn the above insights and logs into human-friendly explanations [117].
-Instead of expecting end-users to read raw CoT logs or analyze feature weights, an “explainer” agent (which could be another LLM) or a human analyst can compile the reasoning trace into a narrative.
-In fact, an important aspect of explainability is communicating the reasons behind decisions in an intelligible way.
-By using natural language, the system can answer user questions like “Why did the agents decide to do X?” in a conversational manner.
-ModelOps : Lifecycle Management for LLM Agents
-ModelOps (Machine Learning Operations) is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [118].
-Within AMAS, ModelOps covers not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59].
-Effective ModelOps helps maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
-AMAS require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations.
-A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations.
-Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified.
-Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
-Application Security for LLM Agents
-To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents.
-This includes prompt hygiene [19] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions.
-Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [119].
-It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [120].
-Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
-Recent frameworks, such as LangChain [121], AutoGen [122], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
-Moreover, training LLM agents with adversarial robustness in mind (e.g.
-fine-tuning on adversarial examples, employing adversarial training regimes) [19] can improve their resistance to malicious inputs.
-Model Privacy and Data Protection
-The model privacy addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [21].
-Since multiple agents may share information with each other (e.g.
-via a shared memory store or message passing) to collaborate.
-Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs.
-To preserve privacy in AMAS, several technical measures can be employed.
-For example, Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [123].
-Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [124].
-Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [125].
-Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [126].Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [127].
-Governance
-Governance of AI systems necessitates compliance with evolving legal and policy frameworks, including the General Data Protection Regulation (GDPR) [103] and the California Consumer Privacy Act (CCPA) [110].
-Effective governance extends beyond regulatory adherence to cover proactive management of legal, ethical, and societal risks.
-This includes establishing transparent accountability structures, maintaining auditable documentation, and ensuring the protection of data subject rights throughout the AI lifecycle.
-Several global frameworks provide guiding principles that align with the TRiSM pillars.
-The EU AI Act [106], particularly Article 14, mandates transparency, traceability, and human oversight in high-risk systems; principles embedded within TRiSM’s Explainability and ModelOps dimensions.
-Similarly, the NIST AI Risk Management Framework (AI RMF) [107] identifies: Explainability, Governability, and Validity as key components of AI risk mitigation, which directly correspond to TRiSM’s lifecycle control mechanisms.
-The OECD AI Principles [109] reinforce the importance of robustness, transparency, and accountability, echoing TRiSM’s attention to secure tool invocation, memory traceability, and multi-agent auditability.
-Together, these frameworks underscore the need for governance architectures in Agentic AI systems that support real-time monitoring, role-based access, verifiable logs, and intelligible outputs.
-As illustrated in Figure 6, TRiSM offers a practical scaffolding for operationalizing these requirements in LLM-based multi-agent environments.
-7 Trust and Explainability in Agentic AI Systems
-Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
-Mechanism Purpose Example System(s) Trust Contribution
-Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [128], MetaGPT [129] Enhances interpretability and debuggability
-Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [13], AutoGen [122] Mimics human team transparency; reduces uncertainty
-Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [130], enterprise dashboards Prevents unsafe execution; reinforces control
-Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [131, 132] Improves predictability; prevents overreach
-Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [133, 134] Enhances user comfort and trust in uncertain scenarios
-Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms.
-Figure 7 illustrates the layered relationship between trust mechanisms, explainability techniques, and evaluation facets.
-Trust-enabling mechanisms, such as transparency and human-in-the-loop (HIIL) oversight, provide the foundation for user confidence.
-These feed into explainability techniques like decision provenance graphs, which make agent behaviors interpretable.
-Finally, evaluation (Section 8) facets ensure that trust and explainability are quantifiable, enabling iterative improvements in system design.
-7.1 Building Trust in Autonomous Agents
-In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [135].
-Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [136].
-Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
-Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
-Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
-Transparency and Decision Traceability.
-Transparency is one of the core enablers of trust.
-For users to understand and evaluate AMAS decisions; the reasoning chains, decision states, and action triggers of agents must be made observable.
-Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules.
-For example, SciAgent [128] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG).
-Similarly, MetaGPT[129] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
-Status Reporting and Progress Visibility.
-Clear reporting of intent and intermediate status is also essential.
-Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions.
-Tools like AutoGen[122] and ChatDev[13] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
-Human-in-the-Loop Oversight.
-Human oversight and intervention mechanisms further reinforce trust.
-Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [137, 138, 139].
-Many systems adopt a human-in-the-loop (HITL) paradigm [140], where agents request confirmation before executing high-risk or irreversible actions.
-Behavioral Consistency and Bounded Autonomy.
-Trust requires predictability in models’ behaviour.
-Agents should follow defined roles, output in expected formats, and remain within delegated authority.
-For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material.
-Similarly, in autonomous research assistants like ChemCrow [130], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
-Social Trust Cues and Language Behaviors.
-Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial.
-If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it.
-Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [141, 12, 142].
-For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries .
-Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [143].
-Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [134].
-This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability.
-These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
-7.2 Explainability Techniques in AMAS
-Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system.
-In platforms such as ChatDev [13] or AutoGen [53], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial.
-While local explainability techniques such as LIME [113] and SHAP [114] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory.
-Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [144], causal influence chains [145], and multi-agent SHAP adaptations [146].
-These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
-Symbolic and Hybrid Architectures.
-Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5].
-These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding.
-Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
-Lightweight Interpretability: Attention and Prompts.
-Attention map visualizations have been used to highlight focus areas in multi-modal language agents [147, 148, 149], offering lightweight but informative views into model behavior.
-Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [150, 151].
-These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
-Table 6: Explainability Techniques in Agentic AI Systems
-Technique Purpose Challenges in Agentic AI Example Use Cases
-LIME [113] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
-SHAP [114] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
-Decision Provenance Graphs [144, 152] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
-Multi-agent SHAP [146] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
-Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
-Attention Maps [147, 149] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
-Prompt Audit Trails [150, 151] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
-The integration of trust-enabling mechanisms and explainability techniques is critical for the adoption of Agentic AI systems in high-stakes domains.
-While transparency, human oversight, and interpretable architectures improve user confidence, challenges in capturing system-level dynamics and scaling explainability remain.
-Future research should focus on developing standardized frameworks for multi-agent explainability, real-time interactive querying, and longitudinal trust evaluation to ensure robust and responsible deployment.
-A comparative summary of current techniques is shown in Table 6.
+  • Natural Language Explanations: Multi-agent systems can include a separate explainer agent that synthesizes logs and feature attributions into human-readable narratives.
+This agent (often another LLM) translates the technical reasoning traces into conversational explanations that answer questions like “Why did the agents choose action X?” Integrated workflows, such as DSM5AgentFlow [147], demonstrate that embedding such explanation directly in the pipeline improves transparency and trust.
+  • Role-oriented modular architectures: Assigning clear roles to agents (e.g., reasoning, verification, summarization, credit assignment) enhances explainability because each stage of decision-making is distinct and can be inspected independently [148].
+Modular designs also allow domain experts to adjust or replace specific agents without affecting the entire system.
+  • Explainable reward critics: In multi-agent reinforcement learning, LLM-based critics can decompose global rewards and provide natural-language justifications, making credit assignment transparent [149] .
+This approach helps developers understand how credit is distributed across agents and why certain behaviors are encouraged or discouraged.
+5.2.2 ModelOps : Lifecycle Management for LLM Agents
+ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management and continuous oversight for all types of AI models [131].
+Unlike MLOps, ModelOps manages not only individual models but also composite systems of agents and heterogeneous models across different business units.
+ModelOps provides end-to-end oversight for AI in production by enforcing model governance and compliance: tracking lineage, ownership, and metadata to meet regulations such as GDPR and HIPAA , and by using policy-driven deployment gates so only approved models go live, with controlled roll-outs and roll-backs [132].
+It continuously monitors risk and bias, detecting drift, fairness issues, hallucinations, and robustness regressions.
+ModelOps also manages heterogeneous assets: rule-based systems, LLMs, and composite multi-agent architectures; ensuring consistent visibility and lifecycle control across the portfolio.
+Finally, it integrates tightly with enterprise IT, wiring AI workflows into CI/CD, IT service management (ITSM), identity and access management (IAM), and cloud-security tooling [132].
+For generative AI and AMAS, ModelOps must handle prompt versioning, guardrails and content filtering enforcement, fine-tuning lineage tracking, usage and cost observability, and ethical compliance.
+Effective ModelOps thus requires:
+  • Version control and CI/CD pipelines: Track versions of each agent’s model and prompt configuration.
+Automated pipelines should test safety, performance and cost impacts whenever an agent’s logic or model is updated, using multi-agent simulations to validate that new behaviours do not introduce regressions or unsafe interactions.
+  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [128] decompose tasks into subtasks and dynamically form groups of agents.
+They employ hierarchical monitoring, where each agent tracks its actions with a checklist, group-level administrators oversee their agents, and system-level supervisors review outputs.
+This layered monitoring ensures that changes to one agent do not compromise the entire system.
+  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks and updating measurement approaches regularly; it suggests using gradient-based attributions, occlusion and counterfactual prompts to improve explainability and transparency, and establishing policies for oversight across the lifecycle [126].
+5.2.3 Application Security for LLM Agents
+Security in AMAS systems must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools.
+According to OWASP [137, 138], prompt injection manipulates an LLM through malicious instructions, potentially causing the model to leak data, execute unauthorized commands or make harmful decisions.
+Multimodal models face cross-modal injection where malicious content in one modality influences the output of another.
+To mitigate these risks:
+  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure prefixes, and validate instructions before they reach an agent.
+Constraining the output format and employing guardrails can limit the model’s response surface [38].
+  • Defense-in-depth design patterns: Existing techniques such as adversarial training and user confirmation provide partial protection, but research shows they do not guarantee safety.
+Design patterns [133] like Action-Selector and Plan-Then-Execute restrict an agent’s capabilities after processing untrusted input; the agent selects from a predefined list of allowed actions or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [58].
+  • Least-privilege access and authentication: Enforce strong authentication for both human users and agents and apply role-based access controls so agents can only invoke the minimum set of tools necessary to perform their task.
+Continuous monitoring should detect anomalous requests; if an agent deviates from normal patterns, automated systems should flag or suspend it [152].
+  • Post-processing and anomaly detection: Implement filters to detect and remove sensitive or inaccurate content from outputs and regularly retrain models to reduce hallucinations.
+Maintain robust logging and retention policies so that investigators can audit interactions without exposing private data.
+Recent frameworks also emphasis hierarchical monitoring and cross-checking among agents to detect inconsistencies [128].
+LangChain [153], AutoGen [98], and CrewAI [87] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
+Training agents with adversarial examples [134] and red-teaming [154] scenarios further strengthens robustness against malicious inputs.
+5.2.4 Model Privacy and Data Protection
+AMAS also poses unique privacy challenges because agents often share information via memory stores and message passing.
+Without strict controls, an agent may inadvertently leak sensitive data [40].
+A layered privacy strategy should include:
+  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re-identified and reduces the risk that models memorize or regurgitate personal information [139].
+DP should be assessed on a case-by-case basis, since heavy noise can harm accuracy.
+  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [140].
+Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
+  • Secure Multi-Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs.
+In privacy-preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7-billion-parameter model within minutes) [141].
+  • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy.
+Recent work demonstrates an encryption-friendly LLM architecture that uses LoRA fine-tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [142].
+  • Trusted Execution Environments (TEEs): TEEs provide hardware-enforced isolation that protects code and data [143].
+They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone.
+In secure LLM inference, TEEs ensure that models run only inside authenticated enclaves, with attestation verifying model integrity.
+  • Continuous monitoring and access controls: Encrypt data at rest and in transit, restrict access to authorized personnel, and use monitoring to detect data poisoning or unusual behaviour.
+These techniques, combined with careful consent management and compliance with local regulations, help maintain confidentiality when agents share or process sensitive information.
+Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST AI RMF, ISO/IEC 42001, 42005, and OECD Principles.
+Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
+Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
+13–14) [125] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [3]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [155].
+Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [119].
+ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [3] and ISO/IEC 42001:2023 (AI management system) [156]; secure SDLC controls via NIST SP 800-218/218A [^5] NIST, 2024.
+Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates.
+Risk Management & Impact Assessment: EU AI Act (Art.
+9, 72–73) [125], ISO/IEC 23894 (AI risk management) [^6] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [129].
+Iterative hazard analysis, impact registers, incident response, and disclosure playbooks.
+Data Governance & Quality: EU AI Act (Art.
+10) [125] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^7] ISO/IEC 5259 series.
+Data minimization, lineage, coverage/bias checks, and sampling documentation.
+Robustness & Safety: EU AI Act (Art.
+15) [125] and ISO/IEC 24029-1 (robustness assessment) [32].
+Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria.
+Application Security: Guided by OWASP LLM Top-10 [137] and OECD AI Principles on robustness [127].
+Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails.
+Model Privacy: Aligned with GDPR (Art.
+25) [34], HIPAA §164 [35], and OECD AI Principles [127].
+Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds.
+Documentation & Traceability: EU AI Act (Arts.
+11–12; Annex IV) [125].
+Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages.
+Post-Market Monitoring: EU AI Act (Arts.
+72–73) [125].
+Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications.
+Governance: Grounded in GDPR, CCPA [144], and OECD AI Principles [127].
+Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments.
+Fairness & Human Rights: OECD AI Principles [127] and ISO/IEC TR 24027 (bias) [^8] ISO/IEC TR 24027.
+Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
+5.2.5 Governance
+Governance integrates ethical, legal and societal considerations into the design and operation of AI systems.
+Regulations such as the EU AI Act [125], NIST AI RMF [3] and OECD AI [127] provide guidance that aligns with TRiSM pillars of Explainability, ModelOps, Security and Privacy.
+  • EU AI Act - Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety or fundamental rights [1].
+The AI system should enable overseers to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop its operation through a ‘stop’ button.
+For certain identification tasks, decisions must be independently verified by at least two competent individuals.
+  • OECD AI Principles: These intergovernmental principles, updated in 2024, promote human-centric, trustworthy AI [127].
+They emphasise transparency and explainability: AI actors should provide meaningful information about data sources, factors and logic to allow affected individuals to understand and challenge outputs.
+Robustness, security and safety require AI systems to function appropriately under normal and adverse conditions and include mechanisms to override or decommission systems to prevent harm.
+Accountability means AI actors are responsible for the proper functioning of AI systems; they must ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
+  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile outline four core functions: govern, map, measure and manage [3].
+For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (gradient-based attributions, occlusion, counterfactual prompts) to improve transparency, and establishing policies for oversight across the AI lifecycle.
+  • Auditability and role-based access: Agentic AI systems should maintain auditable logs of agent actions, decisions and data flows and assign clear roles (e.g., reasoning, verification, summarisation) so that responsibilities are transparent [157].
+Role-based access controls help enforce accountability and prevent unauthorized agents from accessing sensitive functions.
+  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data subject rights, consent and privacy by design.
+Frameworks like the OECD Principles and NIST AI RMF promote international cooperation and interoperability.
+These governance frameworks (as illustrated in Figure 6) underscore the need for transparent accountability structures, human oversight, robust privacy controls, and documented processes in AMAS.
+TRiSM provides a practical scaffolding for operationalizing these requirements, aligning lifecycle management, security, privacy and governance to build trustworthy Agentic AI.
+Next, we discuss the evaluations in AMAS.
 6 Evaluation of Agentic AI Systems
-Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy.
-We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an AMAS performance and its real-world implications.
-Below, we discuss these metrics and summarize in Table 7.
+Agentic AI systems demand evaluation beyond traditional accuracy.
+We outline five categories: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct facet of performance and real-world impact.
+Compared to a related work [158], which focuses on proportional impacts across general AI, our work targets Agentic AI and links risks to their causal sources (e.g., prompt injection, memory poisoning).
+We further introduce an evaluation template with new metrics the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE) to capture multi-agent dynamics.
+A summary appears in Table 5.
 Trustworthiness.
-This criteria evaluates the reliability, safety, and ethical alignment of the AI agent.
-A trustworthy agent consistently produces correct and unbiased results, adheres to safety constraints, and avoids harmful or unpredictable behavior.
-Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [153].
-Some approaches combine such factors into an overall trust score.
-We can define a trustworthiness score $T$ for an Agentic AI system as :
-$T=\frac{C+R+I}{S}\,,$ (1)
-where $C$ is the agent’s credibility (accuracy and correctness of its outputs), $R$ is reliability (consistency of performance over time), $I$ is the level of user alignment or rapport (analogous to “intimacy” in trust modeling), and $S$ is self-orientation (the degree to which the agent pursues its own goals over the user’s goals).
-A higher $T$ indicates an agent that is accurate, consistent, user-aligned, and not self-serving, which corresponds to greater trustworthiness.
-In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
+This criterion evaluates reliability, safety, and ethical alignment [159].
+Core signals include task success rates under distributional shifts (robustness), violation rate of safety/ethical rules (lower is better), and calibration quality (e.g., Expected Calibration Error, ECE).
+To avoid pathological behavior in scoring, we define a normalized trustworthiness index
+$T\;=\;\frac{w_{C}\,C+w_{R}\,R+w_{I}\,I}{\,1+\lambda\,S\,}\,$, (1)
+where $C$ (credibility/accuracy), $R$ (reliability over time), and $I$ (user alignment/rapport) are normalized to $[0,1]$, $S\!\geq\!0$ is self-orientation (higher is worse), $w_{C},w_{R},w_{I}\!\geq\!0$ with $w_{C}{+}w_{R}{+}w_{I}\!=\!1$, and $\lambda\!\geq\!0$ controls how strongly self-orientation penalizes trust.
 Explainability.
-Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [154].
-These metrics focus on the clarity and completeness of the rationale the agent provides for its actions.
-For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
-Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process.
-Quantitatively, methods like OpenXAI [155] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations.
-High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process.
-In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
+Explainability metrics assess how well human evaluators can understand and trace decisions.
+Useful measures include coverage (fraction of outputs with an explanation), fidelity/faithfulness (agreement with the model’s actual decision basis), stability (consistency across similar cases), and human interpretability ratings.
+Suites like OpenXAI [160] report faithfulness, stability, and fairness of explanations.
+High explainability builds user trust and aids debugging; in regulated domains (e.g., healthcare, finance) it is often essential for compliance.
 User-Centered Performance.
-User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs.
-These criteria emphasize the user’s experience and outcomes [156].
-Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
-Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation.
-Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions.
-Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
-Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
+User-centered metrics capture how effectively the agent satisfies end-user needs [161].
+Signals include post-interaction satisfaction scores, goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness.
+Human-in-the-loop (HITL) studies typically rate helpfulness, clarity, and instruction adherence.
+Ultimately, a user-centered agent aligns actions with user intent and preferences.
+Table 5: Summary of key evaluation metrics for agentic AI systems. Each category addresses a distinct facet of performance and behavior.
 Metric Aspect Evaluation Focus (Examples of Metrics)
-Trustworthiness Reliability and safety of the agent’s behavior.
-Example metrics: task success rate across diverse scenarios; frequency of rule or safety violations (lower is better); calibration of confidence vs.
-outcomes; fairness and bias indices.
-Explainability Transparency and interpretability of the agent’s decisions.
-Example metrics: percentage of outputs with an accompanying explanation; explanation fidelity to actual model reasoning; consistency of explanations for similar cases; human interpretability ratings of explanations.
-User-Centered User satisfaction and alignment with user needs.
-Example metrics: user satisfaction score (post-interaction survey); rate of fulfilling user-defined goals; number of clarification questions needed (lower is better); naturalness and coherence of dialog from the user’s perspective.
-Coordination Effective collaboration in multi-agent or modular systems.
-Example metrics: multi-agent task completion rate; communication overhead (e.g., messages exchanged); consistency of shared plans or beliefs among agents; synergy score quantifying complementary actions.
-Composite Overall performance across metrics.
-Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
+Trustworthiness Reliability, safety, and alignment.
+Examples: success rate under shifts; safety/ethics violation rate (lower is better); calibration (ECE/Brier  [^9] https://en.wikipedia.org/wiki/Brier_score); fairness indices; red-team/jailbreak success (lower is better).
+Explainability Transparency and traceability. Examples: explanation coverage; faithfulness/fidelity; stability across similar cases; human interpretability scores; explanation time/cost.
+User-Centered Experience and outcomes.
+Examples: satisfaction (Likert/CSAT  [^10] https://www.ibm.com/think/topics/csat-customer-satisfaction-score); goal-completion rate; clarification turns; perceived coherence/naturalness; task time; cognitive load (NASA-TLX).
+Coordination Collaboration in multi-agent or modular setups.
+Examples: team task completion; coordination overhead (messages/tokens, rounds to converge); plan consistency; deadlock/conflict rate (lower is better); synergy score for complementary actions.
+Composite Holistic summaries. Examples: weighted aggregate across categories (weights sum to 1); Tool-Use Efficacy (TUE, see definition); cross-domain benchmark aggregates.
 Coordination.
-In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics measure how effectively these parts work together.
-Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently.
-Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
-One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis).
-MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
-Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings.
-For instance, if one agent generates a plan and another executes it, a coordination metric would assess if the executing agent follows the planner’s intent correctly and whether both agents remain in agreement throughout the process.
-High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
+In multi-agent or modular systems, coordination captures whether components maintain a shared plan, minimize redundant work, and resolve dependencies.
+Quantify with team success rate, communication efficiency (messages/tokens and rounds to consensus), plan/belief consistency, deadlock/conflict rate, and a Component Synergy Score (CSS) that credits inter-agent actions which enable or improve peers’ outcomes (higher is better).
 Composite Metrics.
-Composite metrics aggregate multiple evaluation aspects into a single overall score.
-These are useful for summarizing agents’ performance holistically, especially when comparing different systems.
-A composite metric is often a weighted combination of the above categories, for example:
+Composite metrics are often presented as a weighted scalar summary across dimensions:
 $M_{\text{composite}}\;=\;w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,$, (2)
-where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application.
-The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
-An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure.
-By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems.
-For instance, AgentBench [157] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent.
-Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others.
-It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability).
-Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
+where $M_{T},M_{E},M_{U},M_{C}\in[0,1]$ are normalized category scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}\!\geq\!0$ with $\sum w_{\bullet}=1$ encode domain priorities.
+While this provides a scalar index for ease of interpretation, it risks masking over-performance in one dimension with under-performance in another.
+To address this, we treat the evaluation more rigorously as a vector of metrics,
+$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{C})\,$, (3)
+and recommend selection based on $\arg\max$ over $\mathbf{M}$ rather than on any scalarization:
+$\arg\max_{i}\;\mathbf{M}_{i}\,.$ (4)
+In this view, Eq.
+2 can be interpreted as a limiting case, useful for communicating results to non-technical stakeholders, while Eq.
+3 and Eq.
+4 form the principled basis for evaluation and comparison.
+Definition (Tool-Use Efficacy, TUE). In this work, we introduce Tool-Use Efficacy (TUE) as a specialized composite for evaluating how correctly and efficiently an agent uses external tools. Let
+[…]
+[…]
+$\mathrm{Exec}$ $\in[0,1]$ $\text{(execution success: no exceptions/timeouts, correct API usage)},$
+$\mathrm{Out}$ $\in[0,1]$ $\text{(post-call outcome correctness on tool-gated sub-tasks)},$
+$\mathrm{Eff}$ $\in[0,1]$ $\text{(efficiency: normalized inverse cost for latency/tokens/\#calls)}.$
+We define two aggregators (both reportable; choose per application):
+(a) Linear aggregator (auditable weights).
+$\mathrm{TUE}_{\mathrm{lin}}\;=\;\alpha_{\!s}\,\mathrm{Sel}+\alpha_{\!a}\,\mathrm{Arg}+\alpha_{\!e}\,\mathrm{Exec}+\alpha_{\!o}\,\mathrm{Out}+\alpha_{\!f}\,\mathrm{Eff}\,,\quad\alpha_{\!*}\!\geq\!0,\;\sum\alpha_{\!*}=1.$ (5)
+(b) Geometric aggregator (“all-of-the-above” strictness).
+$\mathrm{TUE}_{\mathrm{geo}}\;=\;\Big{(}\mathrm{Sel}^{\beta_{s}}\,\mathrm{Arg}^{\beta_{a}}\,\mathrm{Exec}^{\beta_{e}}\,\mathrm{Out}^{\beta_{o}}\,\mathrm{Eff}^{\beta_{f}}\Big{)}^{\!1/(\beta_{s}{+}\beta_{a}{+}\beta_{e}{+}\beta_{o}{+}\beta_{f})},\quad\beta_{\!*}\!>\!0.$ (6)
+Overhead penalty. If the preference is a separate cost term, then we can define normalized overhead $\mathrm{OH}\in[0,1]$ (increasing in time/tokens/calls) and report the penalized score
+$\mathrm{TUE}^{\star}\;=\;\frac{\mathrm{TUE}}{\,1+\lambda\,\mathrm{OH}\,}\,,\quad\lambda\!\geq\!0,$ (7)
+where $\mathrm{TUE}$ is either (5) or (6). It is recommended to always report the components (Sel, Arg, Exec, Out, Eff) with confidence intervals to ensure reproducibility.
 7 Security and Privacy in Agentic AI Systems
 In this section, we discuss the security and privacy in AMAS.
-Table 8: Governance Dimensions for Agentic AI Systems
-Governance Component Description Example Tools or Methods
-Regulatory Compliance Aligning with legal and industry standards for responsible AI [158] NIST AI RMF [99], EU AI Act [159], GDPR [103], HIPAA [104]
-Auditability and Logging Capturing traceable records of agent actions and decisions [160] Decision provenance, immutable logs, blockchain-based traceability
-Policy Enforcement Enforcing system-level operational, ethical, and security constraints [52, 161] RBAC/ABAC policies [162], formal logic rules, dynamic sandboxing [60, 163, 164]
-Human Oversight Human-in-the-loop [140] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
-Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [165] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
-Explainability Governance Ensuring interpretability standards are met in decision pipelines [166] LIME/SHAP [167, 112, 168], decision trees [169, 170], interpretable surrogate models [171, 172, 173]
-Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [133, 135] Governance-as-code [174, 175], auto-updated rule engines [176, 177, 178], learning-based policy adaptation [90, 179]
-Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
 7.1 Security Mechanisms
@@
 Access control.
-Access conrol becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
+Access conrol is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
 For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [57].
@@
 Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution.
-For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data.
+For instance, temporary memory buffers used in systems like ChatDev [30] or ReAct-based pipelines [75] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data.
 Furthermore, anonymization and pseudony [167] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
 Secure computation.
-Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
+Techniques including secure multi-party computation (SMPC) [168], homomorphic encryption [169], and trusted execution environments (TEEs) [^11] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
 In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage.
@@
 As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
+Next, we discuss the compliance to AI goverance and policy in Agentic AI systems.
+Table 6: Governance controls for agentic AI systems: pillar $\rightarrow$ control $\rightarrow$ operationalization $\rightarrow$ evidence $\rightarrow$ standards.
+Pillar Control Operationalization (How) Evidence / Artifacts Standards & Articles
+Governance Org.
+accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [171]; ISO/IEC 42001 [2]
+Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts.
+11–14 [125]; ISO/IEC 42001 [2]
+Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [129]; EU AI Act Art.
+9 [125]
+Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts.
+72–73 [125]; NIST AI RMF (Manage) [171]
+Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [34]; ISO/IEC 5259 series
+Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [125]; ISO/IEC TR 24027
+Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [155]; NIST AI RMF (Map/Measure) [171]
+Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [125]
+Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts.
+12–13 [125]; ISO/IEC 42001 [2]
+Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art.
+14 [125]; NIST AI RMF (Manage) [171]
+Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art.
+15 [125]; ISO/IEC 24029-1 [32]
+Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [137]; HIPAA §164 [35]
+Security Tool access control RBAC/ABAC per agent/role; per-tool API tokens; argument validation; allowlists Access matrices; token scopes; arg-validator tests NIST SP 800-218/218A; NIST AI RMF (Manage)
+Security Isolation/sandboxing Containers/VMs; syscall/network/file egress policies; hardware TEEs for sensitive code Sandbox configs; egress policy; attestation logs NIST SP 800-218A; TEEs best-practice notes
+Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [34]; OECD principles [127]
+Supply Chain Provenance & SBOM Track sources and hashes for models/tools/datasets; dependency scanning SBOM; third-party attestations; license checks NIST SSDF (SP 800-218/218A); ISO/IEC 42001
+Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [171]; ISO/IEC 42001
+Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [32]
+Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [137]; NIST AI RMF (Manage)
+Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [127]; EU AI Act transparency
+Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [171]; ISO/IEC TR 24027
+Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [125]; ISO/IEC 42001
+GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [125]; ISO/IEC 42001
+Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art.
+52 [125]; C2PA121212c2pa.org
+Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [34]
+Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [34]
+Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [171]
+Supply Chain Third-party/vendor risk Vendor questionnaires; DPAs; penetration/attestation reports; license compliance Vendor risk register; DPA file; attestation/SBOM NIST SSDF (SP 800-218/218A); ISO/IEC 42001
+Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210131313ISO 9241-210; OECD principles [127]
+Sustainability Energy/CO2 telemetry Track/normalize energy use and CO2; report per release; set budgets Energy/CO2 logs; budget vs. actual; disclosure note Org policy / reporting guidance
 8 Compliance to AI Governance and Policy in Agentic AI Systems
-As Agentic AI systems grow more autonomous and capable, ensuring they remain compliant with regulations and governed by robust policies becomes imperative.
-Unlike traditional AI, these multi-agent or highly autonomous systems operate with persistent memory and complex decision-making loops, which introduces new risks around safety, ethics, and legality.
-Recent industry surveys indicate that enterprise leaders rank security, trust, compliance, and oversight as top concerns when deploying agentic AI at scale.
-To address these concerns, governance of agentic AI spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
-Below, we expand on each dimension and highlight emerging best practices for secure and compliant agentic AI deployments.
+Recent industry surveys indicate that enterprise leaders deploying Agentic AI at scale rank security, trust, compliance, and oversight among their top concerns.
+A survey [^14] TechRadar (2024) of IT professionals found that while 98% plan to expand AI agent usage, 96% view them as growing security threats, with only 54% having full visibility into their data access.
+In parallel, 67% of business leaders are increasing budgets for security oversight [^15] Cybersecurity Dive (2024), and 53% specifically cite data privacy and compliance as their primary challenge in scaling Agentic AI [^16] Cloudera (2024).
+Table LABEL:tab:agentic_governance operationalizes the governance controls into concrete implementations and evidence and next we detail the compliance to AI goverace and policy in Agentic AI systems.
 8.1 Regulatory Standards
-Regulatory frameworks provide baseline requirements that all AI systems – including agent-based architectures – must meet.
-Key frameworks outline principles of transparency, accountability, safety, and fairness which are especially pertinent for autonomous multi-agent systems operating in high-stakes domains (e.g.
-healthcare, finance, transportation).
-Notable examples include:
-  • NIST AI Risk Management Framework (AI RMF) [188], a voluntary U.S.
-framework emphasizing risk identification, mitigation, and trustworthy AI practices across the AI lifecycle.
-It calls for governance functions that ensure AI systems are transparent, accountable, and monitored for risks.
-Adopting the NIST AI RMF “Govern” and “Explain” functions helps organizations embed policies and documentation that align with trustworthiness goals.
-  • EU AI Act [106], a comprehensive EU regulation (set to fully apply in 2026) that classifies AI systems by risk level.
-High-risk AI systems (a category that would likely include certain autonomous or agentic AI in critical fields) are subject to strict requirements: rigorous risk management, human oversight, transparency, and continuous monitoring.
-For example, the Act explicitly mandates ongoing risk assessment and traceability for high-risk AI, and it requires that such systems be designed with human-in-the-loop or human supervision capabilities (per Article 14 on oversight).
-Agentic AI developers must therefore build in features for traceability and human override to remain compliant with these upcoming rules.
-  • Domain-Specific Regulations In addition to AI-focused standards, Agentic AI systems must comply with sectoral data and safety laws.
-For instance, an autonomous multi-agent healthcare assistant must adhere to health privacy laws like HIPAA [104], and any system handling personal data of EU citizens must follow GDPR [103] requirements for data protection by design.
-These regulations often require data minimization and user consent, which Agentic AI systems need to respect (e.g.
-by expiring or anonymizing stored user data to avoid retention of sensitive information).
+Regulatory frameworks provide baseline requirements that all AI systems, including agent-based architectures, must meet. Key frameworks include:
+  • NIST AI Risk Management Framework (AI RMF) [171].
+A voluntary U.S.
+framework organized around the Govern, Map, Measure, Manage functions.
+It emphasizes organizational governance, risk identification and measurement, and operational risk treatments.
+Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
+  • EU AI Act [125].
+A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026.
+High-risk systems require ongoing risk assessment, technical documentation (Annex IV), logging/traceability (Arts.
+11–12), data governance (Art.
+10), human oversight (Art.
+14), and robustness/accuracy (Art.
+15).
+  • ISO/IEC management and assessment standards.
+ISO/IEC 42001:2023 (AI management systems) sets organization-level governance requirements; ISO/IEC 42005:2025 (AI impact assessment) provides system-level impact-assessment guidance; ISO/IEC 23894 addresses AI risk management; ISO/IEC 24029-1 supports robustness evaluation; ISO/IEC TR 24027 addresses bias.
+These align well with AI RMF and the EU AI Act’s lifecycle duties.
+  • Domain-specific laws.
+Sectoral/privacy regimes still apply: GDPR (e.g., data protection by design, Art.
+25), HIPAA (45 CFR §164), and regional data-residency rules.
+Agentic systems should enforce data minimization, purpose limitation, and consent bases; consider DPIA for high-risk processing and AI impact assessments per ISO/IEC 42005 in parallel.
 8.2 Auditability
-Auditability is critical for transparency and post-hoc accountability in agentic AI.
-Given the complexity of multi-agent decision-making (with potentially emergent or opaque interactions), organizations must be able to reconstruct and scrutinize the chain of events leading to any outcomes.
-Achieving robust auditability involves implementing the following practices:
-  • Comprehensive Logging: Each decision, action, plan, and inter-agent message should be logged with rich metadata – including timestamps, the initiating agent’s identity/role, the context or prompt, and a rationale or justification for the action.
-Such detailed logs create a decision provenance trail, enabling auditors to trace how and why a particular outcome was reached.
-  • Action Traceability: Techniques like decision provenance link each agent’s actions to preceding events or inputs, so one can follow the sequence of steps across agents.
-In complex agent workflows (e.g.
-a chain of a Planner agent delegating to a Coder agent, then to a Tester agent), traceability means every hand-off and decision node is recorded and can be reviewed in context.
-This helps identify where an error or policy violation might have occurred.
-  • Role-Specific Audit Trails: Modern agentic frameworks often assign specialized roles to different agents (for example, in AutoGen or MetaGPT, agents may take on roles like Researcher, Coder, or Reviewer to collaborate on tasks).
-The audit system should capture actions at the granularity of each role.
-By tagging log entries with the agent role, one can detect anomalies or biases in specific functions – for instance, flagging if a “Reviewer” agent consistently approves outputs that the “Researcher” agent flagged as doubtful, or if a certain role’s actions lead to security issues
-  • Immutable Logging: To ensure the integrity of audit logs (so that records cannot be tampered with after the fact), organizations are exploring append-only, tamper-proof logging mechanisms.
-For example, using blockchain or other immutable ledger technology to record multi-agent interactions can provide verifiable audit trails.
-This means any decision or message recorded in the system is cryptographically secured against alteration, which is valuable for compliance checks and forensic analysis in case of an incident.
-Robust auditability not only helps with internal accountability but is also increasingly expected by regulators.
-Transparent record-keeping allows external auditors or regulators to verify compliance with regulations and ethical standards.
-In the event of a problematic outcome, detailed audit logs ensure that responsibility can be determined and corrective action taken.
-In fact, governance frameworks stress that organizations should implement such audit trails and explicable decision logs so that any harmful or biased result can be traced back and attributed appropriately.
-Overall, auditability in agentic AI builds the foundation for trust – both for internal stakeholders and for users or regulators – by making the system’s behavior observable and explainable after the fact.
+Multi-agent systems can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
+  • Comprehensive logging.
+Record prompts/contexts, plans, actions, tool calls, inputs/outputs, timestamps, and agent role/identity; capture rationales where feasible.
+Preserve hashes and version IDs for models, tools, and data snapshots.
+[…]
+  • Role-granular trails. Tag entries by agent role (e.g., Researcher, Coder, Reviewer) to localize anomalies and bias. Require reviewer countersignatures for high-impact actions.
+  • Immutability. Use append-only, tamper-evident storage (e.g., WORM buckets, cryptographic sealing) for audit logs; rotate keys and attest integrity regularly.
 8.3 Policy Enforcement
 While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds.
-Policy in this context refers to the codified rules and constraints that the AI agents must obey – covering operational limits, ethical guidelines, and security restrictions set by developers or governance boards.
+Policy in this context refers to the codified rules and constraints that the AI agents must obey.
 In practice, effective policy enforcement in agentic AI includes:
-  • Orchestrator-Level Governance: Agentic AI systems typically have an orchestration layer or a meta-controller that coordinates multiple agents.
-This layer should be programmed to enforce global policies on what agents can do and under what conditions.
-For example, a policy might dictate that any plan an agent comes up with must receive approval from a supervisory agent (or a human) if it involves deleting or altering important data.
-The orchestration logic would then require that checkpoint before execution, overriding agent autonomy when necessary.
-  • Memory and Data Retention Policies: To comply with privacy rules and avoid unchecked growth of an agent’s knowledge base, policies can enforce memory expiration or sanitization.
-Agents might be required to forget or redact certain information after a task is complete or after a time limit, preventing long-term retention of sensitive data.
-This guards against data misuse and helps comply with regulations like GDPR’s “data minimization” principle.
-For instance, if an agent system stores user queries or personal details, a policy could mandate purging those from the agent’s memory after fulfilling the query (unless explicitly allowed to retain).
-  • Tool and Resource Access Control: Agentic AI often involves agents calling external tools, APIs, or databases to accomplish tasks.
-Role-Based Access Control (RBAC) and Attribute-Based Access Control (ABAC) are crucial for limiting what each agent (or agent role) is permitted to do.
-Under RBAC, each agent role has permissions aligned with its responsibilities – e.g.
-a “Coder” agent might be allowed to call a code compilation service but not to fetch customer data; a “Data Analyst” agent might read a database but not invoke external network calls.
-ABAC can add context-sensitive rules (attributes like time of day, the specific project, or the agent’s trust level) to further restrict actions.
-  • Dynamic Sandboxing and Execution Constraints: Each agent’s actions (especially those that can change the environment or have side effects) should run in a sandbox or constrained context where possible.
-For example, if an agent is meant to write to a file, sandboxing can ensure it only writes to a designated directory and not to system-critical files.
-Similarly, agents might be prohibited from executing certain operations (like sending network requests to unapproved domains) through whitelists/blacklists coded into the environment.
-Dynamic sandboxing means these execution rules can adjust based on the agent’s role or current task, providing flexible yet controlled operation.
-  • Real-time Monitoring and Intervention: Even with policies encoded, agents might encounter novel situations.
-Real-time monitoring systems serve as a safety net – they watch agent behavior and can halt or flag any activity that deviates from predefined ethical constraints or operational bounds.
-For instance, if an agent’s output is about to violate a known policy (say, it’s about to generate disallowed content or leak PII), a monitoring module can catch this and either censor the output or stop the agent’s process.
-Some approaches involve using formal verification and runtime checking tools (like TLA+ specifications or symbolic execution) to prove certain unsafe actions will not occur, or to continuously verify that invariants hold during execution.
-10.4 Emerging Governance Practices
-Because the field of Agentic AI is evolving rapidly, governance strategies are also adapting.
-Beyond the foundational controls above, organizations are adopting new best practices to strengthen oversight and ensure long-term compliance as these AI systems scale.
-Some of the emerging practices include:
-  • AI Governance Boards: Many organizations are establishing dedicated governance boards or committees for AI oversight.
-These boards typically include stakeholders from legal, ethics, security, and domain experts.
-Their role is to continuously review the Agentic AI system’s objectives, assess risks, and update policies or strategies as needed.
-For example, a governance board might conduct regular audits of an AI’s decisions for bias or fairness, ensure alignment with the latest regulations, and approve any major changes to the system’s capabilities.
-This human layer of oversight creates accountability at the organizational level and sends a signal that the AI’s operations are being closely supervised by responsible parties.
-  • Governance-as-Code: Inspired by the infrastructure-as-code paradigm in DevOps, governance-as-code means encoding governance policies and ethical rules directly into the software and configuration that drive the AI system.
-Instead of governance being only a manual or paperwork process, it becomes part of the system’s logic.
-For Agentic AI, this could involve using policy engines or rule-based systems that automatically check each agent action against a set of coded rules.
-For instance, a rule repository might specify “Agent of type X cannot use tool Y without permission Z” and this is enforced by the orchestration code.
-  • Adaptive Governance Layers: Agentic AI systems may change over time – they might learn new skills, add more agents, or move into new deployment contexts.
-Adaptive governance refers to frameworks that evolve along with the AI system.
-This can include auto-updating rule sets or ML-driven policy adjustments.
-For example, if an agent starts operating in a new country, an adaptive governance layer might automatically incorporate that country’s AI regulations into the policy set.
-Or if a system update gives agents new capabilities, the governance rules adapt (or prompt human governors to adapt them) to cover those capabilities.
-  • Standards Alignment and Certification: To have trust, organizations are aligning their Agentic AI governance with emerging international standards.
-A prime example is ISO/IEC 42001:2023 [189], and ISO/IEC 42005:2025 [101] the management system standard, which specify requirements for effective AI governance and risk management .
-ISO 42001 and 42005 emphasizes many of the same principles discussed above – such as transparency, accountability, bias mitigation, safety, and privacy in AI development and deployment.
-Forward-looking organizations are therefore proactively building those capabilities (e.g.
-implementing rigorous risk management pipelines, bias audits, etc.) so that by the time new laws fully kick in, their agentic AI will be in natural compliance.
-By integrating these enforcement measures, Agentic AI systems can operate with much tighter alignment to intended guidelines and ethical norms.
-Policy enforcement translates high-level governance principles into day-to-day operational safeguards, reducing the chance of an autonomous agent straying outside its allowed domain or causing a compliance breach.
-Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
+  • Orchestrator-level controls. A meta-controller enforces allow/deny policies for plans and tool calls; escalate to human approval for sensitive operations (e.g., data deletion, external writes).
+  • Memory & retention. Implement TTLs, redaction, and scoped memory (task- and tenant-scoped) to satisfy minimization and purpose limits; support subject-access and deletion.
+  • Tool & data access. Enforce RBAC/ABAC for agents/roles; use per-tool API tokens, least privilege, and environment scoping (dev/stage/prod). Validate/normalize tool arguments.
+  • Isolation. Run untrusted code/tooling in sandboxes (e.g., containers/VMs) with syscall/network/file egress policies; prefer hardware isolation (TEEs) for sensitive workloads.
+  • I/O mediation. Apply content filters and egress controls on model outputs; mediate RAG inputs to prevent injection/exfiltration; detect sensitive data (PII/PHI, secrets) before release.
+  • Real-time monitoring. Track anomaly signals (policy violations, unusual tool chains, data exfiltration patterns) and enforce runtime kill-switches, rollbacks, and incident triggers.
+8.4 Documentation & Record-Keeping (Evidence)
+To meet audit and regulator expectations, maintain:
+  • Technical documentation (EU AI Act [1]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
+  • Model/Data cards with lineage, licenses, and known hazards; evaluation reports with stress tests (robustness, red-teaming), calibration/fairness metrics, and residual risk.
+  • Versioned artifacts: model versions, datasets, prompts, agents’ policies, and tool catalogs with change logs; signed releases and rollback points.
+8.5 Post-Market Monitoring & Incident Response
+Agentic systems must be monitored after deployment and respond to harm quickly.
+  • Telemetry and drift. Continuously measure real-world performance, jailbreak/abuse rates, and data distribution shifts; trigger retraining or guard updates.
+  • Serious-incident handling. Define thresholds, triage workflows, and reporting timelines; keep case files linking logs, affected users, and remediation.
+  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [129], ISO/IEC 23894 [130] ) when capabilities, data, or context change materially.
+8.6 Supply-Chain & Secure Development (GenAI)
+Strengthen provenance and development hygiene for agents and tools.
+  • Secure SDLC for GenAI.
+Follow NIST SP 800-218 (SSDF) [^17] NIST SP 800-218 (SSDF) and the GenAI profile SP 800-218A  [^18] GenAI profile SP 800-218A for model/system development; integrate threat modeling for agent tool-chains and LLM-specific risks.
+  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [137] Top 10 for LLM applications.
+  • Provenance. Track model/tool/dataset origins, licenses, and checksums; require attestations for third-party components and run dependency/secret scans in CI.
+Bringing Agentic systems into compliance is not a single checklist; it is a lifecycle program that couples standards (NIST AI RMF, ISO/IEC 42001/42005), regulation (EU AI Act), secure development (NIST SP 800-218/218A), and runtime enforcement (policies, isolation, monitoring).
+This combination creates evidenceable trust and reduces the blast radius of inevitable failures.
 9 Discussion
@@
 9.5 Future Roadmap for Agentic AI TRiSM
-Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
-Figure 8: Agentic AI Roadmap: Mindmap Representation
+Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Controllability Benchmarking & Evaluation Cognitive Capability Expansion
+Figure 7: Agentic AI Roadmap: Mindmap Representation.
 Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 7.
@@
 We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
+A cross-cutting concept relevant to both Governance and Lifecycle Management, controllability captures the ability of humans and oversight systems to steer or override agentic AI.
+Mechanisms include stop/pause functions, rollback gates role-based oversight, and auditability of inter-agent actions.
+In practice, controllability is critical to ensuring accountability and preventing runaway behaviors in multi-agent workflows.
 10 Conclusion
-The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services.
+The rapid advancement of Agentic AI, driven by AMAS, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services.
 However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations.
-This review has explored these challenges, proposing a TRiSM framework structured around four pillars: Explainability, ModelOps, Security, Privacy and Governance, adapted for LLM-based AMAS.
+This review has explored these challenges, proposing a TRiSM framework structured around five pillars: Explainability, ModelOps, Security, Privacy and their lifecycle Governance, adapted for LLM-based AMAS.
 Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency.
@@
 By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
+Acknowledgements
+Funding for the research was provided through Horizon Europe project AIXpert: An agentic, multi-layer, GenAI-powered backbone to make an AI system explainable, accountable, and transparent (ID:101214389).
 Declaration of Competing Interest

### v5

@@
 A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation.
-To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows.
+To make coordination and tool use measurable in practice, we propose two metrics: the Component Synergy Score (CSS), which captures inter-agent enablement, and the Tool Utilization Efficacy (TUE), which evaluates whether tools are invoked correctly and efficiently.
 We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance.
@@
 AI governance has moved from aspiration to obligation.
-The EU Artificial Intelligence Act [1] entered into force on 1 August 2024 with a staged application (e.g., prohibitions in early 2025 and obligations for general-purpose AI from 2 August 2025  [^3] EU AI Act); meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [2] and the NIST AI Risk Management Framework (AI RMF 1.0) [3] to operationalize risk controls.
-In parallel, enterprise use of AI continues to expand: recent global surveys [4, 5, 6] report that a substantial majority of organizations use AI in at least one business function.
-These trends sharpen demand for system-level trust, risk, and security management tailored to agentic systems.
-Figure 1: TRiSM taxonomy for Agentic AI (AMAS) presented in this review.
-We cover: (i) Fundamentals of Agentic AI (Agentic vs.
-traditional agents; LLM-based architectures), (ii) AI Trust, Risk, and Security Management (TRiSM) pillars: Explainability, ModelOps, Security, Privacy, and Lifecycle Governance, (iii) Trust management via explainability, (iv) Threats and risks, and (v) Evaluation metrics, including human-centered and composite measures.
-An AI agent is a computational entity that perceives its environment and acts to achieve goals [7].
-What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [8].
+The EU Artificial Intelligence Act [EU_AI_Act2025] entered into force on 1 August 2024 and applies progressively (e.g., prohibited AI practices and AI literacy obligations from 2 February 2025, and governance rules plus obligations for general-purpose AI models from 2 August 2025) [^3] EU AI Act; meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [iso42001] and the NIST AI Risk Management Framework (AI RMF 1.0) [ai2023artificial] to operationalize risk controls.
+In parallel, enterprise use of AI continues to expand, for example, recent global surveys [McKinsey_StateOfAI2025, Stanford_HAI_AIIndex2025, Gartner_AI_TRiSM2024] report that a substantial majority of organizations use AI in at least one business function.
+Critically, the emerging paradigm of AI agents, which are systems that autonomously plan, reason, and execute complex tasks, introduces qualitatively different risks than traditional AI applications.
+These trends demand system-level trust, risk, and security management tailored to agentic systems.
+Figure 1: Agentic AI research growth on arXiv (2019–2024), showing multi-agent systems and LLM-based agent publications with marked acceleration following ChatGPT’s release in November 2022.
+Figure 2: Google search interest for “AI Agents” vs “Agentic AI” (2022–2024), illustrating rising public awareness with key milestones including ChatGPT launch and Gartner’s recognition of Agentic AI as the top technology trend for 2025.
+An AI agent is a computational entity that perceives its environment and acts to achieve goals [russell1997rationality].
+What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [sapkota2025ai].
 In this paper, we refer to such LLM-based, coordinating agent systems as Agentic Multi-Agent Systems (AMAS).
-These systems orchestrate multiple specialized agents to address long-horizon tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents.
-The shift from single-agent or rule-based pipelines to AMAS unlocks new capabilities but complicates coordination, control, transparency, and risk management across heterogeneous components and contexts.
-Why TRiSM for Agentic AI?
-LLM-based AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [9].
+These systems orchestrate multiple specialized agents to address complex tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents.
+TRiSM for Agentic AI
+AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [yang2024multi].
 These risks are amplified when agents plan, coordinate, and call external tools.
-Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [10, 11].
-To address these system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks emphasize governance, explainability, security, privacy, and lifecycle controls [12, 13].
-We argue that a TRiSM perspective is necessary to make AMAS deployable in sensitive domains.
-Scope and Objectives
-Our focus is LLM-based AMAS that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning, with or without human oversight.
-[^4] Throughout, AMAS denotes multi-agent systems built on LLMs.
-We synthesize recent work on AMAS and AI TRiSM to provide a coherent, actionable reference for researchers and practitioners.
-Specifically, we map TRiSM pillars :Explainability, ModelOps, Security, Privacy, and their Lifecycle Governance, to AMAS workflows and failure modes.
-We also surface gaps where current ML governance and model-centric controls do not readily extend to multi-agent settings.
-Necessity of this Survey
-Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively under-addressing adversarial robustness, lifecycle governance, decision provenance, and system-level explainability.
+Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [lim2012memory, fang2025trustworthyaisafetybias].
+To address these multifaceted system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks provide integrated approaches emphasizing governance, explainability, security, privacy, and lifecycle controls [IBM2025AITRiSM, habbal2024artificial].
+Integrating TRiSM principles into the design and deployment of Agentic AI systems is increasingly urgent.
+Figure 1 demonstrates explosive growth in arXiv publications, with multi-agent systems papers increasing from 890 in 2019 to 18,500 in 2024, while LLM-based agent papers surged from near zero to 9,800 following ChatGPT’s release.
+This academic interest is mirrored by public awareness, as shown in Figure 2, where Google search trends for both “AI Agents” and “Agentic AI” reached peak levels by late 2024.
+Gartner’s designation of Agentic AI as a top strategic technology trend for 2025, alongside AI governance platforms within its TRiSM framing [^4] Gartner Market Guide for AI TRiSM, and its 2026 emphasis on multiagent systems, digital provenance, and AI security platforms  [^5] Gartner Top Technology Trends 2026 reflect a growing consensus: as AI systems gain autonomy, robust trust, risk mitigation, and security controls become indispensable.
+We emphasize that a TRiSM perspective is essential to make AMAS deployable in sensitive domains.
+1.1 Necessity and Scope of this Survey
+Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively less attention is given to adversarial robustness, lifecycle governance, decision provenance, and system-level explainability.
+This survey addresses this gap by focusing on AMAS [^6] Throughout, AMAS denotes multi-agent systems built on LLMs.
+that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning.
 As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks.
-To our knowledge, few surveys examine LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance; Table 1 situates our contribution relative to prior reviews.
-Table 1: Comparison with related surveys.
-Survey Threats Lifecycle Gov. Explainability TRiSM Integration LLM-Specific Applications Actionable Guidance
-Guo et al. (2024) [14] ~
-Chen et al. (2025) [15] ~ ~
-Yan et al. (2025) [16] ~
-Tran et al. (2025) [17] ~
-Lin et al. (2025) [18] ~ ~
-Fang et al. (2025) [11] ~ ~ ~
-Xi et al. (2025) [19] ~
-Luo et al. (2025) [20] ~ ~ ~
-Zou et al. (2025) [21] ~ ~
-Wang et al. (2025) [22] ~ ~ ~ ~
+To our knowledge, no prior survey examines LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance.
+Table 1 situates our contribution relative to prior work.
+1.2 Contributions
+The contributions of this study are manifold:
+  • Unified TRiSM framework for Agentic AI.
+We present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to AMAS (Section 4).
+The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualizing each pillar to AMAS workflows and failure modes.
+  • Risk taxonomy for AMAS.
+We synthesize a taxonomy of threats unique to AMAS, including prompt injection, memory poisoning, collusive failure, emergent misbehavior and tool-use abuse, and map these risks to corresponding controls (Section 3).
+  • Evaluation template and new metrics.
+We propose a holistic evaluation template covering trustworthiness, explainability, user-centered performance and inter-agent coordination.
+To operationalize this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilization Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 5).
+  • Technique mapping and gap analysis.
+We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting, revealing where existing approaches can be adapted and where significant gaps remain (Section 7).
+Building on these contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems.
+Key directions include adversarial robustness, governance protocols, standardized benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
+1.3 Related Work
+Prior surveys on LLM-based multi-agent systems mainly emphasize architectures, agent capabilities, communication, and domain applications [guo2024largelanguagemodelbased, chen2025surveyllmbasedmultiagentsystem, yan2025selftalkcommunicationcentricsurveyllmbased, tran2025multiagentcollaborationmechanismssurvey].
+In parallel, broader trustworthy/responsible AI surveys discuss themes aligned with TRiSM, such as alignment, fairness, and privacy defenses; but typically focus on general ML systems and do not model AMAS-specific dynamics (e.g., inter-agent information sharing, tool-mediated actions, and emergent coordination) [fang2025trustworthyaisafetybias, raza2024exploring].
+As a result, key gaps remain in adversarial robustness, governance and accountability for multi-agent deployments, and explainability at scale.
+In comparison to these works (Table 1), our study addresses these gaps by mapping TRiSM principles to AMAS and operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings.
+Table 1: Comparison with representative related surveys on LLM-based agentic systems.
+Survey Threats Lifecycle Explain. TRiSM LLM-Spec. Apps. Actionable
+Guo et al. (2024) [guo2024largelanguagemodelbased] ~ ~
+Chen et al. (2024) [chen2025surveyllmbasedmultiagentsystem] ~ ~
+Yan et al. (2025) [yan2025selftalkcommunicationcentricsurveyllmbased] ~ ~
+Tran et al. (2025) [tran2025multiagentcollaborationmechanismssurvey] ~
+Lin et al. (2025) [lin2025creativityllmbasedmultiagentsystems] ~ ~
+Fang et al. (2025) [fang2025trustworthyaisafetybias] ~ ~ ~
+Xi et al. (2025) [xi2025rise] ~ ~
+Luo et al. (2025) [luo2025large] ~ ~ ~
+Zou et al. (2025) [zou2025llm] ~ ~
+Wang et al. (2025) [wang2025surveyllmbasedagentsmedicine] ~ ~
+Wang et al. (2025) [11081880] ~
+Yu et al. (2025) [yu2025survey] ~
 This Survey (2025)
 Legend: ✓ = explicitly addressed; ~ = partially addressed; ✗ = not addressed.
-Contributions
-We make the following contributions:
-  • Unified TRiSM framework for Agentic AI.
-We propose and present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to large-language-model-based multi-agent systems (Section 5).
-The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualising each pillar to AMAS workflows and failure modes.
-  • Risk taxonomy for AMAS.
-We synthesise a taxonomy of threats unique to LLM-driven multi-agent systems, such as prompt injection, memory poisoning, collusive failure, emergent misbehaviour and tool-use abuse; and map these risks to corresponding controls (Section 4).
-  • Evaluation template and new metrics.
-We propose a holistic evaluation template that covers trustworthiness, explainability, user-centred performance and inter-agent coordination.
-To operationalise this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilisation Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 6).
-  • Technique mapping and limitations.
-We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting.
-This mapping reveals where existing approaches can be adapted and where significant gaps remain (Section 8).
-Building on the above contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems.
-Key directions include adversarial robustness, governance protocols, standardised benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
-2 Background, Related Work and Literature Review Methodology
-2.1 AI Agents vs. Agentic AI
-An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals [7], has undergone a rapid transformation in recent years.
-While early agents were task-specific and deterministic, modern systems have evolved substantially, increasing operational autonomy on benchmark tasks [23].
-These agentic systems integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [8].
-This evolution presents new opportunities but also raises challenges in coordination, control, transparency, and risk assessment across heterogeneous agents.
-As the shift toward agentic architectures accelerates, long-standing concerns around safety, fairness, interpretability, and systemic bias resurface [11], especially in high-stakes domains such as healthcare, law, and public services, where decisions can carry profound societal consequences.
-Autonomous software agents date back to the 1990s [24]; however, the advent of LLM-driven multi-agent ecosystems since 2023 [9] has made “Agentic AI” a distinct and rapidly expanding research focus [8].
-Traditional agents primarily automated limited tasks—such as information retrieval [25], data summarization [26], and dialogue response [27]—often via single-step logic [28] or scripted rules.
-These approaches lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem-solving.
-In contrast, agentic systems coordinate collaborative agents with specialized roles (e.g., planner, coder, analyst) enabled by LLMs and tool use [29].
-They dynamically decompose tasks, share context, and pursue high-level goals over long horizons [30].
-This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent, decentralized behavior.
-2.2 Trust, Risk, and Security Challenges in Agentic AI
-The rise of Agentic AI introduces challenges that extend beyond those of traditional ML systems.
-Unlike deterministic agents [31], LLM-based multi-agent systems (AMAS) exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors.
-These characteristics necessitate robust approaches to building trust and ensuring trustworthiness in high-stakes domains such as healthcare, finance, and law.
-Below, we discuss several key challenges:
-  1. 1.
-    Trust and trustworthiness.
-In Agentic AI, trust refers to a user’s willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner.
-According to ISO [32], trustworthiness is a quality characteristic of systems and is understood as the ‘ability to meet stakeholders’ expectations in a verifiable way’.
-This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications.
-The stochastic nature of LLMs [33] further complicates trustworthiness: randomness in output generation, while sometimes promoting creativity, introduces inconsistency and limits repeatability.
-  2. 2.
-    Security risks in multi-agent systems.
-Risk surfaces expand as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [10].
-For example, a multi-agent supply-chain optimizer may coordinate procurement and logistics agents yet inadvertently leak sensitive data or violate compliance protocols (e.g., [34], HIPAA [35]) if safeguards are inadequate.
-  3. 3.
-    Prompt injection, spoofing, and impersonation.
-AMAS are susceptible to security threats that exploit language-based interfaces and cooperative behaviors.
-One prominent threat is prompt injection [36], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior.
-Recent studies [37] describe “prompt infection”, in which malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [38].
-Corrupted outputs can cascade, leading to data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
-    Another critical vulnerability involves spoofing and impersonation [39].
-Spoofing refers to falsifying identity or credentials (e.g., using a false name or address) to deceive a system, whereas impersonation involves pretending to be a specific agent or user to gain trust, access, or privileges.
-In AMAS, where agents rely on credentials or tokens for authentication during coordination, an adversary who steals an API key or mimics a trusted peer can issue unauthorized commands or access sensitive information.
-  4. 4.
-    Privacy risks in multi-agent LLM systems.
-LLM-based agents often process user data, proprietary business information, and other sensitive inputs to fulfill their tasks [40].
-In a multi-agent context, this challenge is amplified by inter-agent information sharing (e.g., via a shared memory store or message passing), which complicates data minimization, access control, and purpose limitation.
-Traditional evaluation and safety frameworks, designed for deterministic, single-agent pipelines; are inadequate for the autonomy and control properties of AMAS.
-To address these gaps, AI Trust, Risk, and Security Management (AI TRiSM) [12, 13] proposes lifecycle controls spanning explainability, secure model/tool orchestration, and privacy management.
-Rather than focusing solely on technical performance, TRiSM integrates governance practices to strengthen reliability and robustness, protect data, and align system behavior with ethical and regulatory expectations.
-2.3 Survey Landscape
-LLM-based multi-agent surveys (technical focus)
-Prior surveys on LLM-driven AMAS predominantly address system architectures, agent capabilities, and domain-specific applications of agentic systems.
-For example, [14, 15] examine simulated environments, inter-agent communication protocols, and performance benchmarks.
-These works clarify how agents coordinate and scale but give limited treatment to adversarial threats (e.g., prompt injection, collusion), lifecycle governance (e.g., continuous monitoring, audit, regulatory compliance), and system-wide explainability, especially for emergent behaviors in complex settings.
-More specialized surveys, such as [16], focus on natural-language communication, while [17] details collaboration mechanisms and practical use cases in domains like question answering, 5G networks, and Industry 5.0.
-Although [16] highlights security risks associated with communication channels and [17] discusses deployment contexts, both remain narrowly centered on coordination strategies.
-Deeper analysis is needed on robustness to adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability at scale.
-Trustworthy/responsible AI surveys (broad trust focus)
-Review articles addressing TRiSM-related themes (e.g., [11]; see also fairness-focused work [41]) discuss alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage).
-However, these surveys typically target general ML systems and overlook dynamics unique to LLM-based multi-agent settings.
-For instance, while they advocate privacy-preserving techniques such as differential privacy, they rarely address privacy risks stemming from inter-agent data sharing or the distributed nature of LLM-driven workflows.
-Moreover, critical TRiSM aspects—such as governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents), are often under-specified [42].
-2.3.1 Positioning of this study
-This review maps TRiSM principles to LLM-based AMAS, operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings.
-Relative to prior surveys (Table 1), we foreground system-level risks (e.g., prompt infection, impersonation, memory poisoning), propose evaluation dimensions and composite measures for coordination and tool use, and offer actionable guidance for researchers, engineers, and policymakers.
 1.4 Literature Review Methodology
 We conducted a structured literature review to synthesize AI TRiSM principles for AMAS.
-Guided by established systematic-review practices [43, 44], this section states our research objectives and describes the data sources and search strategy, the inclusion/exclusion criteria and screening procedure, and the classification schema used for analysis.
-(a) Agentic AI papers per year (2022–2025).
-(b) Papers by TRiSM pillar.
-(c) Topic trends in Agentic AI (agents, security, explainability).
-(d) Bibliography entries per year (1987–2025).
-Figure 2: Bibliographic analysis of Agentic AI research.
-(a) Annual trend in Agentic AI publications, showing a sharp rise post-2023.
-(b) Distribution of TRiSM-related publications across governance and safety topics.
-(c) Topic-specific trends in agents, security, and explainability.
-(d) Overall growth in bibliographic entries, highlighting acceleration since 2020.
+Following established systematic review guidance [boland2017doing, keele2007guidelines], we describe our data sources, search strategy, screening criteria, and classification schema.
 Research Questions (RQs)
 The review was guided by the following three questions:
-  • RQ1: What trust, risk, and security challenges are specific to LLM-based AMAS?
+  • RQ1: What trust, risk, and security challenges are specific to AMAS?
   • RQ2: Which governance, explainability, security, privacy, and lifecycle (TRiSM) controls are proposed or adapted for AMAS?
@@
 We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar.
-The search covered publications from January 2022 to May 2025, with select seminal works prior to 2022 included for foundational context.
-We used Boolean queries and expanded keywords to capture the intersection of LLM-based multi-agent systems and TRiSM considerations.
-Example query template:
-("Agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR "distributed AI" OR "LLM-based agents" OR "agent coordination" OR "agent-based modeling" OR "swarm intelligence")
+The search covered publications from January 2022 to December 2025, with influential foundational works prior to 2022 included for theoretical context.
+We used a two-block Boolean query: (agentic/AMAS terms) AND (TRiSM terms).
+("agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents"
+OR  "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR
+ "distributed AI" OR "LLM-based agents" OR "agent coordination" OR
+ "agent-based modeling" OR "swarm intelligence")
 AND
-("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection" OR "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR "system reliability")
+("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR
+ "governance" OR "oversight" OR "compliance" OR "explainability" OR
+ "interpretability" OR "transparency" OR "privacy" OR "data protection" OR
+ "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR
+ "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR
+ "system reliability")
 Screening and Eligibility
-Two-stage screening (title/abstract, then full-text) used these criteria. Inclusion:
-  • Studies published 2022–2025 (capturing the post-2022 LLM wave), plus select pre-2022 foundational work.
-  • Studies explicitly discussing Agentic AI, multi-agent, or LLM-powered systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
-  • Peer-reviewed articles and widely cited preprints or reports from credible sources (e.g., NeurIPS, IEEE, ACM, Nature, governmental/standards bodies, arXiv).
-Exclusion:
-  • Papers focused solely on traditional, rule-based, or symbolic agents without LLM integration or emergent coordination.
-  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) with no agentic behavior or TRiSM concerns.
-  • Non-English papers or those lacking sufficient metadata (e.g., missing abstract or publication details).
-For empirical papers, we assessed clarity of threat models, reproducibility (code/data), evaluation rigor (baselines, ablations), and external validity.
-For surveys, we assessed coverage breadth, structure, and treatment of risks/governance.
-After screening titles and abstracts, we shortlisted 250 papers; full-text review yielded 180 primary studies addressing one or more TRiSM pillars.
-We also incorporated relevant technical reports and authoritative web resources to contextualize the rapidly evolving Agentic-AI/TRiSM landscape.
-A bibliometric summary appears in Figure 2.
+We applied a two-stage screening process (title/abstract, then full text).
+Inclusion criteria:
+  • Published between 2022–2025 (plus influential foundational papers published before 2022).
+  • Focused on agentic AI, multi-agent, or LLM-powered systems and addressed at least one TRiSM aspect (trust, risk, security, governance, explainability, or privacy).
+  • Peer-reviewed papers, widely cited preprints, technical reports from reputable venues or organizations (e.g., IEEE/ACM conferences and journals, standards bodies, and government sources).
+Exclusion criteria:
+  • Traditional rule-based or symbolic agents without LLM-based autonomy or coordination.
+  • Papers limited to low-level model training/architecture details with no agentic behavior or TRiSM relevance.
+  • Non-English papers or items with insufficient bibliographic information (e.g., missing abstract or venue details).
+For empirical studies, we verified that the threat model and evaluation were clearly described and assessed reproducibility when code or data were available.
+For survey papers, we evaluated coverage, structure, and treatment of risks and governance.
+After title/abstract screening, we shortlisted 250 papers; full-text screening yielded 180 primary studies covering one or more TRiSM pillars, supplemented by relevant standards and technical reports for regulatory context.
 Quality Assurance
-To ensure reliability, we adapted quality criteria from established guidelines [45].
-Each paper was evaluated on: (1) clear objectives for agentic or LLM-based systems; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to Agentic-AI governance.
+To ensure reliability, we adapted quality criteria from established guidelines [kitchenham2004procedures].
+Each paper was evaluated on: (1) clear objectives for AMAS; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to agentic AI governance.
 Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only.
-Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion (with inter-rater agreement checked on a subset).
+Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion.
+Bibliometric Overview
+The results of our screening process reveal the rapid acceleration of agentic AI research.
+The publication landscape (Figure 3) shows journal articles dominating at 61.2%, with leading venues and publishers detailed in Figure 4.
+Key research themes, including large language models, multi-agent systems, and security, are visualized in Figure 5a, while Figure 6 confirms the unprecedented growth of research activity in recent years.
+(a) Publication trends (1985–2025)
+(b) Publication type distribution
+Figure 3: Publication landscape: (a) Long-term trends showing accelerating research output over four decades; (b) Distribution of publication types, with journal articles dominating at 61.2%.
+(a) Top publication venues
+(b) Publisher distribution
+Figure 4: Publication outlets: (a) Top venues including NeurIPS, ACM Computing Surveys, and AAAI; (b) Publisher distribution led by Elsevier, Springer, and IEEE.
+(a) Research topics and keywords
+(b) Publication type evolution over time
+Figure 5: Research themes and evolution: (a) Word cloud highlighting key topics including large language models, multi-agent systems, security, and explainability; (b) Temporal distribution of publication types.
+Figure 6: Comparison of recent publications (2020–2025) with historical distribution by period, demonstrating the rapid acceleration of agentic AI research in recent years.
+Figures 7 and 8 provide a roadmap for the paper, which progresses from architectural foundations to risk analysis, evaluation, and governance.
+TRiSM for Agentic AIFundamentalsThreats & RisksTRiSM FrameworkEvaluation
+$\circ$ AMAS Architecture
+$\circ$ LLM Core
+$\circ$ Planning/Reasoning
+$\circ$ Memory Module
+$\circ$ Tool Interface
+$\circ$ Adversarial Attacks
+$\circ$ Data Leakage
+$\circ$ Agent Collusion
+$\circ$ Emergent Behavior
+$\circ$ Memory Poisoning
+$\circ$ Explainability
+$\circ$ ModelOps
+$\circ$ Application Security
+$\circ$ Model Privacy
+$\circ$ Governance
+$\circ$ Trustworthiness
+$\circ$ CSS Metric
+$\circ$ TUE Metric
+$\circ$ Compliance
+$\circ$ Auditability
+Figure 7: Taxonomy of TRiSM for LLM-based Agentic Multi-Agent Systems.
+Figure 8: TRiSM taxonomy for Agentic AI (AMAS) presented in this review.
+The paper progresses through six phases: (i) Fundamentals of Agentic AI: distinguishing AMAS from traditional agents and establishing LLM-based architectures (Section 2); (ii) AI Trust, Risk, and Security Management (TRiSM) framework pillars: Explainability, ModelOps, Application Security, Model Privacy, and Lifecycle Governance, introduced conceptually (Section 4); (iii) Trust management via explainability techniques (Section 4.2.1); (iv) Threats and risks taxonomy with real-world case studies (Section 3); (v) Evaluation metrics encompassing trustworthiness, coordination, and composite measures (Section 5); (vi) Technical implementation of security/privacy mechanisms (Section 6); and (vii) Compliance and governance integration (Section 7).
 2 Fundamentals of Agentic AI Systems
 In this section, we present the fundamentals of Agentic AI systems.
-Table 2: Traditional AI Agents vs. Agentic AI Systems
+Table 2: Comparison of Traditional AI Agents and Agentic AI Systems
 Dimension Traditional AI Agents Agentic AI Systems
-Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
-$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
-Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
-$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
-Intelligence Scope × Narrow & task-specific Broad & compositional
-$\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
-Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
-$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
-Collaboration × Isolated or limited coordination Role-specialized coordination
-$\bullet$ Manual decomposition $\bullet$ Automated hierarchy
-Temporal Context × Episodic, often stateless Persistent memory
-$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
-Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
-$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
-Tool Utilization × Static or domain-specific tools Planned API invocation
-$\bullet$ Handcrafted interfaces $\bullet$ Function calling
-Context Awareness × Bounded or heuristic context Memory-augmented
-$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
-Learning Mechanism × Limited or rule-based learning Emergent learning
-$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
-Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
-$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
-Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
-= Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
-3.1 Traditional AI Agents vs Agentic AI
-Traditional agents operate through predefined rules [46], heuristic workflows [47], or deterministic logic [48], and perform well in narrow, well-defined environments.
-Cognitive Agents [49], model human-like cognitive functions, such as integrating perception, memory, and decision-making, often using cognitive systems architectures like [50] or ACT-R [51].
-According to a recent work [52], authors distinguish between 4 types of knowledge: Concepts, Skills, Processes, Motives for cognitive agents.
-A key difference between Agentic AI and such cognitive architecture approaches is that the latter were mostly applied to robot planning and control (so applies to ‘physical operational environments’) but Agentic AI applies more broadly, also to non-physical operational environments (which could be made to operate entities in physical also environments).
-In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities.
-Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [30, 53].
-(2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [54].
-(3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [55].
-This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [56].
-For example, systems like AutoGen [57] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [58].
-This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [53].
-A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
-For the scope of this work, we rely on LLM-based Agentic AI systems.
+Autonomy Reactive with fixed action sequences Goal-driven with adaptive behavior
+Cognitive Foundation Symbolic logic, finite state machines Foundation models (large language models)
+Domain Scope Narrow, task-specific applications Broad, cross-domain capabilities
+Reasoning Mechanism Rule-based, deterministic inference Multi-step, contextual reasoning (chain-of-thought)
+Collaboration Isolated agents, manual task decomposition Role-specialized agents, automated coordination
+Memory Architecture Episodic, session-bound storage Persistent memory (vector databases, long-term memory)
+Orchestration Hard-coded, predefined workflows Dynamic orchestration via meta-agents
+Tool Integration Static, hand-crafted interfaces Dynamic planning with API/function calling
+Context Handling Bounded context with heuristic rules Memory-augmented retrieval (RAG)
+Learning Paradigm Rule updates, fixed knowledge base Emergent learning (fine-tuning, reinforcement learning)
+User Interaction Scripted, template-based dialogue Natural language understanding and generation
+Representative Systems MYCIN [daniel1997cadiag], ELIZA [walters2025eliza], SOAR [laird1987soar], ACT-R [taatgen2006modeling] AutoGen [autogen_agents_tutorial], ChatDev [qian2023chatdev], MetaGPT [hong2023metagpt]
+2.1 Traditional AI Agents vs. Agentic AI
+Traditional agents rely on predefined rules, heuristics, or deterministic logic and typically perform well in narrow, well-specified environments [muscettola1998remote, hannebauer1999formal, govaerts2010development].
+Cognitive agents extend this paradigm by modeling human-like functions (perception, memory, decision-making) through cognitive architectures such as SOAR and ACT-R [langley2006cognitive, laird1987soar, anderson1993act]; recent work further categorizes their knowledge into concepts, skills, processes, and motives [langley2025spatial].
+However, these approaches were largely developed for physical operational settings (e.g., robotic planning and control), whereas Agentic AI generalizes autonomy to both physical and non-physical environments.
+In contrast, Agentic AI systems leverage foundation models (primarily LLMs) to support adaptive, goal-directed behavior via (i) multi-agent coordination, (ii) persistent context through memory, and (iii) dynamic orchestration of roles and tool use [qian2023chatdev, hong2023metagpt, borghoff2025human, schneider2025generative].
+This enables long-horizon task execution and cross-domain generalization beyond traditional reactive pipelines [de2025open].
+Table 2 summarizes these differences; in this survey, we focus on LLM-based Agentic AI systems.
 2.2 Multi-Agentic AI System (AMAS) Architecture
-AMAS represents an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks.
-As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem.
-At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [59].
-These agents access a Shared Toolchain Interface [60] to execute code, perform searches, or interact with domain-specific APIs.
 Figure 9: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms).
-The modular design highlights adaptive intelligence and inter-agent synergy.
-As shown in Figure 3), communication and coordination are facilitated through a Communication Middleware [61] that allows agents to share goals, observations, or intermediate results.
-A Task Manager or Orchestrator [62] governs high-level planning and delegates subtasks to agents based on their roles or specializations.
-Agents can read from and write to a World Model or Shared Memory [14], which stores contextual knowledge, system state, or evolving task data.
-Human oversight is supported through a Human-in-the-Loop Interface [63], enabling users to prompt, correct, or halt agent behavior.
-To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [64].
-However, this modular and distributed structure introduces significant TRiSM [65] challenges.
-With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [66].
-Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [67], especially when data traverses multiple agents or tools.
-Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [68].
-Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors.
-Below, we discuss the architecture of AMAS.
-Language Model Core (Agent Brain).
-At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [69].
-The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access).
-It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats.
-In many agent frameworks, such as AutoGPT [70], Baby AGI [71] and GPT Engineer [72], the LLM governs the full control loop, orchestrating the overall system behavior.
-Planning and Reasoning Module.
-To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals.
-This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [73], where the model performs intermediate reasoning before arriving at a final decision.
-Reasoning is also one aspect of cognitive functionalities [74].
-Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making.
-Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations.
-Techniques like Reasoning and Acting (ReAct) [75] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
-Memory Module.
-Agentic AI systems integrate memory to maintain context across iterations.
-This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [9].
-Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search.
-By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning.
-Effective memory management enables adaptive, learning-driven behavior.
+As shown in Fig.
+9, AMAS coordinate multiple LLM-based agents through a Communication Middleware [weyns2009agent], a Task Manager/Orchestrator [langgraph2024], and a World Model/Shared Memory [guo2024largelanguagemodelbased] that stores system state and task artifacts.
+Human oversight is provided via a Human-in-the-Loop (HITL) interface [cranshaw2017calendar].
+To support trustworthy operation, a Trust and Audit module records actions and tool usage [kaur2024building], while a Security Gateway and Privacy Management Layer enforce access control and reduce leakage across agents and tools [zhong2023towards, feretzakis2024privacy].
+An Explainability Interface provides interpretable rationales for multi-agent decisions to support transparency and trust calibration [singh2025explainable].
+Communication & Collaboration (Agent Communication and Coordination)
+Beyond basic messaging, recent work formalizes AMAS communication through explicit protocols.
+Google’s Agent-to-Agent (A2A) [^7] A2A Protocol supports discovery, capability advertisement, delegation, and result exchange, improving interoperability and auditability.
+Related efforts such as the Agent Network Protocol (ANP) [^8] Agent Network Protocol take a network-centric view with standardized interfaces for coordination and trust establishment.
+From a TRiSM perspective, protocolized communication exposes control points for authentication, authorization, logging, and lifecycle governance.
+Emerging semantic-aware communication [guo2024survey, wang2025internet] can improve coordination by encoding intent, but introduces risks such as semantic spoofing and privacy leakage.
+Cognitive/Reasoning Layer (LLM Core + Planning)
+At the center of an AMAS, a foundation model (typically an LLM) acts as the decision controller, taking a user goal plus role and tool specifications as input and generating the next actions (e.g., messages, tool calls, or structured function outputs) [Saravia_LLM_Agents_2024].
+In agent frameworks, this controller drives an iterative loop of planning, acting, and updating context based on tool results and feedback.
+Planning and Reasoning Module
+To achieve complex goals, AMAS employ planning mechanisms that decompose objectives into sub-tasks and interleave reasoning with execution and feedback.
+This can be implemented through prompting strategies (e.g., CoT/ToT) or through explicit plan–act–observe loops such as ReAct [yao2023reactsynergizingreasoningacting, Saravia_Prompt_Engineering_Guide_Web_2024].
+The plan is typically revised as new observations arrive, enabling adaptive long-horizon behavior.
+Memory Module
+AMAS integrate memory to maintain context across steps and sessions.
+Short-term memory is maintained in the working context, while long-term memory stores salient facts and past events for later retrieval (often via embedding-based search) [yang2024multi].
+Memory improves continuity and reduces repetition, but also increases exposure to stale context and poisoning risks.
+Learning Module
+Some AMAS include learning components that adapt behavior over time (e.g., from feedback, rewards, or task outcomes), such as updating policies, prompt templates, or routing strategies based on past performance.
+In multi-agent settings, learning can amplify errors through feedback loops and introduce risks such as drift, poisoning, and emergent misalignment, motivating continuous monitoring and gated updates under TRiSM controls.
+Data Storage & Retrieval
+AMAS typically combine structured stores for system state and task metadata with vector-based retrieval to support RAG and long-term memory.
+Because data may be shared across agents and tools, storage and retrieval require strong access control, encryption, and audit logging to reduce leakage and support accountability [he2024security].
+Tool-Use Interface
+To extend capabilities beyond text generation, AMAS agents use a tool-use interface that enables structured calls to external systems (e.g., web search, APIs, code execution, and databases) [schick2023toolformerlanguagemodelsteach].
+Tools and call schemas are specified in the agent configuration; when needed, the LLM emits a structured command that is executed externally, and the returned output is incorporated as a new observation for subsequent reasoning.
+Tool-routing frameworks such as MRKL [karpas2022mrklsystemsmodularneurosymbolic] exemplify this design by dispatching requests to specialized expert modules under an LLM controller, while training-based approaches (e.g., Toolformer) teach models to generate API calls during generation [schick2023toolformerlanguagemodelsteach].
+Perception and Environment Interface
+For agents interacting with dynamic environments (e.g., web interfaces, simulated worlds, or physical systems), an observation–action interface is essential [liu2025agent].
+Perception modules translate raw inputs (sensor streams, images, or textual state) into representations the LLM can process; the agent’s selected actions are then executed in the environment, and resulting state changes are returned as new observations.
+This loop supports iterative sense–plan–act behavior until a task is completed or halted.
+In multimodal or robotic settings, additional perception models (e.g., vision encoders) may be included, while the core interaction pattern remains the same.
+Monitoring and Governance Layer
+A monitoring and governance layer provides system-wide oversight by collecting logs, traces, and metrics, enforcing policy constraints, and triggering alerts or interventions when anomalous behavior is detected.
+From a TRiSM perspective, this layer enables auditability and lifecycle governance across agent interactions, tool calls, and memory updates.
+Integration and Autonomy
+Together, these modules form an integrated architecture in which the LLM coordinates planning, tool use, and memory to act in an environment and update context from feedback [acharya2025agentic].
+Table LABEL:tab:agentic_architectures summarizes representative systems and maps them to five design axes (LLM core, planning/reasoning, memory, tool use, and environment interface).
+Key Insights
+Overall, AMAS combine planning, tool use, memory, and environment interaction into a closed-loop system, which expands the attack surface and complicates accountability.
+Protocolized communication and monitoring create practical control points for authentication, logging, and governance, motivating TRiSM-oriented analysis for multi-agent deployments.
 Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
-Acronyms: CoT = Chain-of-Thought reasoning, FSM = Finite State Machine, PDDL = Planning Domain Definition Language, DAG = Directed Acyclic Graph, DB = Database, REPL = Read–Eval–Print Loop, API = Application Programming Interface, KV = Key–Value store, CLI = Command Line Interface, SDK = Software Development Kit, MCP = Model Context Protocol.
 Framework Core LLM Planning Memory Tool Use Notable Features
-AutoGPT [76] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
-BabyAGI [71] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
-GPT Engineer [77] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
-LangGraph [78] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
-AutoGen [57] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
-MRKL [79] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
-Reflexion [80] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
-MetaGPT [81] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
-Voyager [82] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
-WebVoyager [83] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
-HuggingGPT [84] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
-CAMEL [85] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
-ChatDev [86] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
-CrewAI [87] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
-AgentVerse [88] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
-OpenAgents [89] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
-SuperAGI [90] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
-Semantic Kernel [91] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
-OpenAI Swarm [92] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
-OpenAI Agents SDK [93] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
-Strands Agents [94] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
-LlamaIndex Agents [95] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
-Tool-Use Interface.
-To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [96].
-This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases.
-The available tools are specified in the agent prompt with command schemas.
-When the LLM determines a tool is needed, it emits a structured command, which is executed externally.
-The result is fed back into the LLM as a new observation.
-This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically.
-Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [79] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router.
-Similarly, approaches like Toolformer [96] train the model to insert API calls in its text generation.
-Perception and Environment Interface.
-For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [97].
-Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process.
-Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations.
-This loop supports sense-plan-act cycles that continue until the task is completed or halted.
-In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
-Integration and Autonomy.
-These modules together form a closed-loop architecture.
-The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [29].
-Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback.
-This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
-Table  LABEL:tab:agentic_architectures presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
-4 Threats and Risks in LLM-based Agentic Multi-Agent Systems
-In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
+AutoGPT [yang2023auto] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
+BabyAGI [nakajima_babyagi_archive_2024] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
+GPT Engineer [gpt-engineer] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
+LangGraph [wang2024agent] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
+AutoGen [wu2023autogen] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
+MRKL [karpas2022mrklsystemsmodularneurosymbolic] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
+Reflexion [shinn2023reflexionlanguageagentsverbal] GPT-3.5/4 Retry-reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
+MetaGPT [hong2024metagpt] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
+Voyager [wang2023voyageropenendedembodiedagent] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
+WebVoyager [he2024webvoyager] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
+HuggingGPT [shen2023hugginggpt] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
+CAMEL [li2023camel] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
+ChatDev [qian2024chatdevcommunicativeagentssoftware] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
+CrewAI [crewAI] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
+AgentVerse [chen2023agentverse] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
+OpenAgents [xie2023openagentsopenplatformlanguage] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
+SuperAGI [superagi] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
+Semantic Kernel [microsoft_semantickernel] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
+OpenAI Swarm [openai_swarm] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
+OpenAI Agents SDK [openai_agents_sdk] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
+Strands Agents [strands_agents] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
+LlamaIndex Agents [llamaindex_agents] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
+3 Threats and Risks in Agentic AI
+This section summarizes major threats in Agentic AI and organizes emergent AMAS risks into a compact taxonomy.
+3.1 Threats in Agentic AI
+Agentic AI systems introduce security and reliability concerns beyond single-agent LLM pipelines:
+  • Autonomy abuse. Agents with delegated authority may misinterpret objectives or execute harmful plans due to reasoning errors or manipulated inputs.
+  • Persistent memory.
+Long-term memory improves continuity but creates a contamination channel: injected or low-quality content can persist and propagate through shared stores without strong versioning and audit controls.
+  • Agent orchestration. Centralized or distributed orchestrators mediate roles and workflows; compromise or misconfiguration can misroute information and trigger cascading failures.
+  • Goal misalignment and exposure. Poorly scoped or exposed goal representations can enable prompt extraction, misalignment, or adversarial steering, especially when third-party tools are invoked.
+  • Tool misuse and external API exploits.
+Tool-enabled agents introduce new attack surfaces (e.g., unsafe code execution, costly actions, policy violations); sandboxing and invocation tracing are often inconsistent across open-source frameworks.
+  • Multi-agent collusion or drift.
+Agents can converge on mutually reinforcing errors (groupthink/mode collapse) or drift in self-reinforcing loops, complicating interpretability and safety alignment.
+3.2 Risks Taxonomy in AMAS
+We group AMAS risks into four classes—adversarial attacks, data leakage, agent collusion, and emergent behaviors—as shown in Fig. 10.
 Figure 10: Overview of key risks and threats in agentic AI systems.
-The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration.
-The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.
-4.1 Threats in Agentic AI
-Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures. Below, we discuss some of these major threats.
-  • Autonomy abuse.
-The foremost threat is the autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs .
-Unlike deterministic models, these Agentic AI systems dynamically generate actions, complicating efforts to define and enforce safe operational states .
-  • Persistent memory.
-Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential prompt injections.
-Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms .
-  • Agent orchestration.
-This risk involves central or distributed control mechanisms for role assignment and workflow mediation.
-A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [53] and AutoGen [98] .
-These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
-  • Goal Misalignment and Exposure.
-Agentic AI systems often operate with internalized or user-specified goals.
-Improper scoping or exposure of goal representations can lead to prompt extraction, misalignment, or competitive agent behaviors, particularly when third-party tools or APIs are invoked dynamically during task completion.
-  • Tool Misuse and External API Exploits.
-Agents capable of invoking external tools (e.g., via APIs, code execution, or web access) present novel attack surfaces.
-Misuse of these tools, intentional or accidental, can trigger financial costs, Distributed Denial of Service (DDoS) behavior, or legal violations.
-Sandboxing and invocation tracing remain limited in most open-source frameworks.
-  • Multi-Agent Collusion or Drift.
-Emergent coordination among agents, while useful for collaboration, can also lead to collusion, competitive drift, or polarization of outputs, especially in self-reinforcing memory loops or agent swarms.
-Such behavior complicates interpretability and may undermine safety alignment goals.
-4.2 Taxonomy of Risks
-We propose a taxonomy of risks associated with AMAS, categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 4.
-These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS.
-Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
+The left panel outlines system-level threats (e.g., autonomy abuse, persistent memory, orchestration).
+The right panel summarizes emergent AMAS risk classes arising from multi-agent interaction, memory persistence, and decentralized decision-making.
   • Adversarial attacks.
-Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps.
-These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [99].
-For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [30].
-Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
+AMAS are vulnerable to prompt injection, engineered reasoning traps, and other manipulations that can cascade across agents; e.g., role-swapping attacks in ChatDev exploit inter-agent dependencies [miao2025autonomous, qian2023chatdev].
   • Data leakage.
-The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information.
-In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [100].
-For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls.
-Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
-  • Agent Collusion and Mode Collapse.
-Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [101].
-This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies.
-For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance.
-Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
+Shared memory and inter-agent communication increase the chance of unintended disclosure of sensitive information, particularly when boundaries, sanitization, or access controls are weak [tran2025multi].
+  • Agent collusion and mode collapse. Collaboration mechanisms can produce groupthink, where agents reinforce the same flawed assumptions and converge on suboptimal outputs [feng2025multi].
   • Emergent behaviors.
-The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [102].
-Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability.
-Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [103], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals .
-These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
+Complex interactions among agents, tools, and memory can yield unpredictable strategies that evade conventional testing (e.g., shortcutting verification), motivating continuous monitoring and adaptive controls [dawid2025agentic].
 Real-World Examples
-Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
-  • Prompt Leakage in Agentic AI Systems.
-Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [104], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information.
-In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [105].
-Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks.
-  • Collusive Failure in ChatDev.
-In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [86].
-Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations.
-This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [37].
-  • Simulation Attack in Swarm Robotics.
-In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [106].
-This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences.
-The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [107].
-  • Memory Poisoning in Multi-Agent Chatbots.
-In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer.
-This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones.
-This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [108].
-  • System Prompt Drift in Autonomous Memory Agents.
-In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated.
-This led to hallucinated goals and emergent behaviors misaligned with initial intentions [109].
-Next, we discuss the AI TRiSM framework and map its core principles to a AMAS and how it can mitigate these aforementioned risks.
-5 The TRiSM Framework for Agentic Multi-Agent Systems
-5.1 Overview of AI TRiSM and Its Core Components
-The TRiSM Framework, short for Trust, Risk, and Security Management, offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors.
-In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
+The following examples illustrate how these risks can appear in practice:
+  • Prompt leakage.
+In AutoGPT-style setups, recursive prompt augmentation and weak memory controls can expose sensitive content when stored tokens later surface in summaries or logs [Euler2023AutoGPTRCE, invicti_prompt_injection_ebook].
+  • Collusive failure.
+In ChatDev-like workflows, shared planning errors can propagate when agents validate each other without external checks or objective feedback [qian2024chatdevcommunicativeagentssoftware].
+  • Simulation attack in swarms.
+Misleading environment assumptions can induce coordination failures (e.g., congestion and task incompletion) in multi-agent planning settings [StrDorFri2024, ligot2022using].
+  • Memory poisoning.
+Feedback buffers can be polluted (e.g., sarcastic or adversarial signals) and later used for policy updates, degrading behavior without validation gates [atkins2023memorypoisoning].
+  • System prompt drift. Self-appended context without versioning can shift prompts over time and induce hallucinated goals or misaligned behaviors [liu2023lostmiddlelanguagemodels].
+Key Insights
+We identified AMAS-specific risks that arise from autonomy, shared memory, orchestration, and tool use, and organized them into four emergent classes (Fig.
+10).
+Across the taxonomy, failures can cascade across agents and persist through memory substrates, while emergent behaviors may evade conventional test suites.
+These observations motivate defense-in-depth controls spanning prompt hygiene, memory sanitization, orchestrator authentication, tool sandboxing, and continuous monitoring.
+4 TRiSM Framework for Agentic Multi-Agent Systems
+4.1 Overview of AI TRiSM
+The TRiSM Framework offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. Below, we discuss core components of TRiSM framework.
   • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness.
-Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [54].
+Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [borghoff2025human].
 This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
   • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies.
-This includes technical risks such as algorithmic bias [41], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [11].
+This includes technical risks such as algorithmic bias [raza2024exploring], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [fang2025trustworthyaisafetybias].
 Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
   • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access.
-This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [38].
+This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [raza2025justhumansneedvaccines].
 Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
-Originally highlighted in industry guidelines for AI governance [110], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems.
-The TRiSM framework addresses trust concerns in Agentic AI by focusing on key pillars, which are: Explainability, ModelOps, Application Security, Model Privacy and Goverance [64, 13].
-Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
+Table 4: TRiSM pillars (Part 1): Core controls, techniques, and key risks mitigated.
+Pillar Core Controls Techniques / Patterns Key Risks Mitigated
+Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [sanwal2025layered]; LIME/SHAP [ribeiro2016should, lundberg2017unified]; decision-provenance graphs [tupayachi2024towards, bai2024derived]; multi-agent SHAP [tian2025outlook]; attention maps [liu2024ava, hu2024overview]; role-based interpretability [tomsett2018interpretable]; RAG-linked justifications [ghafarollahi2024sciagents]; prompt/action logs [gu2024artificial, chin2025automating] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration
+ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [wang2025megaagent] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps
+Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [beurer2025design, qiao2024agent]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [standen2025adversarial]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks
+Privacy & Data Protection DP; anonymization; pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [dwork2006calibrating]; k-anonymity [sweeney2002k]; MPC/PUMA [lindell2009secure]; HE/FHE pipelines [gentry2009fully]; TEEs [costan2016intel]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention
+Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions
+Table 5: TRiSM pillars (Part 2): Evaluation facets, example systems, and governing standards.
+Pillar Evaluation Facets & Metrics Example Systems Standards / Frameworks
+Explainability / Trust Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [raza_fair_2024] SciAgent [ghafarollahi2024sciagents]; MetaGPT [ke2025meta]; ChatDev [qian2023chatdev]; AutoGen [wu2023autogen] EU AI Act Arts 13–14 [eu_ai_act]; NIST AI RMF / GAI profile [ai2023artificial, ai2024artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 24029-1 [iso24029-1]
+ModelOps (Lifecycle) Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent [wang2025megaagent] ISO/IEC 42001 (AIMS); ISO/IEC 42005 [ISO_IEC_42005_2025]; ISO/IEC 23894 [ISOIEC23894_2023]; NIST AI RMF [ai2023artificial, ai2024artificial]; ModelOps [lefevre2022modelops, Sinha2025ModelOps]
+Application Security Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [chen2025human]; enterprise dashboards [chen2025human]; ChemCrow safety pauses [bran2023chemcrow] OWASP LLM Top-10 [owasp2024llmtop10, owasp2025agentic]; OECD Robustness [oecd_ai_catalogue]
+Privacy & Data Protection Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [gdpr25]; CCPA/CPRA [oag_ccpa_2024]; HIPAA §164 [hipaa164]; ISO/IEC 27001/27701
+Governance User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [eu_ai_act]; NIST AI RMF [ai2023artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 42001/42005; ISO/IEC 23894 [ISOIEC23894_2023]
+4.2 Mapping Pillars of TRiSM with Agentic AI
+Lifecycle Governance Plane (TRiSM Pillar) policies $\cdot$ accountability $\cdot$ compliance $\cdot$ oversight $\cdot$ traceability
+Trust & Explainability (TRiSM Pillar) reasoning/provenance traces $\cdot$ explainer agent $\cdot$ user feedback $\cdot$ audit logs ModelOps / Lifecycle Management (TRiSM Pillar) prompt+agent config versioning $\cdot$ CI/CD safety gates $\cdot$ monitoring/rollback $\cdot$ incident workflow Application Security (TRiSM Pillar) prompt hygiene $\cdot$ allowlists $\cdot$ tool sandboxing $\cdot$ least privilege (RBAC) $\cdot$ anomaly detection Model Privacy & Data Protection (TRiSM Pillar) encrypted/scoped memory $\cdot$ data minimization $\cdot$ DP where applicable $\cdot$ access logs $\cdot$ retention/consent Planning & Reasoning (Agent Core Layer; non-TRiSM) multi-agent planning $\cdot$ task decomposition $\cdot$ CoT/ToT/ReAct $\cdot$ inter-agent coordination
+User / HITL UI
+confirm $\cdot$ override $\cdot$ stop
+feedback loop
+Secure API Gateway
+authN/authZ $\cdot$ egress control
+tool isolation $\cdot$ policy checks
+External Tools
+web $\cdot$ DB $\cdot$ code
+enterprise APIs
 Figure 11: TRiSM-aligned architecture for Agentic AI systems.
-The system is organized into four layers: Trust and Explainability (e.g., reasoning traces, user feedback), ModelOps (prompt versioning, CI/CD, orchestration), Planning and Reasoning (multi-agent CoT), and Model Privacy (encrypted memory, differential privacy, audit logs).
-A cross-cutting Governance Plane enforces oversight via policies and logging.
-External tools are accessed through a secure API gateway.
-The user interface connects back to the trust and governance layers for human-in-the-loop (HITL) oversight.
-Abbreviations used: TRiSM = Trust, Risk, and Security Management; CoT = Chain-of-Thought; CI/CD = Continuous Integration and Deployment; DP = Differential Privacy.
-Table 4: TRiSM pillars for Agentic (LLM-based) multi-agent systems: core controls, techniques, risks, evaluation facets, example systems, and governing frameworks.
-Pillar Core controls (keywords) Techniques / patterns / artifacts Key risks mitigated Evaluation facets & example metrics Example systems / patterns Primary standards / frameworks
-Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [111]; LIME/SHAP [112, 113]; decision-provenance graphs [114, 115]; multi-agent SHAP [116]; attention maps [117, 118]; role-based interpretability [119]; RAG-linked justifications [120]; prompt/action logs [121, 122] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [123] SciAgent [120]; MetaGPT [124]; ChatDev [30]; AutoGen [57] EU AI Act Arts 13–14 [125]; NIST AI RMF / GAI profile [3, 126]; OECD AI [127]; ISO/IEC 24029-1 [32]
-ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [128] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent (hierarchical monitoring) [128] ISO/IEC 42001 (AIMS) [ISO_IEC_42001_2023]; ISO/IEC 42005 [129]; ISO/IEC 23894 [iso23894 or [130]]; NIST AI RMF / GAI profile [3, 126]; ModelOps [131, 132]
-Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [133, 58]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [134]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [135]; enterprise dashboards (override/stop) [135]; ChemCrow safety pauses [136] OWASP LLM Top-10 [137, 138]; OECD Robustness [127]
-Model Privacy & Data Protection DP; anonymization/pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [139]; k-anonymity [140]; MPC/PUMA [puma2024 or [141]]; HE/FHE pipelines [142]; TEEs [143]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [34]; CCPA/CPRA [144]; HIPAA §164 [35]; ISO/IEC 27001/27701 [iso27001, iso27701]
-Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [125]; NIST AI RMF [3]; OECD AI [127]; ISO/IEC 42001/42005 [align keys]; ISO/IEC 23894 [130]
-5.2 Mapping Pillars of TRiSM with Agentic AI
-In this section, we examine how the pillars of TRiSM align with Agentic AI (or AMAS) and summarize our syntheis in Table LABEL:tab:trism-pillars.
-For each pillar, we outline its relevance and implications within the context of Agentic AI.
+The five TRiSM pillars are shown explicitly and in the same order as defined in the framework: Trust & Explainability, ModelOps, Application Security, Model Privacy & Data Protection, with a cross-cutting Lifecycle Governance plane.
+Planning & Reasoning is included as the agent-core execution layer (non-TRiSM).
+Originally highlighted in industry guidelines for AI governance [litan2024ai], TRiSM also provides a structured approach to managing the unique challenges of AMAS.
+In this work, we focus on five key TRiSM pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance [kaur2024building, habbal2024artificial].
+These five pillars also reflect established risk-surface decompositions in the NIST AI RMF [ai2023artificial], ISO/IEC 42001 [iso42001], and the EU AI Act [eu_ai_act], where explainability, lifecycle management, security, privacy, and governance are treated as distinct and non-overlapping risk domains.
+We summarize this mapping in Tables LABEL:tab:trism-pillars-1 and LABEL:tab:trism-pillars-2.
+Figure 11 illustrates how these five TRiSM pillars map onto the core architectural layers of an LLM-based agentic AI system.
+Next, we discuss these pillars and how they can help in AMAS.
 4.2.1 Explainability in Multi-Agent Decision Making
-Explainability refers to making the inner workings and decisions of AI systems interpretable to humans [145].
-In AMAS settings, outcomes emerge from the interactions among specialized agents rather than a single model prediction [146].
-Therefore, explainability must cover both individual agent decisions and the inter-agent dynamics that lead to emergent results.
-Recent research emphasises that user trust hinges on understanding how each agent contributes and how their collaboration is orchestrated [111].
-Traditional single-model XAI techniques often ignore this multi-agent dimension, so new methods are required.
-Challenges Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information they used, and how their interactions influenced the final outcome [56].
-Additional challenges include: (1) Emergent behaviour: interactions among agents can lead to behaviour that was not explicitly programmed; explaining these emergent patterns requires capturing interaction chains.; (2) Verification of intermediate steps: vanilla CoT prompting in LLMs often produces plausible-sounding but unchecked reasoning.
-Without verification, small errors can propagate
-Recent directions Recent works introduce new frameworks and techniques for multi-agentic explainability:
-  • Layered Chain-of-Thought prompting: Instead of producing a single monolithic reasoning trace, Layered-CoT divides reasoning into discrete layers [111].
-Each layer answers a sub-question, and a specialized Reasoning Agent generates the partial chain-of-thought while Verification Agents and User-Interaction Agents cross-check the intermediate steps against knowledge graphs or human feedback.
-This agent-based decomposition localizes errors, prevents contradictions, and ensures that each partial solution is validated before the model proceeds.
-  • Integrated multi-agent workflows: Multi-agent workflows embed explainability into the process itself rather than as a post-hoc add-on.
-For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves relevant DSM-5 criteria and explicitly links conversational utterances to diagnostic rules.
-This integration produces transparent, step-by-step diagnostic predictions and shows why specific recommendations are made [147].
-The study notes that step-by-step reasoning improves perceptions of fairness and reliability and that embedding a diagnostician agent moves beyond general synthetic dialogues by providing a self-documenting rationale.
-  • Modular resume-screening systems: In recruitment, a multi-agent architecture divides the process into extraction, evaluation, summarization and formatting.
-The evaluation agent dynamically retrieves company-specific hiring criteria via Retrieval-Augmented Generation (RAG), and each stage remains clearly defined.
-This modularity allows recruiters to trace how a candidate was assessed and why a particular score was assigned [148].
-  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide individualized credit assignments to agents and explain its feedback strategy.
-Unlike earlier deep RL methods, this approach uses the language-based critic to justify why each agent receives a particular reward and to show how it decomposes the global reward [149].
-Approaches for explainability in multi-agent LLM systems The following techniques adapt traditional XAI methods to the multi-agent setting:
-  • Feature Attribution (LIME/ SHAP) for agents: Techniques like LIME [112] or SHAP [113] can be extended to attribute an agent’s output to specific input components (e.g., particular context tokens or inter-agent messages).
-Highlighting which parts of the input or which inter-agent signals most influenced an agent’s action helps illuminate why a specific agent behaved as it did [150].
-  • Counterfactual and Causal Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability.
-Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [151].
-For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes.
-This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
-  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency.
-Logging the step-by-step reasoning process – for example, an agent CoT prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [75].
-By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision.
-This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
-  • Natural Language Explanations: Multi-agent systems can include a separate explainer agent that synthesizes logs and feature attributions into human-readable narratives.
-This agent (often another LLM) translates the technical reasoning traces into conversational explanations that answer questions like “Why did the agents choose action X?” Integrated workflows, such as DSM5AgentFlow [147], demonstrate that embedding such explanation directly in the pipeline improves transparency and trust.
-  • Role-oriented modular architectures: Assigning clear roles to agents (e.g., reasoning, verification, summarization, credit assignment) enhances explainability because each stage of decision-making is distinct and can be inspected independently [148].
-Modular designs also allow domain experts to adjust or replace specific agents without affecting the entire system.
-  • Explainable reward critics: In multi-agent reinforcement learning, LLM-based critics can decompose global rewards and provide natural-language justifications, making credit assignment transparent [149] .
-This approach helps developers understand how credit is distributed across agents and why certain behaviors are encouraged or discouraged.
+In AMAS, explainability is less about interpreting a single model and more about understanding how decisions emerge from interacting agents [yossef2024explainable].
+In AMAS settings, outcomes often emerge from interactions among specialized agents rather than from a single model prediction [rosenfeld2019explainability].
+As a result, explainability must account for both (i) individual agent decisions and (ii) the inter-agent dynamics that produce emergent system-level behavior.
+Recent work emphasizes that user trust depends on understanding each agent’s contribution and how collaboration is orchestrated [sanwal2025layered].
+Traditional single-model XAI methods often overlook this multi-agent dimension, motivating new approaches tailored to AMAS.
+Challenges.
+Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information it used, and how inter-agent exchanges influenced the final outcome [de2025open].
+Key challenges include:
+  • Emergent behavior: agent interactions can produce behaviors that were not explicitly specified; explaining these patterns requires capturing and summarizing interaction chains.
+  • Verification of intermediate steps: vanilla CoT prompting can yield plausible but unverified reasoning.
+Without systematic verification, small errors may propagate across agents and compound into incorrect final decisions.
+Approaches for explainability in AMAS The following techniques adapt classical XAI ideas to AMAS:
+  • Feature attribution (LIME/SHAP) for agents: LIME [ribeiro2016should] and SHAP [lundberg2017unified] can be extended to attribute an agent’s output to specific inputs (e.g., context spans, retrieved evidence, or inter-agent messages), clarifying which signals most influenced an agent’s action [bilal2025llms].
+  • Counterfactual and causal analysis: “What-if” analysis evaluates how outcomes change when an agent’s contribution or a key condition is modified [gyevnar2023causal].
+This helps isolate agent roles and identify causal dependencies in collaborative decision-making.
+  • Reasoning-trace logging: Logging intermediate reasoning artifacts (e.g., CoT traces, tool calls, and inter-agent dialogues) creates an auditable trail of how decisions emerged [yao2023reactsynergizingreasoningacting].
+Such traces reveal which agent contributed what and how partial conclusions accumulate into final outputs.
+  • Natural-language explanations: An explainer agent can synthesize traces and attributions into human-readable narratives that answer questions such as “Why did the system choose action X?” Embedded explainers in workflows (e.g., DSM5AgentFlow) can improve transparency and trust [ozgun2025trustworthy].
+  • Role-oriented modular architectures: Assigning explicit roles (e.g., reasoning, verification, summarization, credit assignment) improves inspectability because each stage is separable and reviewable [lo2025ai].
+  • Explainable reward critics: In multi-agent RL, LLM-based critics can provide transparent reward decomposition and justification, supporting debugging and governance of agent incentives [nagpal2025leveraging].
+Recent directions. Recent work proposes frameworks and techniques that embed explainability directly into multi-agent operation:
+  • Layered Chain-of-Thought prompting: Layered-CoT decomposes reasoning into discrete layers, where each layer answers a sub-question [sanwal2025layered].
+A specialized Reasoning Agent produces partial reasoning, while Verification and User-Interaction Agents cross-check intermediate outputs (e.g., against knowledge graphs or human feedback).
+This decomposition can localize errors, reduce contradictions, and promote validation before proceeding.
+  • Integrated multi-agent workflows: Some systems integrate explanation into the workflow rather than adding it post hoc.
+For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves DSM-5 criteria and explicitly links conversational evidence to diagnostic rules, producing a self-documenting rationale [ozgun2025trustworthy].
+  • Modular resume-screening systems: In recruitment, multi-agent pipelines separate extraction, evaluation, summarization, and formatting.
+The evaluation agent retrieves company-specific criteria via RAG, enabling recruiters to trace how candidate evidence maps to scores [lo2025ai].
+  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide per-agent credit assignment and natural-language justification, explaining how global rewards are decomposed across agents [nagpal2025leveraging].
 4.2.2 ModelOps : Lifecycle Management for LLM Agents
-ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management and continuous oversight for all types of AI models [131].
-Unlike MLOps, ModelOps manages not only individual models but also composite systems of agents and heterogeneous models across different business units.
-ModelOps provides end-to-end oversight for AI in production by enforcing model governance and compliance: tracking lineage, ownership, and metadata to meet regulations such as GDPR and HIPAA , and by using policy-driven deployment gates so only approved models go live, with controlled roll-outs and roll-backs [132].
-It continuously monitors risk and bias, detecting drift, fairness issues, hallucinations, and robustness regressions.
-ModelOps also manages heterogeneous assets: rule-based systems, LLMs, and composite multi-agent architectures; ensuring consistent visibility and lifecycle control across the portfolio.
-Finally, it integrates tightly with enterprise IT, wiring AI workflows into CI/CD, IT service management (ITSM), identity and access management (IAM), and cloud-security tooling [132].
-For generative AI and AMAS, ModelOps must handle prompt versioning, guardrails and content filtering enforcement, fine-tuning lineage tracking, usage and cost observability, and ethical compliance.
-Effective ModelOps thus requires:
-  • Version control and CI/CD pipelines: Track versions of each agent’s model and prompt configuration.
-Automated pipelines should test safety, performance and cost impacts whenever an agent’s logic or model is updated, using multi-agent simulations to validate that new behaviours do not introduce regressions or unsafe interactions.
-  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [128] decompose tasks into subtasks and dynamically form groups of agents.
-They employ hierarchical monitoring, where each agent tracks its actions with a checklist, group-level administrators oversee their agents, and system-level supervisors review outputs.
-This layered monitoring ensures that changes to one agent do not compromise the entire system.
-  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks and updating measurement approaches regularly; it suggests using gradient-based attributions, occlusion and counterfactual prompts to improve explainability and transparency, and establishing policies for oversight across the lifecycle [126].
+ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management, and continuous oversight across AI systems [lefevre2022modelops].
+In production, ModelOps provides end-to-end oversight by enforcing governance and compliance controls, including lineage, ownership, and metadata tracking to support regulatory requirements (e.g., GDPR and HIPAA), as well as policy-driven deployment gates so that only approved models go live with controlled rollouts and rollbacks [Sinha2025ModelOps].
+It also enables continuous risk monitoring to detect drift, bias and fairness issues, hallucinations, and robustness regressions.
+For AMAS, ModelOps must additionally support prompt and toolchain versioning, enforcement of guardrails and content filtering, fine-tuning lineage tracking, usage and cost observability, and ethical compliance.
+Effective ModelOps for LLM agents therefore requires:
+  • Version control and CI/CD pipelines: Track versions of each agent’s model, prompts, tools, and configuration.
+Automated pipelines should test safety, performance, and cost impacts whenever agent logic changes, using multi-agent simulations to validate that updates do not introduce regressions or unsafe interactions.
+  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [wang2025megaagent] decompose tasks into subtasks and dynamically form agent groups.
+They apply hierarchical monitoring, where agents log actions (e.g., via checklists), group-level administrators oversee agent teams, and system-level supervisors review final outputs.
+This layered oversight helps ensure that changes in one agent do not compromise system-wide behavior.
+  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks, updating measurement approaches regularly, and establishing lifecycle oversight policies.
+It also highlights techniques such as gradient-based attributions, occlusion tests, and counterfactual prompting to strengthen transparency and explainability [ai2024artificial].
 4.2.3 Application Security for LLM Agents
-Security in AMAS systems must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools.
-According to OWASP [137, 138], prompt injection manipulates an LLM through malicious instructions, potentially causing the model to leak data, execute unauthorized commands or make harmful decisions.
-Multimodal models face cross-modal injection where malicious content in one modality influences the output of another.
+Security in AMAS must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools.
+OWASP identifies prompt injection as a core risk in LLM applications: attackers manipulate the model with malicious instructions, potentially causing data leakage, unauthorized tool use, or unsafe decisions [owasp2024llmtop10, owasp2025agentic].
+Multimodal models further face cross-modal injection, where malicious content in one modality (e.g., text in an image) influences behavior in another.
 To mitigate these risks:
-  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure prefixes, and validate instructions before they reach an agent.
-Constraining the output format and employing guardrails can limit the model’s response surface [38].
-  • Defense-in-depth design patterns: Existing techniques such as adversarial training and user confirmation provide partial protection, but research shows they do not guarantee safety.
-Design patterns [133] like Action-Selector and Plan-Then-Execute restrict an agent’s capabilities after processing untrusted input; the agent selects from a predefined list of allowed actions or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [58].
-  • Least-privilege access and authentication: Enforce strong authentication for both human users and agents and apply role-based access controls so agents can only invoke the minimum set of tools necessary to perform their task.
-Continuous monitoring should detect anomalous requests; if an agent deviates from normal patterns, automated systems should flag or suspend it [152].
-  • Post-processing and anomaly detection: Implement filters to detect and remove sensitive or inaccurate content from outputs and regularly retrain models to reduce hallucinations.
-Maintain robust logging and retention policies so that investigators can audit interactions without exposing private data.
-Recent frameworks also emphasis hierarchical monitoring and cross-checking among agents to detect inconsistencies [128].
-LangChain [153], AutoGen [98], and CrewAI [87] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise.
-Training agents with adversarial examples [134] and red-teaming [154] scenarios further strengthens robustness against malicious inputs.
+  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure system prefixes, and validate instructions before they reach an agent.
+Constraining output formats and enforcing guardrails can reduce the attack surface [raza2025justhumansneedvaccines].
+  • Defense-in-depth design patterns: Techniques such as adversarial training and user confirmation provide partial protection, but do not guarantee safety.
+Design patterns [beurer2025design] such as Action-Selector and Plan-Then-Execute restrict capabilities after untrusted input is processed: the agent selects from a predefined set of allowed actions, or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [qiao2024agent].
+  • Least-privilege access and authentication: Enforce strong authentication for both users and agents, and apply role-based access control so agents can invoke only the minimum set of tools required for their task.
+Continuous monitoring should detect anomalous requests; if an agent deviates from expected behavior, automated systems should flag or suspend it [barbera2025ai].
+  • Post-processing and anomaly detection: Apply output filters to detect sensitive or inaccurate content, and update or retrain models to reduce hallucinations.
+Maintain robust logging and retention policies to support audits and incident response while minimizing exposure of private data.
+Recent frameworks also emphasize hierarchical monitoring and cross-checking among agents to detect inconsistencies [wang2025megaagent].
+Tooling ecosystems such as LangChain [langchain2024function], AutoGen [autogen_agents_tutorial], and CrewAI [crewAI] operationalize verification through agent-level checks (e.g., reviewers/critics) and, in some cases, trust or reputation signals that encourage agents to validate each other’s outputs and flag potential compromise.
+Finally, adversarial training with attack examples [standen2025adversarial] and systematic red-teaming [ganguli_red_2022] can further strengthen robustness against malicious inputs.
 4.2.4 Model Privacy and Data Protection
-AMAS also poses unique privacy challenges because agents often share information via memory stores and message passing.
-Without strict controls, an agent may inadvertently leak sensitive data [40].
-A layered privacy strategy should include:
-  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re-identified and reduces the risk that models memorize or regurgitate personal information [139].
-DP should be assessed on a case-by-case basis, since heavy noise can harm accuracy.
-  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [140].
+Without strict controls, an agent may inadvertently leak sensitive data [ko2025sevensecuritychallengessolved]. A layered privacy strategy should include:
+  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re‑identified and reduces the risk that models memorize or regurgitate personal information [dwork2006calibrating].
+DP should be assessed on a case‑by‑case basis, since heavy noise can harm accuracy.
+  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [sweeney2002k].
 Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
-  • Secure Multi-Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs.
-In privacy-preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7-billion-parameter model within minutes) [141].
+  • Secure Multi‑Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs.
+In privacy‑preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7‑billion‑parameter model within minutes) [lindell2009secure].
   • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy.
-Recent work demonstrates an encryption-friendly LLM architecture that uses LoRA fine-tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [142].
-  • Trusted Execution Environments (TEEs): TEEs provide hardware-enforced isolation that protects code and data [143].
+Recent work demonstrates an encryption‑friendly LLM architecture that uses LoRA fine‑tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [gentry2009fully].
+  • Trusted Execution Environments (TEEs): TEEs provide hardware‑enforced isolation that protects code and data [costan2016intel].
 They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone.
@@
   • Continuous monitoring and access controls: Encrypt data at rest and in transit, restrict access to authorized personnel, and use monitoring to detect data poisoning or unusual behaviour.
-These techniques, combined with careful consent management and compliance with local regulations, help maintain confidentiality when agents share or process sensitive information.
+4.2.5 Governance
+Governance integrates ethical, legal, and societal considerations into the design and operation of AI systems. Below, we summarize how major governance frameworks inform the TRiSM principles.
+  • EU AI Act — Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety, or fundamental rights [EU_AI_Act2025].
+Oversight mechanisms should enable users to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop operation (e.g., via a “stop” mechanism).
+In certain high-risk identification contexts, decisions may require independent verification by multiple competent individuals.
+  • OECD AI Principles: These intergovernmental principles (updated in 2024) promote human-centric, trustworthy AI [oecd_ai_catalogue].
+They emphasize transparency and explainability: AI actors should provide meaningful information about data sources, factors, and logic so affected individuals can understand and, where appropriate, challenge outputs.
+Robustness, security, and safety require systems to function appropriately under normal and adverse conditions, including mechanisms to override or decommission systems to prevent harm.
+Accountability requires that AI actors ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
+  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile define four core functions: Govern, Map, Measure, and Manage [ai2023artificial].
+For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (e.g., gradient-based attributions, occlusion tests, and counterfactual prompting) to improve transparency, and establishing policies for oversight across the AI lifecycle.
+  • Auditability and role-based access: AMAS should maintain auditable logs of agent actions, decisions, and data flows, and assign clear roles (e.g., reasoning, verification, summarization) so responsibilities remain transparent [khan2024role].
+Role-based access control (RBAC) helps enforce accountability and prevents unauthorized agents from invoking sensitive functions.
+  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data-subject rights, consent, and privacy by design.
+Frameworks such as the OECD Principles and NIST AI RMF also promote international cooperation and interoperability.
+These governance frameworks (Figure 12) highlight the need for clear accountability structures, human oversight, robust privacy controls, and documented processes in AMAS.
+TRiSM provides a practical scaffolding to operationalize these requirements by aligning lifecycle management, security, privacy, and governance to support trustworthy agentic AI.
 Figure 12: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST AI RMF, ISO/IEC 42001, 42005, and OECD Principles.
@@
 Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art.
-13–14) [125] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [3]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [155].
-Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [119].
-ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [3] and ISO/IEC 42001:2023 (AI management system) [156]; secure SDLC controls via NIST SP 800-218/218A [^5] NIST, 2024.
+13–14) [eu_ai_act] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [ai2023artificial]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [phillips2021four].
+Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [tomsett2018interpretable].
+ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [ai2023artificial] and ISO/IEC 42001:2023 (AI management system) [thiers2024emerging]; secure SDLC controls via NIST SP 800-218/218A [^9] NIST, 2024.
 Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates.
 Risk Management & Impact Assessment: EU AI Act (Art.
-9, 72–73) [125], ISO/IEC 23894 (AI risk management) [^6] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [129].
+9, 72–73) [eu_ai_act], ISO/IEC 23894 (AI risk management) [^10] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [ISO_IEC_42005_2025].
 Iterative hazard analysis, impact registers, incident response, and disclosure playbooks.
 Data Governance & Quality: EU AI Act (Art.
-10) [125] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^7] ISO/IEC 5259 series.
+10) [eu_ai_act] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^11] ISO/IEC 5259 series.
 Data minimization, lineage, coverage/bias checks, and sampling documentation.
 Robustness & Safety: EU AI Act (Art.
-15) [125] and ISO/IEC 24029-1 (robustness assessment) [32].
+15) [eu_ai_act] and ISO/IEC 24029-1 (robustness assessment) [iso24029-1].
 Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria.
-Application Security: Guided by OWASP LLM Top-10 [137] and OECD AI Principles on robustness [127].
+Application Security: Guided by OWASP LLM Top-10 [owasp2024llmtop10] and OECD AI Principles on robustness [oecd_ai_catalogue].
 Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails.
 Model Privacy: Aligned with GDPR (Art.
-25) [34], HIPAA §164 [35], and OECD AI Principles [127].
+25) [gdpr25], HIPAA §164 [hipaa164], and OECD AI Principles [oecd_ai_catalogue].
 Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds.
 Documentation & Traceability: EU AI Act (Arts.
-11–12; Annex IV) [125].
+11–12; Annex IV) [eu_ai_act].
 Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages.
 Post-Market Monitoring: EU AI Act (Arts.
-72–73) [125].
+72–73) [eu_ai_act].
 Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications.
-Governance: Grounded in GDPR, CCPA [144], and OECD AI Principles [127].
+Governance: Grounded in GDPR, CCPA [oag_ccpa_2024], and OECD AI Principles [oecd_ai_catalogue].
 Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments.
-Fairness & Human Rights: OECD AI Principles [127] and ISO/IEC TR 24027 (bias) [^8] ISO/IEC TR 24027.
+Fairness & Human Rights: OECD AI Principles [oecd_ai_catalogue] and ISO/IEC TR 24027 (bias) [^12] ISO/IEC TR 24027.
 Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
-5.2.5 Governance
-Governance integrates ethical, legal and societal considerations into the design and operation of AI systems.
-Regulations such as the EU AI Act [125], NIST AI RMF [3] and OECD AI [127] provide guidance that aligns with TRiSM pillars of Explainability, ModelOps, Security and Privacy.
-  • EU AI Act - Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety or fundamental rights [1].
-The AI system should enable overseers to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop its operation through a ‘stop’ button.
-For certain identification tasks, decisions must be independently verified by at least two competent individuals.
-  • OECD AI Principles: These intergovernmental principles, updated in 2024, promote human-centric, trustworthy AI [127].
-They emphasise transparency and explainability: AI actors should provide meaningful information about data sources, factors and logic to allow affected individuals to understand and challenge outputs.
-Robustness, security and safety require AI systems to function appropriately under normal and adverse conditions and include mechanisms to override or decommission systems to prevent harm.
-Accountability means AI actors are responsible for the proper functioning of AI systems; they must ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
-  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile outline four core functions: govern, map, measure and manage [3].
-For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (gradient-based attributions, occlusion, counterfactual prompts) to improve transparency, and establishing policies for oversight across the AI lifecycle.
-  • Auditability and role-based access: Agentic AI systems should maintain auditable logs of agent actions, decisions and data flows and assign clear roles (e.g., reasoning, verification, summarisation) so that responsibilities are transparent [157].
-Role-based access controls help enforce accountability and prevent unauthorized agents from accessing sensitive functions.
-  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data subject rights, consent and privacy by design.
-Frameworks like the OECD Principles and NIST AI RMF promote international cooperation and interoperability.
-These governance frameworks (as illustrated in Figure 6) underscore the need for transparent accountability structures, human oversight, robust privacy controls, and documented processes in AMAS.
-TRiSM provides a practical scaffolding for operationalizing these requirements, aligning lifecycle management, security, privacy and governance to build trustworthy Agentic AI.
+Key Insights
+This section operationalized the TRiSM framework for AMAS across five pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance.
+Multi-agent explainability requires novel techniques including Layered CoT decomposition, decision provenance graphs, and role-based interpretability frameworks to trace distributed decision processes.
+ModelOps extends lifecycle management through hierarchical monitoring, prompt versioning, and CI/CD safety gates with pre-deployment simulation.
+Security strategies emphasize defense-in-depth via design patterns (Plan-Then-Execute), least-privilege access control, and cross-agent verification.
+Privacy preservation leverages differential privacy, secure multi-party computation, homomorphic encryption, and trusted execution environments to protect sensitive data during inter-agent communication.
+Governance integration with regulatory frameworks (EU AI Act, NIST AI RMF, ISO/IEC 42001) ensures that TRiSM implementation satisfies transparency, accountability, and human oversight requirements for high-stakes deployments.
 Next, we discuss the evaluations in AMAS.
-6 Evaluation of Agentic AI Systems
-Agentic AI systems demand evaluation beyond traditional accuracy.
-We outline five categories: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct facet of performance and real-world impact.
-Compared to a related work [158], which focuses on proportional impacts across general AI, our work targets Agentic AI and links risks to their causal sources (e.g., prompt injection, memory poisoning).
-We further introduce an evaluation template with new metrics the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE) to capture multi-agent dynamics.
-A summary appears in Table 5.
-Trustworthiness.
-This criterion evaluates reliability, safety, and ethical alignment [159].
-Core signals include task success rates under distributional shifts (robustness), violation rate of safety/ethical rules (lower is better), and calibration quality (e.g., Expected Calibration Error, ECE).
-To avoid pathological behavior in scoring, we define a normalized trustworthiness index
-$T\;=\;\frac{w_{C}\,C+w_{R}\,R+w_{I}\,I}{\,1+\lambda\,S\,}\,$, (1)
-where $C$ (credibility/accuracy), $R$ (reliability over time), and $I$ (user alignment/rapport) are normalized to $[0,1]$, $S\!\geq\!0$ is self-orientation (higher is worse), $w_{C},w_{R},w_{I}\!\geq\!0$ with $w_{C}{+}w_{R}{+}w_{I}\!=\!1$, and $\lambda\!\geq\!0$ controls how strongly self-orientation penalizes trust.
-Explainability.
-Explainability metrics assess how well human evaluators can understand and trace decisions.
-Useful measures include coverage (fraction of outputs with an explanation), fidelity/faithfulness (agreement with the model’s actual decision basis), stability (consistency across similar cases), and human interpretability ratings.
-Suites like OpenXAI [160] report faithfulness, stability, and fairness of explanations.
-High explainability builds user trust and aids debugging; in regulated domains (e.g., healthcare, finance) it is often essential for compliance.
-User-Centered Performance.
-User-centered metrics capture how effectively the agent satisfies end-user needs [161].
-Signals include post-interaction satisfaction scores, goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness.
-Human-in-the-loop (HITL) studies typically rate helpfulness, clarity, and instruction adherence.
-Ultimately, a user-centered agent aligns actions with user intent and preferences.
-Table 5: Summary of key evaluation metrics for agentic AI systems. Each category addresses a distinct facet of performance and behavior.
-Metric Aspect Evaluation Focus (Examples of Metrics)
-Trustworthiness Reliability, safety, and alignment.
-Examples: success rate under shifts; safety/ethics violation rate (lower is better); calibration (ECE/Brier  [^9] https://en.wikipedia.org/wiki/Brier_score); fairness indices; red-team/jailbreak success (lower is better).
-Explainability Transparency and traceability. Examples: explanation coverage; faithfulness/fidelity; stability across similar cases; human interpretability scores; explanation time/cost.
-User-Centered Experience and outcomes.
-Examples: satisfaction (Likert/CSAT  [^10] https://www.ibm.com/think/topics/csat-customer-satisfaction-score); goal-completion rate; clarification turns; perceived coherence/naturalness; task time; cognitive load (NASA-TLX).
-Coordination Collaboration in multi-agent or modular setups.
-Examples: team task completion; coordination overhead (messages/tokens, rounds to converge); plan consistency; deadlock/conflict rate (lower is better); synergy score for complementary actions.
-Composite Holistic summaries. Examples: weighted aggregate across categories (weights sum to 1); Tool-Use Efficacy (TUE, see definition); cross-domain benchmark aggregates.
-Coordination.
-In multi-agent or modular systems, coordination captures whether components maintain a shared plan, minimize redundant work, and resolve dependencies.
-Quantify with team success rate, communication efficiency (messages/tokens and rounds to consensus), plan/belief consistency, deadlock/conflict rate, and a Component Synergy Score (CSS) that credits inter-agent actions which enable or improve peers’ outcomes (higher is better).
-Composite Metrics.
+5 TRiSM Evaluation of Agentic AI Systems
+Accuracy alone is insufficient for Agentic AI, where failures often arise from coordination breakdowns, unsafe tool use, or misaligned autonomy rather than incorrect predictions.
+We organize evaluation into five dimensions: trustworthiness, explainability, user-centered performance, coordination, and composite reporting, each capturing a distinct facet of real-world performance and impact.
+We summarize these dimensions in Table 6.
+Trustworthiness
+This dimension evaluates reliability, safety, and ethical alignment [liu2023gevalnlgevaluationusing].
+Core signals include success under distributional shift (robustness), safety/ethical violation rate (lower is better), and calibration quality (e.g., ECE/Brier).
+To avoid pathological scoring, we define a normalized trustworthiness index:
+$T\;=\;\frac{w_{\text{acc}}\,A+w_{\text{rob}}\,R+w_{\text{align}}\,L}{1+\lambda\,V}\,$, (1)
+where $A$ is calibrated task success (accuracy adjusted for calibration), $R$ is robustness under shifts, $L$ is alignment/safety compliance (e.g., 1violation rate1{-$\text${violation rate}), all normalized to $[0,1]$, and $V\!\geq\!0$ is a normalized violation/self-serving penalty term (higher is worse).
+The weights satisfy $w_{\text{acc}}{+}w_{\text{rob}}{+}w_{\text{align}}=1$ and $\lambda\!\geq\!0$ controls how strongly violations penalize trust.
+Explainability
+Explainability metrics assess how well humans can understand and trace decisions.
+Useful measures include coverage (fraction of outputs with explanations), faithfulness/fidelity (agreement with the model’s actual decision basis), stability (consistency under small perturbations), and human interpretability ratings.
+Suites such as OpenXAI [agarwal2024openxaitransparentevaluationmodel] report faithfulness, stability, and fairness-related properties of explanations.
+High explainability supports debugging and, in regulated domains (e.g., healthcare and finance), is often necessary for compliance.
+User-Centered Performance
+User-centered metrics capture how effectively the agent satisfies end-user needs [rodden2010measuring].
+Signals include post-interaction satisfaction (e.g., CSAT), goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness.
+Human-in-the-loop (HITL) studies commonly rate helpfulness, clarity, and instruction adherence.
+A user-centered agent aligns actions with user intent and preferences.
+Table 6: Summary of key evaluation dimensions for Agentic AI systems.
+Aspect Focus Strengths Limitations
+Trustworthiness Reliability, safety, alignment (OOD success, safety violations, calibration, fairness).
+Deployment readiness; regulatory alignment (EU AI Act); statistically grounded; captures safety failures.
+Assumes stationarity; fairness trade-offs across groups [raza_fair_2024]; incomplete edge-case coverage.
+Explainability Transparency and traceability (coverage, fidelity, stability, human interpretability).
+Supports debugging and compliance [ai2023artificial]; improves calibrated trust; standardized benchmarks (OpenXAI [agarwal2024openxaitransparentevaluationmodel]).
+Expensive fidelity checks; faithful $\neq$ interpretable; weak multi-agent attribution.
+User-Centered User experience and outcomes (CSAT, task success, clarification turns, NASA-TLX).
+Reflects real-world utility; identifies usability bottlenecks; validated human factors measures.
+Subjective and high variance; costly user studies; limited scalability.
+Coordination Multi-agent collaboration (task success, communication cost, plan consistency, deadlocks).
+Captures emergent AMAS behavior; measures coordination efficiency; decomposable contribution metrics.
+No standard benchmarks; sensitive to orchestration; may penalize useful redundancy.
+Composite Aggregate system performance (weighted scores, tool-use efficacy).
+Single actionable summary; enables cross-system comparison; supports thresholding.
+Subjective weighting; can mask single-dimension failures; prone to Goodhart effects.
+Coordination
+In multi-agent or modular systems, coordination measures whether components maintain a shared plan, minimize redundant work, and resolve dependencies.
+Common signals include team success rate, communication efficiency (messages/tokens and rounds-to-consensus), plan/belief consistency, and deadlock/conflict rate.
+We additionally report: (i) a Component Synergy Score (CSS) to credit enabling contributions across agents, and (ii) Tool Utilization Efficacy (TUE) to assess correctness and efficiency of tool calls.
+Composite Reporting
 Composite metrics are often presented as a weighted scalar summary across dimensions:
-$M_{\text{composite}}\;=\;w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,$, (2)
-where $M_{T},M_{E},M_{U},M_{C}\in[0,1]$ are normalized category scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}\!\geq\!0$ with $\sum w_{\bullet}=1$ encode domain priorities.
-While this provides a scalar index for ease of interpretation, it risks masking over-performance in one dimension with under-performance in another.
-To address this, we treat the evaluation more rigorously as a vector of metrics,
-$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{C})\,$, (3)
-and recommend selection based on $\arg\max$ over $\mathbf{M}$ rather than on any scalarization:
-$\arg\max_{i}\;\mathbf{M}_{i}\,.$ (4)
-In this view, Eq.
-2 can be interpreted as a limiting case, useful for communicating results to non-technical stakeholders, while Eq.
-3 and Eq.
-4 form the principled basis for evaluation and comparison.
-Definition (Tool-Use Efficacy, TUE). In this work, we introduce Tool-Use Efficacy (TUE) as a specialized composite for evaluating how correctly and efficiently an agent uses external tools. Let
-[…]
-[…]
-$\mathrm{Exec}$ $\in[0,1]$ $\text{(execution success: no exceptions/timeouts, correct API usage)},$
-$\mathrm{Out}$ $\in[0,1]$ $\text{(post-call outcome correctness on tool-gated sub-tasks)},$
-$\mathrm{Eff}$ $\in[0,1]$ $\text{(efficiency: normalized inverse cost for latency/tokens/\#calls)}.$
-We define two aggregators (both reportable; choose per application):
-(a) Linear aggregator (auditable weights).
-$\mathrm{TUE}_{\mathrm{lin}}\;=\;\alpha_{\!s}\,\mathrm{Sel}+\alpha_{\!a}\,\mathrm{Arg}+\alpha_{\!e}\,\mathrm{Exec}+\alpha_{\!o}\,\mathrm{Out}+\alpha_{\!f}\,\mathrm{Eff}\,,\quad\alpha_{\!*}\!\geq\!0,\;\sum\alpha_{\!*}=1.$ (5)
-(b) Geometric aggregator (“all-of-the-above” strictness).
-$\mathrm{TUE}_{\mathrm{geo}}\;=\;\Big{(}\mathrm{Sel}^{\beta_{s}}\,\mathrm{Arg}^{\beta_{a}}\,\mathrm{Exec}^{\beta_{e}}\,\mathrm{Out}^{\beta_{o}}\,\mathrm{Eff}^{\beta_{f}}\Big{)}^{\!1/(\beta_{s}{+}\beta_{a}{+}\beta_{e}{+}\beta_{o}{+}\beta_{f})},\quad\beta_{\!*}\!>\!0.$ (6)
-Overhead penalty. If the preference is a separate cost term, then we can define normalized overhead $\mathrm{OH}\in[0,1]$ (increasing in time/tokens/calls) and report the penalized score
-$\mathrm{TUE}^{\star}\;=\;\frac{\mathrm{TUE}}{\,1+\lambda\,\mathrm{OH}\,}\,,\quad\lambda\!\geq\!0,$ (7)
-where $\mathrm{TUE}$ is either (5) or (6). It is recommended to always report the components (Sel, Arg, Exec, Out, Eff) with confidence intervals to ensure reproducibility.
+$M_{\text{comp}}\;=\;\omega_{T}M_{T}+\omega_{E}M_{E}+\omega_{U}M_{U}+\omega_{Co}M_{Co}\,$, (2)
+where $M_{T},M_{E},M_{U},M_{Co}\in[0,1]$ are normalized scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $\omega_{\bullet}\geq 0$ with $\sum\omega_{\bullet}=1$ encode domain priorities.
+Since scalarization can hide critical failures, we recommend always reporting the full metric vector:
+$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{Co})\,.$ (3)
+Model selection should then be based on (i) minimum thresholds (e.g., $M_{T}\geq\tau_{T}$ and $M_{Co}\geq\tau_{Co}$) and/or (ii) Pareto-optimality (non-dominated models) rather than relying only on a single weighted score.
+Component Synergy Score (CSS)
+Definition: CSS quantifies the quality of inter-agent collaboration by measuring how effectively one agent’s actions enable or enhance a peer’s downstream performance.
+Formal expression:
+$\text{CSS}=\frac{1}{N_{\text{pairs}}}\sum_{i}\sum_{j\neq i}\Big[\text{Impact}(a_{i}\!\rightarrow\!a_{j})\cdot\text{Quality}(a_{j}\mid a_{i})\Big],$ (4)
+where $N_{\text{pairs}}$ is the number of ordered agent pairs $(i,j)$ included in the evaluation.
+$\text{Impact}(a_{i}\!\rightarrow\!a_{j})$ measures how $a_{i}$ changes $a_{j}$’s subsequent performance (e.g., relative improvement), and $\text{Quality}(a_{j}\mid a_{i})$ is the quality of $a_{j}$’s output conditioned on $a_{i}$’s contribution.
+Illustrative example: In a software AMAS with Planner, Coder, and Tester:
+  • Planner’s decomposition speeds up Coder by 35% ($\text{Impact}=0.35$).
+  • Tester achieves 92% coverage conditioned on Coder’s output ($\text{Quality}=0.92$).
+If we score only this link, then $\text{CSS}=0.35\times 0.92\approx 0.32$ (and increases as additional enabling links are included).
+Tool Utilization Efficacy (TUE)
+Definition: TUE evaluates how correctly and efficiently agents invoke external tools (APIs, databases, code interpreters) within their workflows.
+Components:
+  • $\text{Sel}\in[0,1]$ — Selection (decision precision/recall for tool use)
+  • $\text{Arg}\in[0,1]$ — Argument validity (well-formed parameters)
+  • $\text{Exec}\in[0,1]$ — Execution success (no errors/timeouts)
+  • $\text{Out}\in[0,1]$ — Outcome correctness (downstream task correctness)
+  • $\text{Eff}\in[0,1]$ — Efficiency (normalized inverse of latency/token/call cost)
+Linear aggregator:
+$\text{TUE}_{\text{lin}}=\alpha_{s}\text{Sel}+\alpha_{a}\text{Arg}+\alpha_{e}\text{Exec}+\alpha_{o}\text{Out}+\alpha_{f}\text{Eff},$ (5)
+where $\alpha_{\bullet}\geq 0$ and $\sum\alpha_{\bullet}=1$.
+Key Insights
+This section defined a multi-dimensional evaluation framework covering trustworthiness (robustness, violation rates, calibration), explainability (coverage, faithfulness, stability), user-centered performance (satisfaction and goal fulfillment), and coordination quality.
+CSS captures positive inter-agent enablement, while TUE decomposes tool use into selection, argument validity, execution success, outcome correctness, and efficiency.
+Finally, the framework emphasizes reporting a metric vector and using thresholding/Pareto analysis to reduce the risk that scalar composites mask critical single-dimension failures.
 6 Security and Privacy in Agentic AI Systems
-In this section, we discuss the security and privacy in AMAS.
+Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness.
+This subsection reviews current evaluation frameworks and security metrics for agentic AI, aligned with the TRiSM pillars.
 6.1 Security Mechanisms
-Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [162].
-Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [163].
-Among the foundational techniques employed are encryption , access control , adversarial defense , and runtime monitoring, each adapted to the unique demands of decentralized multi-agent environments.
+Agentic AI systems (multiple collaborating LLM agents) have a bigger attack surface than traditional agents, so they need layered defenses: protecting data, execution, communication, and model robustness; using encryption, access control, adversarial defenses, and runtime monitoring tailored to multi-agent setups.
 Encryption.
-Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [67].
-Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts.
+Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [feretzakis2024privacy].
+Agentic workflows often include the inter-agent handoff of partially processed results, models, or prompts.
 Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
 Access control.
-Access conrol is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
-For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [57].
-Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [157] and Attribute-Based Access Control (ABAC) [157] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
+Access control is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities.
+For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [wu2023autogen].
+Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [khan2024role] and Attribute-Based Access Control (ABAC) [khan2024role] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
 Adversarial learning.
-Adversarial learning is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results.
-Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially designing outputs from one compromised agent propagating misleading information to others [164].
-Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities.
-Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
+Adversarial threats are increasing for LLM-based agents, including prompt injection, poisoned tool outputs, and disruption via malformed intermediate results.
+Recent work shows that compromising one agent can propagate misleading outputs and destabilize an entire multi-agent framework [zhang2024breaking].
+Mitigations include adversarial training (e.g., input perturbation, reward shaping, contrastive learning) and practical safeguards such as enforcing safety constraints and validating tool outputs before execution.
 Runtime Monitoring.
 Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity.
-Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [165].
+Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [zhen2025novel].
 For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@
 Differential privacy (DP).
-DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified.
-In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information.
-For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [166].
-DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
+DP protects individuals by adding controlled noise, and in multi-agent LLM systems it can be used during training or inference to limit what agents can reveal to each other.
+Using approaches like DP in federated learning, DP-SGD, and an $\epsilon$ privacy budget, agents can collaboratively train, fine-tune, or coordinate while reducing exposure of raw or identifiable data [lazaros2024federated].
 Data minimization.
 Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution.
-For instance, temporary memory buffers used in systems like ChatDev [30] or ReAct-based pipelines [75] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data.
-Furthermore, anonymization and pseudony [167] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
+For instance, temporary memory buffers used in systems like ChatDev [qian2023chatdev] or ReAct-based pipelines [yao2023reactsynergizingreasoningacting] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data.
+Furthermore, anonymization and pseudony [luzon2024tutorial] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
 Secure computation.
-Techniques including secure multi-party computation (SMPC) [168], homomorphic encryption [169], and trusted execution environments (TEEs) [^11] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
+Techniques including secure multi-party computation (SMPC) [evans2018pragmatic], homomorphic encryption [acar2018survey], and trusted execution environments (TEEs) [^13] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy.
 In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage.
 Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
-Privacy-by-design [170] principles are becoming central to the engineering of next-generation Agentic AI systems.
+Privacy-by-design [schaar2010privacy] principles are becoming central to the engineering of next-generation Agentic AI systems.
 Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share.
 As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
-Next, we discuss the compliance to AI goverance and policy in Agentic AI systems.
 Table 7: Governance controls for agentic AI systems: pillar $\rightarrow$ control $\rightarrow$ operationalization $\rightarrow$ evidence $\rightarrow$ standards.
@@
 Governance Org.
-accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [171]; ISO/IEC 42001 [2]
+accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001 [iso42001]
 Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts.
-11–14 [125]; ISO/IEC 42001 [2]
-Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [129]; EU AI Act Art.
-9 [125]
+11–14 [eu_ai_act]; ISO/IEC 42001 [iso42001]
+Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [ISO_IEC_42005_2025]; EU AI Act Art.
+9 [eu_ai_act]
 Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts.
-72–73 [125]; NIST AI RMF (Manage) [171]
-Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [34]; ISO/IEC 5259 series
-Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [125]; ISO/IEC TR 24027
-Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [155]; NIST AI RMF (Map/Measure) [171]
-Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [125]
+72–73 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
+Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art.
+25 [gdpr25]; ISO/IEC 5259 series
+Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [eu_ai_act]; ISO/IEC TR 24027
+Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [phillips2021four]; NIST AI RMF (Map/Measure) [tabassi2023ai]
+Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts.
+11–12; Annex IV [eu_ai_act]
 Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts.
-12–13 [125]; ISO/IEC 42001 [2]
+12–13 [eu_ai_act]; ISO/IEC 42001 [iso42001]
 Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art.
-14 [125]; NIST AI RMF (Manage) [171]
+14 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
 Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art.
-15 [125]; ISO/IEC 24029-1 [32]
-Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [137]; HIPAA §164 [35]
+15 [eu_ai_act]; ISO/IEC 24029-1 [iso24029-1]
+Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [owasp2024llmtop10]; HIPAA §164 [hipaa164]
 Security Tool access control RBAC/ABAC per agent/role; per-tool API tokens; argument validation; allowlists Access matrices; token scopes; arg-validator tests NIST SP 800-218/218A; NIST AI RMF (Manage)
 Security Isolation/sandboxing Containers/VMs; syscall/network/file egress policies; hardware TEEs for sensitive code Sandbox configs; egress policy; attestation logs NIST SP 800-218A; TEEs best-practice notes
-Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [34]; OECD principles [127]
+Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art.
+25 [gdpr25]; OECD principles [oecd_ai_catalogue]
 Supply Chain Provenance & SBOM Track sources and hashes for models/tools/datasets; dependency scanning SBOM; third-party attestations; license checks NIST SSDF (SP 800-218/218A); ISO/IEC 42001
-Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [171]; ISO/IEC 42001
-Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [32]
-Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [137]; NIST AI RMF (Manage)
-Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [127]; EU AI Act transparency
-Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [171]; ISO/IEC TR 24027
-Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [125]; ISO/IEC 42001
-GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [125]; ISO/IEC 42001
+Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001
+Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [iso24029-1]
+Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [owasp2024llmtop10]; NIST AI RMF (Manage)
+Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [oecd_ai_catalogue]; EU AI Act transparency
+Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [tabassi2023ai]; ISO/IEC TR 24027
+Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [eu_ai_act]; ISO/IEC 42001
+GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [eu_ai_act]; ISO/IEC 42001
 Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art.
-52 [125]; C2PA121212c2pa.org
-Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [34]
-Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [34]
-Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [171]
+52 [eu_ai_act]; C2PA141414c2pa.org
+Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [gdpr25]
+Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [gdpr25]
+Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [tabassi2023ai]
 Supply Chain Third-party/vendor risk Vendor questionnaires; DPAs; penetration/attestation reports; license compliance Vendor risk register; DPA file; attestation/SBOM NIST SSDF (SP 800-218/218A); ISO/IEC 42001
-Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210131313ISO 9241-210; OECD principles [127]
+Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210151515ISO 9241-210; OECD principles [oecd_ai_catalogue]
 Sustainability Energy/CO2 telemetry Track/normalize energy use and CO2; report per release; set budgets Energy/CO2 logs; budget vs. actual; disclosure note Org policy / reporting guidance
-8 Compliance to AI Governance and Policy in Agentic AI Systems
+6.3 Evaluation Benchmarks and Metrics for Agentic AI Security
+Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness.
+This subsection reviews current evaluation frameworks and metrics related to agentic AI security, aligned with the TRiSM pillars.
+Attack Evaluation Benchmarks
+Several benchmark datasets have emerged to evaluate adversarial attacks against LLM-based agents.
+For prompt injection and jailbreak scenarios, HarmBench [mazeika2024harmbench] and JailbreakBench [chao2024jailbreakbench] provide standardized datasets for evaluating robustness against adversarial prompts, including role-playing, obfuscation, and multi-turn manipulation scenarios particularly relevant to agentic systems.
+Tool misuse is addressed by ToolBench [xu2023tool] and API-Bank [li2023apibank], which evaluate agents’ ability to use external tools correctly while resisting adversarial instructions that trigger unauthorized access or malicious function calls.
+Multi-agent adversarial scenarios are covered by AgentBench [liu2023agentbench] and GAIA [mialon2023gaia], which test resilience against coordinated attacks, impersonation, and information manipulation in multi-agent environments.
+Recent datasets also evaluate vulnerabilities specific to agentic architectures, including memory poisoning attacks where adversaries inject malicious information into agent memory stores, and context manipulation where attack payloads are embedded in retrieved documents or tool outputs [raza2025justhumansneedvaccines].
+Defense Evaluation Metrics
+Standard metrics for assessing defense mechanisms have been established across the research community.
+Attack Success Rate (ASR) measures the percentage of attack attempts that successfully compromise agent behavior, bypass safety constraints, or extract sensitive information, with lower ASR indicating more robust defenses [saha2025breaking].
+Robustness Score quantifies the agent’s ability to maintain intended behavior under adversarial conditions, typically measured as the degradation in task performance when subjected to attacks compared to benign scenarios [raza2025humanibench].
+Defense Effectiveness Rate captures the proportion of attacks successfully detected and mitigated by defense mechanisms, including input validation, output filtering, and behavioral monitoring systems.
+For production deployments, False Positive and False Negative Rates are critical, measuring the rate at which legitimate user requests are incorrectly flagged as attacks and actual attacks that evade detection, respectively.
+Our proposed Composite Security Score (CSS), introduced in Section 5 provides a holistic assessment by combining multiple security dimensions including input validation coverage, tool access control effectiveness, memory integrity, and inter-agent authentication strength.
+Table 8 provides a systematic comparison of these benchmarks, including their coverage of attack scenarios, applicable defense metrics, and alignment with TRiSM pillars.
+Table 8: Evaluation Benchmarks for Agentic AI Security
+Benchmark Attack Coverage Metrics TRiSM Pillars Limitations
+HarmBench [mazeika2024harmbench] Prompt injection, jailbreaks ASR, robustness Security, Explainability Limited multi-agent
+ToolBench [xu2023tool] Tool misuse, unauthorized API Effectiveness, FP/FN Security, Privacy Tool-focused only
+AgentBench [liu2023agentbench] Multi-step attacks, coordination Task success, compliance All pillars Limited adversarial
+GAIA [mialon2023gaia] Ambiguous instructions, reasoning Completion, safety Explainability, Governance Early stage
+WebArena [zhou2023webarena] Web attacks, malicious content Success, safety Security, Privacy Domain-specific
+HELM [liang2022helm] Robustness, fairness, bias Comprehensive ModelOps, Governance Not agent-specific
+MLCommons [mlcommons2024] Harmful content, privacy leakage Safety, privacy scores Security, Privacy, Governance Evolving
+Key Insights
+This section presented a defense-in-depth view of AMAS security spanning data confidentiality (encryption and secure enclaves), access control (RBAC/ABAC with least privilege and dynamic trust), adversarial resilience (prompt/tool-output validation and training-based hardening), and runtime monitoring (auditing, anomaly detection, and trust scoring).
+Privacy protection was framed as a system property: differential privacy to limit re-identification risk, data minimization through scoped/ephemeral memory and pseudonymization, and secure computation (SMPC, homomorphic encryption, and TEEs) to enable collaboration without raw-data exposure.
+We also summarized emerging evaluation practice, such as benchmarks for prompt injection, tool misuse, and multi-agent manipulation, together with metrics such as ASR, robustness degradation, defense effectiveness, and FP/FN rates; to quantify both vulnerabilities and mitigation quality in a TRiSM-aligned manner.
+Overall, security and privacy shift from optional add-ons to foundational requirements for regulated deployments.
+7 Compliance with AI Governance and Policy Frameworks for Agentic AI Systems
 Recent industry surveys indicate that enterprise leaders deploying Agentic AI at scale rank security, trust, compliance, and oversight among their top concerns.
@@
 Regulatory frameworks provide baseline requirements that all AI systems, including agent-based architectures, must meet. Key frameworks include:
-  • NIST AI Risk Management Framework (AI RMF) [171].
+  • NIST AI Risk Management Framework (AI RMF) [tabassi2023ai].
 A voluntary U.S.
@@
 Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
-  • EU AI Act [125].
+  • EU AI Act [eu_ai_act].
 A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026.
@@
 7.2 Auditability
-Multi-agent systems can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
+AMAS can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
   • Comprehensive logging.
@@
 To meet audit and regulator expectations, maintain:
-  • Technical documentation (EU AI Act [1]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
+  • Technical documentation (EU AI Act [EU_AI_Act2025]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
   • Model/Data cards with lineage, licenses, and known hazards; evaluation reports with stress tests (robustness, red-teaming), calibration/fairness metrics, and residual risk.
@@
   • Serious-incident handling. Define thresholds, triage workflows, and reporting timelines; keep case files linking logs, affected users, and remediation.
-  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [129], ISO/IEC 23894 [130] ) when capabilities, data, or context change materially.
+  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [ISO_IEC_42005_2025], ISO/IEC 23894 [ISOIEC23894_2023] ) when capabilities, data, or context change materially.
 7.6 Supply-Chain & Secure Development (GenAI)
@@
 Follow NIST SP 800-218 (SSDF) [^19] NIST SP 800-218 (SSDF) and the GenAI profile SP 800-218A  [^20] GenAI profile SP 800-218A for model/system development; integrate threat modeling for agent tool-chains and LLM-specific risks.
-  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [137] Top 10 for LLM applications.
+  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [owasp2024llmtop10] Top 10 for LLM applications.
   • Provenance. Track model/tool/dataset origins, licenses, and checksums; require attestations for third-party components and run dependency/secret scans in CI.
@@
 This combination creates evidenceable trust and reduces the blast radius of inevitable failures.
+Key Insights
+This section operationalized compliance through systematic mappings between TRiSM pillars and regulatory instruments (EU AI Act, NIST AI RMF, ISO/IEC 42001, GDPR, HIPAA).
+Comprehensive auditability via immutable logging and decision provenance graphs enables regulatory audit and incident investigation.
+Policy-as-code mechanisms encode runtime constraints, preventing violations through pre-execution validation and human sign-off requirements for high-impact operations.
+Documentation rigor (technical dossiers, model cards, evaluation reports) provides regulator-ready evidence packages, while post-market monitoring implements continuous telemetry, incident handling workflows, and periodic reassessment protocols.
+Supply chain security addresses provenance verification through Software Bill of Materials documentation and dependency scanning.
+This integrated lifecycle program transforms compliance from checkbox exercise to strategic capability, building evidenceable trust through systematic controls and adaptive risk management.
 8 Discussion
-In this section, we present our findings and outline directions for future work.
+In this section, we summarize key findings and outline implications and future directions for deploying TRiSM in AMAS.
 8.1 Technical Implications of TRiSM for Agentic AI Design
-Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [172].
-For example, there is discussion on designing specialized “guardian agents” within AMAS [42].
-In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g.
-blocking disallowed actions such as outputting personally identifiable information).
-Prior research highlights risks like “excessive agency” [138] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals).
-TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds.
-Likewise, emerging threats specific to Agentic AI systems such as prompt injection attacks, memory poisoning , or cascading hallucinations, highlights the need for more built in risk controls.
-By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g.
-masking a sensitive datum or stopping an unsafe action).
-9.2 Ethical and Societal Ramifications of Multi-Agent AI
-Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions.
-Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [173], all of which are vital for public trust in AI systems.
-Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”.
-This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed .
-Recent guidance on trustworthy AI [174] often highlights accountability and explainability as key pillars of trust.
-In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [175].
-Human oversight is another ethical imperative tightly coupled with accountability.
-TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [176].
-In AMAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly .
-Therisk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [177].
-TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
+Rather than treating LLM agents as black-box decision-makers, TRiSM encourages designing agents with continuous oversight guardrails [guardrails_ai_guardrails_2024].
+Recent discussions also highlight the value of specialized “guardian agents” within AMAS [securiti2025aitrism], which can filter sensitive data, establish baselines of normal behavior, and enforce runtime policies (e.g., blocking disallowed actions such as exposing personally identifiable information).
+Prior work also warns about “excessive agency” [owasp2025agentic], where overly autonomous agents with broad tool access may trigger unintended harmful actions (e.g., hallucination-driven execution or misinterpreted goals).
+TRiSM mitigates these risks by limiting autonomy through role separation, scoped permissions, and safety constraints.
+Emerging AMAS-specific threats such as prompt injection, memory poisoning, and cascading hallucinations further motivate built-in risk controls.
+8.2 Ethical and Societal Ramifications
+Beyond technical risks, deploying networks of autonomous agents raises ethical and societal concerns around accountability, fairness, and human oversight [raza2025responsible].
+TRiSM governance emphasizes that organizations must retain responsibility for AI actions and avoid obscuring accountability behind opaque decision-making.
+In practice, each autonomous agentic decisions should be transparent enough to be understood and challenged by human reviewers when necessary [vayani2025all].
+TRiSM does not remove humans from the loop; instead, it structures human-agent collaboration through defined oversight roles and escalation pathways [raza2025humanibench].
+In AMAS deployments, humans may monitor agent swarms in real time, pause or shut down anomalous behavior, and adjust policies on the fly.
+The risk of user complacency, over-trusting autonomous agents; has also been noted as a hazard [parasuraman2010complacency], and TRiSM counteracts this by formalizing oversight and supervision requirements.
+8.3 Critical Analysis of Evaluation Approaches
+The evaluation framework in Sections 5 and 6.3 supports targeted diagnosis by separating trustworthiness, explainability, user-centered performance, and coordination metrics, rather than relying on monolithic accuracy [raza2024exploring].
+The proposed CSS and TUE further address gaps in capturing multi-agent phenomena that single-agent metrics cannot represent [ribeiro2016should, lundberg2017unified].
+However, several limitations remain.
+Human-centered evaluation is resource-intensive and does not scale to continuous deployment, and AMAS evaluation often lacks ground truth for emergent coordination behaviors.
+Improvements in one dimension may degrade another (e.g., verbose explanations increasing latency), and composite weights vary by domain, requiring clearer guidance for practitioners.
+Current protocols also struggle to capture behavioral drift over time, and subjective measures often omit inter-rater reliability.
+Finally, evaluation metrics may be gamed (Goodhart’s Law [^21] Goodhart’s law.), especially when systems optimize for a single composite score rather than underlying capability.
+8.4 Implementation Challenges and Real-World Constraints
+While TRiSM provides a strong foundation, deploying it in production introduces practical constraints.
+Continuous monitoring, validation, and anomaly detection add computational overhead and latency, which can harm user experience in time-sensitive settings.
+Large-scale enterprise deployments may involve hundreds or thousands of agents, increasing compute, storage, networking, and staffing costs.
+Integration is also challenging because many organizations rely on legacy infrastructure or third-party frameworks not designed for agentic security; adding memory protection, access control, and behavioral monitoring often requires cross-component changes and ongoing maintenance.
+Operationally, monitoring can produce large volumes of alerts, including false positives, which may overwhelm security teams.
+A phased, risk-driven adoption strategy is therefore often more feasible: start with input validation, restricted tool access, and logging, and then add stronger controls as system maturity and risk exposure increase.
 8.5 Alignment with Emerging AI Regulations and Standards
-The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI.
-This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards.
-For example, the European Union’s AI Act [125] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems.
-These are precisely the capabilities that a TRiSM approach promotes.
-Deploying Agentic AI systems responsibly requires a multifaceted governance approach.
-By adhering to established regulatory standards (and keeping abreast of new ones), by ensuring auditability for transparency and accountability, and by enforcing policies through technical and organizational means, we can significantly mitigate the risks posed by autonomous AI agents.
-Moreover, embracing evolving best practices – from governance boards to adaptive, code-driven policies – will help maintain control over these systems as they grow in complexity.
-The end goal is to harness the benefits of Agentic AI (increased efficiency, creativity, and problem-solving at scale) while staying within legal, ethical, and safe bounds, thereby preserving public trust and meeting compliance obligations.
-9.4 Limitations and Current Research Gaps
-While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research.
-First, the classical AI domains that works on standard test suites, there is no consensus on how to measure an AI Agent vs the Agentic AI ability to operate safely under TRiSM principles.
-This makes it difficult to compare different governance strategies or to track progress objectively.
-We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [134] or “red-teaming” exercises [154] that can stress-test Agentic AI systems and score their resilience.
-There is also a lack of real-world validation for many TRiSM-inspired controls.
-Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [65].
-It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
-Another critical gap is adversarial learning.
-As we improve defenses, attackers will inevitably adapt [178].
-Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [38].
-Beyond technical aspects, the organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures.
-Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent.
-Without a strong organizational commitment, even the best technical framework can falter.
+TRiSM principles align closely with emerging regulatory frameworks and standards.
+For example, the EU AI Act [eu_ai_act] emphasizes risk management, transparency, data governance, and human oversight for high-risk systems, for example capabilities directly supported by TRiSM-aligned governance and controls [yang2025agentic, farooq2025evaluating].
+This convergence suggests that TRiSM-based practices can help organizations translate technical safeguards into compliance-ready processes and evidence.
+Despite progress, some gaps remain.
+For example, there is still no consensus on standardized benchmarks and metrics for evaluating agentic systems under TRiSM principles, which limits cross-study comparison and objective progress tracking.
+Many controls also lack real-world validation beyond lab settings or narrowly scoped tasks, and adversarial learning remains an ongoing challenge as attackers adapt to defenses.
+Future work should prioritize standardized stress-testing protocols (e.g., red-teaming), longitudinal evaluation to detect drift, and deployment case studies that report both effectiveness and operational costs.
+Extension to Multimodal and Embodied Agents.
+While this survey focuses on LLM-based AMAS, the Agentic AI landscape increasingly encompasses multimodal agents (e.g., vision-language models such as LLaVA, GPT-4V) and embodied agents [xia2018gibson] operating in physical environments (e.g., robotic systems, autonomous vehicles, industrial automation).
+Future work should systematically extend TRiSM frameworks to address the unique challenges these systems present.
+Multimodal agents introduce additional security considerations including cross-modal adversarial attacks (e.g., adversarial images triggering unsafe text outputs), visual hallucinations, and sensor spoofing vulnerabilities.
+Embodied agents face physical safety risks, real-time decision-making constraints under uncertainty, sim-to-real transfer gaps, and complex human-robot trust calibration requirements.
+Extending the proposed CSS and TUE metrics to capture sensor fusion quality, physical action safety scores, and human-robot collaboration efficacy represents a promising direction for comprehensive agentic AI evaluation.
 8.6 Future Roadmap for Agentic AI TRiSM
@@
 The community should create open benchmarks and challenge environments to test AMAS governance.
-For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed Agentic AI systemthat performs relative to one without such controls.
+For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed ASAM that performs relative to one without such controls.
 This will enable direct comparisons and drive progress on measurable metrics of trust (e.g.
@@
 Techniques from cybersecurity (e.g.
-adversarial training, AI model “penetration testing” [179], and formal verification) should be integrated into the LLM agent development pipeline.
+adversarial training, AI model “penetration testing” [bacudio2011overview], and formal verification) should be integrated into the LLM agent development pipeline.
 Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions.
-Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [180] , can help discover vulnerabilities in a controlled way before real adversaries do.
+Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [schneider2020wargaming] , can help discover vulnerabilities in a controlled way before real adversaries do.
 We encourage designing better interfaces and protocols for human oversight of Agentic AI systems.
-Borrowing from human-computer interaction [181] and cognitive engineering [182], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
+Borrowing from human-computer interaction [AmershiWVFNCSIB19] and cognitive engineering [hexmoor2025behaviour], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
 Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials.
-These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators.
-Insights from such pilots can inform refinements in both technical standards and regulations.
 There is much to learn from other high-stakes domains.
@@
 We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
-A cross-cutting concept relevant to both Governance and Lifecycle Management, controllability captures the ability of humans and oversight systems to steer or override agentic AI.
-Mechanisms include stop/pause functions, rollback gates role-based oversight, and auditability of inter-agent actions.
-In practice, controllability is critical to ensuring accountability and preventing runaway behaviors in multi-agent workflows.
 9 Conclusion
-The rapid advancement of Agentic AI, driven by AMAS, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services.
-However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations.
-This review has explored these challenges, proposing a TRiSM framework structured around five pillars: Explainability, ModelOps, Security, Privacy and their lifecycle Governance, adapted for LLM-based AMAS.
-Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency.
-The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
-Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination.
-By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
-Acknowledgements
-Funding for the research was provided through Horizon Europe project AIXpert: An agentic, multi-layer, GenAI-powered backbone to make an AI system explainable, accountable, and transparent (ID:101214389).
+Agentic AI systems built as AMAS are rapidly changing how autonomous systems collaborate, plan, and make decisions in high-stakes domains such as healthcare, finance, and public services.
+At the same time, these systems expand the risk surface and raise new challenges in trust, risk, and security management (TRiSM).
+This review addressed those challenges and introduced a TRiSM-oriented framework for LLM-based AMAS organized around five pillars: Explainability, ModelOps, Security, Privacy, and lifecycle Governance.
+We contributed a risk taxonomy that captures AMAS-specific threats, including prompt injection, memory poisoning, and collusive or cascading failures.
+We also proposed practical evaluation metrics to measure inter-agent enablement and the correctness and efficiency of tool use.
+In addition, we summarized key technical mechanisms to support TRiSM in practice, including explainability methods and security/privacy protections.
+Looking ahead, progress will depend on stronger adversarial robustness, clearer governance and audit protocols, and standardized benchmarks that evaluate both trustworthiness and coordination under realistic deployment conditions.
+Aligning Agentic AI development with TRiSM principles can help ensure these systems remain not only capable and scalable, but also safe, transparent, and accountable for real-world use.
+Acknowledgments
+This research was funded by the European Union’s Horizon Europe research and innovation programme under the AIXPERT project (Grant Agreement No.
+101214389), which aims to develop an agentic, multi-layered, GenAI-powered framework for creating explainable, accountable, and transparent AI systems.
+Additionally, this work was supported in part by the National Science Foundation (NSF); in part by United States Department of Agriculture (USDA); in part by the National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” Program, Accession Number 1029004 for the Project Titled “Robotic Blossom Thinning with Soft Manipulators” under Award AWD003473, Award AWD004595, and Award USDA-NIFA; and in part by United States Department of Agriculture National Science Foundation (USDANSF), Accession Number 1031712, under the Project “ExPanding University of Central Florida (UCF) AI Research To Novel Agricultural EngineeRing Applications (PARTNER)” under Grant 2024-67022-41788.
 Declaration of Competing Interest
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Author Contributions
+S.R.
+and R.S.
+contributed equally to this work.
+S.R.: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing, Revision R.S: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing M.K.: Supervision, Detailed review, Structuring, Writing.
+C.E.: Detailed review, Writing, Re-writing, Supervision, Revision.
+All authors have read and approved the final manuscript.
 Appendix
-Table A.1: Key terminologies for LLM-based agentic AI systems.
+Table A.1: List of abbreviations used in this paper.
+Abbreviation Description
+AI Artificial Intelligence
+AMAS Agentic Multi-Agent Systems
+CoT Chain-of-Thought
+CSS Component Synergy Score
+DP Differential Privacy
+FHE Fully Homomorphic Encryption
+HITL Human-in-the-Loop
+HE Homomorphic Encryption
+IAM Identity and Access Management
+LLM Large Language Model
+MLOps Machine Learning Operations
+ModelOps Model Lifecycle Operations, Monitoring, and Governance
+MPC Secure Multi-Party Computation
+PII Personally Identifiable Information
+RAG Retrieval-Augmented Generation
+RBAC Role-Based Access Control
+RMF Risk Management Framework
+TEEs Trusted Execution Environments
+TRiSM Trust, Risk, and Security Management
+TUE Tool Utilization Efficacy
+XAI Explainable Artificial Intelligence
+Figure A.1: Traditional AI Agent vs. Agentic AI.
+Table A.2: Key terminologies for LLM‐based agentic AI systems.
 Term Definition
-Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
-Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
-Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
-Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
-Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
-Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
-Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
-ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
-Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
-Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
-Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
-Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
-Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
-Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
-Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
-Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
-Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
-Figure A.1: Traditional AI Agent vs. Agentic AI.
+Agentic AI System A multi-agent artificial intelligence system powered by large language models (LLMs), in which multiple autonomous agents coordinate, communicate, and reason collectively to plan and execute tasks over extended horizons.
+Such systems typically incorporate persistent memory, tool use, and dynamic role assignment to support adaptive, goal-driven behavior.
+Autonomy Model The internal decision-making mechanism that governs how an agent selects goals, plans actions, and executes tasks without continuous human intervention, often leveraging planning, reasoning traces, feedback signals, and environmental context.
+Chain of Thought (CoT) A prompting and reasoning paradigm in which an LLM explicitly generates intermediate reasoning steps before producing a final output, enabling multi-step problem solving, improved transparency, and more controllable agent behavior.
+Counterfactual Analysis An interpretability and diagnostic technique that evaluates how changes to inputs, agent roles, or internal states would alter system outcomes, helping reveal causal dependencies, failure modes, and contribution disparities across agents.
+Explainability The ability of an AI system or individual agent to provide human-understandable rationales for its decisions, actions, or outputs, often supported by techniques such as feature attribution, decision provenance, or structured reasoning traces.
+Foundation Model (LLM) A large pretrained language model (e.g., GPT-4, LLaMA) that serves as the core reasoning and generation component of an agent, providing linguistic competence, contextual understanding, tool invocation, and general-purpose reasoning capabilities.
+Shared Memory (Persistent Memory) A centralized or distributed memory substrate, commonly implemented using vector databases or structured logs, that enables agents to store, retrieve, and share contextual information across time, supporting long-term coherence and coordination.
+ModelOps A set of operational practices for managing models and agent prompts throughout their lifecycle, including versioning, deployment, monitoring, auditing, and retirement, with an emphasis on reliability, reproducibility, and governance.
+Application Security Protective mechanisms and design practices that safeguard agentic systems against adversarial behaviors such as prompt injection, identity spoofing, privilege escalation, and cross-agent manipulation.
+Model Privacy Privacy-preserving techniques that ensure sensitive data remains protected during model training, inference, and inter-agent communication, including differential privacy, secure computation, and trusted execution environments.
+Prompt Injection A class of security attacks in which maliciously crafted inputs introduce hidden or indirect instructions that override or corrupt an agent’s intended behavior, potentially propagating harmful actions across agent interactions.
+Retrieval-Augmented Generation (RAG) An architectural approach in which agents dynamically retrieve relevant external knowledge (e.g., documents or facts) from a datastore and condition LLM generation on this retrieved context to improve factuality and grounding.
+Role-Specialized Coordination A coordination strategy where agents are assigned distinct functional roles (e.g., planner, verifier, executor) and interact via structured communication protocols to collectively solve complex tasks more efficiently and robustly.
+Decision Provenance Graph A structured, graph-based representation that records data flows, reasoning steps, and inter-agent interactions, enabling post-hoc auditing, accountability analysis, and system-level interpretability.
+Tool-Use Interface The mechanism that allows agents to issue structured commands to external tools, APIs, or execution environments and integrate returned results into subsequent reasoning and decision-making steps.
+Trust Score A composite quantitative indicator of an agent’s reliability and alignment, typically combining task performance, safety compliance, behavioral consistency, and calibration of expressed confidence over time.
+Composite Metric An aggregated evaluation measure that combines multiple dimensions—such as trustworthiness, explainability, coordination quality, and user-centered performance—to benchmark and compare agentic AI systems holistically.